### v2

# Generating Shrinkers by Mean Curvature Flow

Abstract.
We prove existence for many examples of shrinkers by producing compact, smoothly embedded surfaces that, under mean curvature flow, develop singularities at which the shrinkers occur as blowups.
Key words and phrases:
Mean curvature flow, shrinkers, self-shrinkers, level set flow
2020 Mathematics Subject Classification:
Primary 53E10
F. Martín’s research is partially supported by the framework of IMAG-María de Maeztu grant CEX2020-001105-M funded by MCIN/AEI/ 10.13039/50110001103. Some of this work was carried out while the authors were visitors at the Simons Laufer Mathematical Sciences Institute (SLMath, formerly MSRI) in Berkeley, CA during the Fall 2024 semester, in a program supported by National Science Foundation Grant No. DMS-1928930.
1. Introduction
The formation of singularities in a mean curvature flow $t\mapsto M(t)$ with smooth, compact initial data is a fundamental phenomenon that involves the underlying geometry and topology of the evolving surfaces. As the flow progresses, regions of high curvature develop, leading to singularities where the surface either pinches off, collapses, or forms more complex structures. If a singularity occurs at a spacetime point $(p,T)$, then, as $t\uparrow T$, suitable rescalings of $M(t)$ about $p$ converge (at least weakly and subsequentially) to a shrinker, i.e., a surface $\Sigma$ such that
$t\in(-\infty,0)\mapsto|t|^{1/2}\Sigma$
is a mean curvature flow. Thus, to understand mean curvature flow singularities, it is natural to try to understand all possible shrinkers. Many interesting examples of shrinkers have been constructed, and there are some notable theorems classifying certain kinds of shrinkers. In this paper, we present a unified approach to producing many of the known examples, as well as a new family of examples.
Almost all of the known examples of shrinkers have been produced by one of the following three methods:
  1. (1)
    Desingularization. The idea is that one takes two embedded shrinkers that intersect along one or more curves, and one tries to replace intersecting ribbons of surface by a surface modelled on Scherks’s singly periodic surface. See [kkm, nguyen]
  2. (2)
    Doubling. Here one takes a known embedded shrinker, and produces new examples that look like two slighly perturbed copies of that shrinker joined by suitably placed catenoidal necks. See [kapouleas-doubling].
  3. (3)
    Minimax. See [ketover-platonic, ketover-fat, buzano].
Those methods are based on the fact that a surface is a shrinker if and only if it is minimal with respect to a certain Riemannian metric (the so-called shrinker metric). All three methods are powerful and have produced many interesting examples, but they have some limitations. In particular, desingularization and doubling only produce examples when some parameter, typically the genus, is very large. Also, all three methods leave open the question of whether the shrinker actually arises as a blowup of the mean curvature flow of an initially smooth, compact surface. This last defect has been somewhat overcome by recent work of Tang-Kai Lee and Xinrui Zhao[lee-zhao], who show that any shrinker that satisfies a natural and rather mild hypothesis does indeed arise as such a blowup. (For $2$-dimensional shrinkers in $\mathbf{R}^{3}$, the only case considered in this paper, the hypothesis is that no end of the shrinker is asymptotically cylindrical.) In practice, it is typically easy to show that the hypothesis holds for the examples one produces provided the genus is sufficiently large. But for low genus, one generally does not know whether the hypothesis holds.
(Of course, if a shrinker is itself a compact, embedded surface, then it trivially arises as such blowup. This is the case for the examples constructed in [kapouleas-doubling].)
In this paper, we produce shrinkers by creating initial surfaces that, under mean curvature flow, develop singularities at which the desired shrinkers occur as blowups. Since we are interested in shrinkers because of mean curvature flow, it is perhaps more natural to produce them by mean curvature flow than by other methods. And, of course, we then know that the resulting shrinkers do arise as blowups since that is how they are produced. In particular, the Platonic shrinkers in Section 4 were first produced by Ketover [ketover-platonic] using minimax, but it was not known until now that they arise as blowups. Similarly, the genus-$g$ examples in Section 7 were produced for high genus by desingularization [kkm, nguyen] and for every genus by minimax [buzano], but, except in the case of high genus, it was not known until now that they arise as blowups. (Indeed, until the recent result of Lee-Zhao [lee-zhao] mentioned above, it was not known whether any of the desingularized examples arose as blowups.)
We now briefly describe the shrinkers produced in this paper. Section 4 produces shrinkers that resemble skeletons of the Platonic solids. See Figure 2. The proof we give also gives shrinkers with the symmetries of a regular $n$-gon. Section 7 produces shrinkers that, for large genus, look like the union of a sphere and a bisecting plane, with the circle of intersection replaced by necks. See Figures 1 and 5. Section 8 produces new examples that are shrinker analogs of some well-known, non-existent minimal surfaces. Specifically, in the 1980s, researchers wondered whether there existed complete embedded minimal surfaces that were, roughly speaking, Costa-Hoffman-Meeks surfaces with additional handles. In particular, for each $g$, the new surface (if it existed) would have the same symmetries as the genus-$g$ Costa-Hoffman-Meeks surface and the same number of ends ($3$), but it would have twice the genus, namely $2g$. Martín and Weber [martin-weber] proved that such minimal surfaces do not exist. By contrast, we show that (for large $g$) there are shrinkers that have the defining properties of those non-existent minimal surfaces. See Figure 7.
We are aware of two other papers that use mean curvature flow to prove existence of shrinkers: [chu-sun] proves existence of shrinkers of genus one with less entropy than the Angenent tori, and [iw-fat] proves existence of certain genus-$g$ shrinkers that cause fattening when $g$ is large. On the other hand, Drugan and Nguyen [drugan-nguyen] used a flow other than mean curvature flow to prove existence of rotationally invariant shrinking tori. Their examples presumably coincide with the Angenent Tori, whose existence was established in [angenent] by a shooting method.
Figure 1. A cross section of a numerical simulation of one of the shrinkers constructed in Section 7. Figure courtesy of Mario Schulz.
2. Preliminaries
[…]
In this paper, “mean curvature flow” means “unit-regular Brakke flow”. Roughly speaking, a unit-regular Brakke flow is an integral Brakke flow that is smooth (without sudden vanishing) in a spacetime neighborhood of each point of Gauss density $1$. (See [white-boundary]*Definition 11.1 for the precise definition.) Any weak limit of unit-regular flows is also unit-regular [white-local].
If $K$ is a closed subset of $\mathbf{R}^{3}$, or, more generally, of a smooth Riemannian $3$-manifold $N$, we let $K(t)$ denote the result of flowing $K$ for time $t$ by the level set flow.
Let $\operatorname{T_{\textnormal{fat}}}:=\operatorname{T_{\textnormal{fat}}}(K)$ be the fattening time, i.e, the infimum of times $t$ such that $K(t)$ has nonempty interior.
Definition 1.
If $M$ is smooth, a regular time for $M$ is a time $t\in[0,\operatorname{T_{\textnormal{fat}}}(M)]$ such that $M(\tau)$ is smooth for all $\tau$ in some neighborhood $[0,\operatorname{T_{\textnormal{fat}}}]\cap(t-\epsilon,t+\epsilon)$ of $t$.
If $M$ is smoothly embedded and compact, then 00 is a regular time. By [white-genus], if $t_{1}<t_{2}<\operatorname{T_{\textnormal{fat}}}(M)$ are regular times, then
(1) $\operatorname{genus}(M(t_{2}))\leq\operatorname{genus}(M(t_{1})).$
By [bamler-kleiner], almost every time in $[0,\operatorname{T_{\textnormal{fat}}}(M)]$ is regular, and all tangent flows at times $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ are given by smoothly embedded shrinkers of multiplicity $1$. Thus if $p\in N$ and $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, then every sequence $t_{i}\uparrow T$ has a subsequence such that
(2) $(T-t_{i})^{-1/2}(M(t_{i})-p)$
converges smoothly and with multiplicity $1$ on compact subsets to a shrinker $\Sigma$. If $\tau_{i}/t_{i}$ is bounded away from 00 and away from $\infty$, then
$(T-\tau_{i})^{-1/2}(M(\tau_{i})-p)$
also converges to the same shrinker $\Sigma$. The shrinker $\Sigma$ is a plane if and only if $(p,T)$ is a regular point of the flow.
(If we are working in a general ambient manifold $N$, then the expression (2) makes sense provided we regard $N$ as isometrically embedded in a Euclidean space.)
Definition 2.
For $t\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, we say $\Sigma$ is a shrinker to $M(\cdot)$ at the spacetime point $(p,t)$ if there exist a sequence $t_{i}\uparrow t$, such that the surfaces (2) converge smoothly to $\Sigma$.
We expect there to be only one shrinker at each spacetime point, but that is not known in all cases. (See [chodosh-schulze] for cases in which uniqueness is known.) However, all the shrinkers at $(p,T)$ have the same genus (by Theorem 64 in the appendix), so the following definition makes sense.
Definition 3.
The genus of a spacetime point $(p,t)$ with $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ is the number $g$ such that each shrinker at $(p,t)$ has genus $g$.
At a singular point of genus 00, the shrinker is either a sphere or a cylinder [brendle]. (Earlier, Kleene and Møller proved the result assuming an axis of rotational symmetry [kleene-moller].) Furthermore, the shrinker is unique, by [huisken] (in the spherical case) and [colding-m] (in the cylindrical case). See [zhu] for a simpler proof in the cylindrical case.
The following result is trivial, but important.
Theorem 4.
If $(p,t)$ is a singular point with $t\leq\operatorname{T_{\textnormal{fat}}}$, and if there is an $\epsilon>0$ such that
$\operatorname{genus}(M(\tau)\cap B(p,\epsilon))\leq g$
for every regular time $\tau\in(t-\epsilon,t)$, then $(p,t)$ has genus $\leq g$.
Proof.
Let $t_{i}<t$ be regular times converging to $t$. By passing to a subsequence, we can assume that
$(t-t_{i})^{-1/2}(M(t_{i})-p)$
converges smoothly to a shrinker $\Sigma$. Hence
$(t-t_{i})^{-1/2}(M(t_{i})\cap B(p,\epsilon)-p)$
also converges smoothly to $\Sigma$. Since the surfaces * have genus $\leq g$, the surface $\Sigma$ also has genus $\leq g$. ∎
Corollary 5.
Suppose $T\leq\operatorname{T_{\textnormal{fat}}}$. Then the sum of the genuses of the singularities at time $T$ is less than or equal to the genus of $M$.
If $\operatorname{T_{\textnormal{fat}}}<\infty$, then there is a spacetime singular point $(p,\operatorname{T_{\textnormal{fat}}})$ of positive genus.
Definition 6.
We let $T_{\textnormal{pos}}(M)$ be the infimum of times $t$ at which there is a singularity of positive genus.
The set of singular spacetime points of genus $>0$ is a closed subset of spacetime. Thus if $T_{\textnormal{pos}}(M)<\infty$, then there is spacetime singularity $(p,T_{\textnormal{pos}}(M))$ of positive genus. (In [ccms]*§11, $T_{\textnormal{pos}}(M)$ is written $T_{\textnormal{gen}}(M)$.)
Theorem 7 (Local Regularity Theorem).
[white-local] Suppose $U$ and $W$ are open subset of a smooth Riemannian manifold and that $\overline{U}$ is a compact subset of $W$. Suppose that
$t\in(a,b]\mapsto M(t),$
is a mean curvature flow in which each $M(t)$ is a smooth, properly embedded submanifold of $W$. Suppose for each $n$ that
$t\in(a,b]\mapsto M_{n}(t)$
a unit-regular Brakke flow in $W$ such that $M_{n}(t)$ converges weakly to $M(t)$ for almost every $t\in(a,b)$. Let $a^{\prime}\in(a,b)$. Then, for large $n$, $M_{n}(t)\cap U$ is smooth and converges smoothly to $M(t)$ for all $t\in[a^{\prime},b]$.
Here, the $M_{n}(\cdot)$ and $M(\cdot)$ can be Brakke flows with respect to Riemannian metrics $\gamma_{n}$ and $\gamma$ on $W$, where $\gamma_{n}$ converges smoothly to $\gamma$.
Theorem 8 (Initial Regularity Theorem).
Suppose $t\in[0,T)\mapsto M(t)$ is a mean curvature flow in an open set $U$ such that $M(0)$ is a smooth, propertly embedded, multiplicity $1$ submanifold of $U$. Let $K$ be a compact region in $U$ with smooth boundary transverse to $M(0)$. There is an $\epsilon>0$ with the following properties.
  1. (1)
    The flow $t\in[0,\epsilon]\mapsto M(t)\cap K$ is smooth.
  2. (2)
    if $t\in[0,T)\mapsto M_{i}(t)$ are mean curvature flows that converge weakly to $M(\cdot)$, then for $0<\epsilon^{\prime}<\epsilon$,
    $t\in[\epsilon^{\prime},\epsilon]\mapsto M_{i}(t)\cap K$
    converges smoothly to
    $t\in[\epsilon^{\prime},\epsilon]\mapsto M(t)\cap K.$
  3. (3)
    Suppose in Assertion (2) that $M_{i}(0)$ converges smoothly to $M(0)$. Then
    $t\in[0,\epsilon]\mapsto M_{i}(t)\cap K$
    converges smoothly to
    $t\in[0,\epsilon]\mapsto M(t)\cap K.$
Proof.
For notational simplicity, we give the proof in Euclidean space. Let
$K^{\prime}$ =$=\{\operatorname{dist}(\cdot,K)\leq\delta\},$
where $\delta>0$ is small enough that $K^{\prime}$ is compact subset of $U$ and that $\partial K^{\prime}$ is transverse to $M(0)$.
Claim 1.
Suppose $p_{i}$ converges to $p\in M(0)$ and that $t_{i}>0$ converges to 00. Let $q_{i}$ be the point in $M(0)$ closest to $p_{i}$. Let
$r_{i}=\max\{|p_{i}-q_{i}|,t_{i}^{1/2}\}.$
Translate the flow $M(\cdot)$ in spacetime by $(-q_{i},0)$ and the dilate parabolically by
$(x,t)\mapsto(x/r_{i},t/r_{i}^{2})$
to get a flow $M_{i}^{\prime}(\cdot)$. Then the flows $M_{i}^{\prime}(\cdot)$ converge to the flow
$t\in[0,\infty)\mapsto P,$
where $P$ is the plane $\operatorname{Tan}(M,p)$ with multiplicity $1$, and the convergence is smooth on compact subsets of $\mathbf{R}^{3}\times(0,\infty)$.
To prove the claim, note that, by passing to a subsequence, we can assume that the $M_{i}^{\prime}(\cdot)$ converge weakly to a mean curvature flow
$t\in[0,\infty)\mapsto M^{\prime}(t).$
Note that $M^{\prime}(0)$ is $\operatorname{Tan}(M(0),p)$ with multiplicity $1$. It follows (by Huisken monotonicity) that $M^{\prime}(t)\equiv M^{\prime}(0)$ for all $t\geq 0$. By the Local Regularity Theorem (Theorem 7), the convergence of $M_{i}(\cdot)$ to $M^{\prime}(\cdot)$ is smooth on compact subsets of $\mathbf{R}^{3}\times(0,\infty)$. This completes the proof of Claim 1.
It follows from Claim 1 and from [sw-local]*Corollary A.3 that there is an $\epsilon>0$ such that
(*) The flow $t\in[0,2\epsilon]\mapsto M(t)\cap K^{\prime}$ is smooth.
Assertion (1) follows immediately. Assertion (2) follows from (*2) and the Local Regularity Theorem (Theorem 7).
Now let
$\Omega=(K^{\prime}\setminus\partial K^{\prime})\times(-2\epsilon,2\epsilon).$
Claim 2.
For $j\geq 2$ and $\alpha\in(0,1)$, there exists $C_{j,\alpha}<\infty$ such that
$K_{j,\alpha;\Omega}(M_{n}(\cdot))\leq C_{j,\alpha}$
for all sufficiently large $n$.
See [white-local] for the definition of $K_{2,\alpha;U(R)}(M(\cdot))$. The proof of the claim is almost identical to the proof of [white-local]*Theorem 3.1, so we omit it. Assertion (3) follows immediatetly from Claim 2. ∎
3. Oriented Flows
Suppose that $t\in I\mapsto M(t)$ is a unit-regular Brakke flow. We let $\operatorname{Reg}(t)$ be the set of points $p$ such that the flow is smooth and with multiplicity $1$ in a spacetime neighborhood of $(p,t)$. Equivalently, $\operatorname{Reg}(t)$ is the set of points $p$ such that the Gauss Density at $(p,t)$ is $1$. We let $\operatorname{Sing}(t)$ be the set of $p$ such that the Gauss Density of at $(p,t)$ is $>1$.
Definition 9.
Suppose $t\in I\mapsto M(t)$ is an integral Brakke flow of $m$-dimensional surfaces in an oriented Riemannian $(m+1)$-manifold. An orientation of $M(\cdot)$ is a continuous function that assigns to $p\in\operatorname{Reg}(t)$ a unit normal $\nu(p,t)$ to $p$ at $M(t)$.
A strong orientation is an orientation $\nu$ with the following property. For each time $t$, there is a closed region $K(t)$ of $N$ such that $K(t)\setminus\operatorname{Sing}(t)$ is a smooth manifold-with-boundary, the boundary being $\operatorname{Reg}(t)$, and such that $\nu(\cdot,t)$ is the unit normal that points out of $K(t)$.
Suppose $M$ is a compact, connected embedded, oriented surface in $\mathbf{R}^{3}$ or, more generally, in a simply connected Riemannian $3$-manifold with Ricci curvature bounded below. Let $\nu$ be the unit normal to $M$ given by the orientation. Now $M$ divides the ambient space into two regions. We let $K$ be the closure of the region such that $\nu$ points out of $K$. Then $\nu$ extends to an orientation of the level set flow
$t\in[0,\operatorname{T_{\textnormal{fat}}}]\mapsto M(t)$
as follows. As usual, we let $K(t)$ be the result of flowing $K$ for time $t$ by level flow. If $(p,t)$ is a regular spacetime point of $M(\cdot)$, then there is a neighborhood $B$ of $p$ such that $K(t)\cap B$ is a smooth manifold-with-boundary, the boundary being $M(t)\cap B$. We let $\nu(p,t)$ be the unit normal to $M(t)$ at $p$ that points out of $K(t)$. Note that the orientation is strong. We say that $\nu(\cdot)$ is the standard orientation induced by the given orientation of $M$.
Theorem 10.
Suppose $M_{i}(\cdot)$ is a sequence of unit-regular integral Brakke flows that converge to a unit-regular Brakke flow $M(\cdot)$. Suppose $\nu_{i}$ is an orientation of $M_{i}(\cdot)$. Then, after passing to a subsequence, the $\nu_{i}(\cdot)$ converge to an orientation $\nu(\cdot)$ of $M(\cdot)$.
If the $\nu_{i}$ are strong, then so is $\nu$.
Proof.
This follows easily from the Local Regularity Theorem (Theorem 7). ∎
4. Platonic and Regular Polygon Shrinkers
In this section, we produce shrinkers that have the symmetries of a Platonic solid and that are disjoint from the rays through the centers of the faces of that solid. For each $n\geq 3$, we also produce a shrinker that has the symmetries of a regular $n$-gon in the $xy$-plane and that is disjoint from the rays through the vertices of the $n$-gon. Specifically, we prove
Theorem 11.
Consider a set $Q$ consisting of a finite collection of rays in $\mathbf{R}^{3}$ from the origin. We assume that the rays intersect the unit sphere in one of the following:
  1. (i)
    The centers of the faces of a Platonic solid, or
  2. (ii)
[…]
Let $g$ be number of rays minus $1$. There exists a smoothly embedded shrinker $\Sigma$ such that
  1. (1)
    $\Sigma$ has the symmetries of $Q$ and is disjoint from $Q$.
  2. (2)
    $\Sigma$ has genus $g$.
  3. (3)
    $\Sigma$ occurs as a shrinker in a mean curvature flow $M(\cdot)$ for which $M(0)$ is a compact, smoothly embedded surface in $Q^{c}$.
We do not know whether any of the shrinkers in Theorem 11 are compact. In the cases of the cube and the octahedron, shrinkers as in Theorem 11 were found numerically by Chopp [chopp], and they are compact. See Figure 2. We would guess that all the Platonic solid shrinkers are compact. In the regular $n$-gon case, one can show that for all sufficiently large $n$, the shrinker must be non-compact. Indeed, one can show that as $n\to\infty$, the shrinker converges to a multiplicity $2$ plane.
Figure 2. A cubical shrinker, found numerically by David Chopp. Figure courtesy of David Chopp.
Using minimax methods, Ketover proved existence of shrinkers with properties (1) and (2) in [ketover-platonic] (the Platonic case (i)) and in [ketover-fat] (the regular polygon case (ii)). Assertion (3) in the Platonic solid case is new.
In the regular $n$-gon case (Case [ii]), Theorem 11 was proved, by similar methods to those in this paper, by Ilmanen and White [iw-fat]. However, the surface $M$ in that paper is quite different than the surface $M$ in our proof of Theorem 11: the $M$ in this proof is disjoint from $Q$ and intersects $Z$ in $4$ points, whereas the corresponding $M$ in [iw-fat] intersects every horizontal ray from the origin, and it intersects $Z$ in a pair of points. The shrinkers $\Sigma$ produced here may well be different from the corresponding shrinkers produced in [iw-fat]. In particular, the shrinkers produced in [iw-fat] are all non-compact, whereas the shrinkers produced here might be compact when $g$ is small.
[…]
$H_{1}(Q^{c};\mathbf{Z}_{2})\cong\mathbf{Z}_{2}{}^{g}.$
We let $G$ be the group of symmetries of $Q$:
$G=\{\sigma\in O(3):\sigma(Q)=Q\}.$
If $K$ is a compact set in $\mathbf{R}^{3}$, we let
$T_{Q}=T_{Q}(K)=\inf\{t\geq 0:Q\cap K(t)\neq\emptyset\}.$
If $T_{Q}<\infty$, then $K(T_{Q})\cap Q$ is nonempty.
The following lemma says that if $M$ separates $K$ from $Q$, then $M$ has to bump into $Q$ before $K$ can:
Lemma 12.
Suppose $K$ and $M$ are disjoint compact sets, and that no connected component of $M^{c}$ contains points of $K$ and of $Q$. Then $T_{Q}(M)<T_{Q}(K)$.
Proof.
The proof of this lemma is an immediate consequence of Theorem 5.2 in [white-95]. ∎
Let $\mathscr{A}$ be the class of compact, smoothly embedded surfaces $M$ in $Q^{c}$ such that
  1. (1)
    $\sigma(M)=M$ for every $\sigma\in G$.
  2. (2)
    $\operatorname{genus}(M)\leq g$.
  3. (3)
    $T_{Q}(M)<\infty$.
Theorem 13.
If $M\in\mathscr{A}$ and if $t<\min\{T_{Q},\operatorname{T_{\textnormal{fat}}}\}$ is a regular time of the flow $M(\cdot)$, then $M(t)\in\mathscr{A}$.
Proof.
The theorem is a trivial consequence of the facts that (1) level set flow preserves symmetries, and (2) for any compact smooth embedded surface, $\operatorname{genus}(M(t))$ is a decreasing function of regular times in $[0,\operatorname{T_{\textnormal{fat}}}]$. (See (1).) ∎
Theorem 14.
If $M\in\mathscr{A}$, then there is a connected component $\Sigma$ of $M$ such that the convex hull of $\Sigma$ contains points of $Q$,
Proof.
Since $M(T_{Q})$ contains a point $p$ of $Q$, there must be a connected component $\Sigma$ of $M$ such that $p\in\Sigma(T_{Q})$. Now $\Sigma(t)$ lies in the convex hull $H$ of $\Sigma$ for all $t$, so $p\in\Sigma(T_{Q})\subset H$. ∎
Thus every $M\in\mathscr{A}$ has the properties described by the following theorem:
Theorem 15.
Let $M$ be a $G$-invariant surface properly embedded in $\mathbf{R}^{3}$, disjoint from $Q$, and of genus $\leq g$. Suppose $M$ has a connected component $\Sigma$ whose convex hull contains points of $Q$. Then $\Sigma$ has genus $g$, and if $\Omega$ is a simply connected open subset of $Q^{c}$, then $\operatorname{genus}(M\cap\Omega)=0$.
Proof.
[…]
Let
$\mathscr{T}=\{\overline{W}:\text{$W$ is a component of $\mathbf{R}^{3}\setminus\cup\mathscr{P}$}\}.$
Let $\Delta\in\mathscr{T}$ be such that the interior of $\Delta$ contains a point of $\Sigma$. Exactly one of the edges of $\Delta$ is a ray of $Q$. Label it $E_{1}$. Let $E_{2}$ and $E_{3}$ be the other two edges of $\Delta$. For $i\neq j$ in $\{1,2,3\}$, let $F_{ij}$ the face bounded by rays $E_{i}$ and $E_{j}$.
[…]
For $i\in\{1,2,3\}$, let $K_{i}$ be the union of the $\Delta^{\prime}\in\mathscr{T}$ such that one edge of $\Delta^{\prime}$ is $E_{i}$. Note that $K_{i}$ is a convex cone. Also,
$Q\cap K_{1}$ $=E_{1},$
$Q\cap\operatorname{interior}(K_{i})$ $=\varnothing\quad(i=2,3).$
Let $\Psi$ be a connected component of $\Sigma\cap\Delta$.
Claim 3.
$\Psi$ contains points in $F_{12}$ and points in $F_{13}$
Suppose not. Then (by relabelling) we may assume that it contains no points in the face $F_{12}$. It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{3}$, that $\Sigma$ lies in $\operatorname{interior}(K_{3})$, a convex open set disjoint from $Q$. But that contradicts the hypothesis on $\Sigma$. Thus the claim is proved.
Let $M^{\prime}$ be the union of $\Sigma$ and its images under $G$. Let $U^{\prime}$ be the region bounded by $M^{\prime}$ that is disjoint from $Q$. (Thus, a point in $(M^{\prime})^{c}$ is in $U^{\prime}$ if and only if there is a path from that point to a point in $Q$ that is transverse to $M^{\prime}$ and that intersects $M^{\prime}$ in an odd number of points.) Let $K=M^{\prime}\cup U^{\prime}$. Thus $K$ is smooth $3$-manifold-with-boundary, the boundary being $M^{\prime}$.
Let $\gamma$ be an embedded curve in $\Psi$ such that one of its endpoints is in $F_{12}$ and the other endpoint is in $F_{13}$. Iterated reflections in the planes of $\mathscr{P}$ containing $E_{1}$ produces a simple closed curve in $M^{\prime}\cap K_{1}$ that winds once around the line containing $E_{1}$. Of course, there is such a curve for each ray of $Q$. Let $\mathscr{G}$ be the set of those $g+1$ curves. Note that any $g$ of those $g+1$ curves form a basis for the vector space $H_{1}(Q^{c};\mathbf{Z}_{2})$.
Thus if $\iota$ is the inclusion of $K$ into $Q^{c}$, then the map
(2) $\iota_{\#}:H_{1}(K;\mathbf{Z}_{2})\to H_{1}(Q^{c};\mathbf{Z}_{2}).$
is surjective.
Now
$g\geq\operatorname{genus}(M^{\prime})=\dim H_{1}(K;\mathbf{Z}_{2})\geq\dim H_{1}(Q^{c};\mathbf{Z}_{2})=g.$
Thus all the inequalities are equalities and the map (2) is an isomorphism.
Now the $g+1$ curves of $\mathscr{G}$ taken together are homologically trivial (mod $2$) in $Q^{c}$ and thus also in $K$. Thus they bound a surface in $K$. That surface must be connected since no subcollection of those curves bounds a surface in $Q^{c}$. Hence $K$ is connected. Therefore $M^{\prime}=\partial K$ is connected, so $\Sigma=M^{\prime}$.
It remains to show that if $W$ is a simply connected open subset of $Q^{c}$, then $M\cap W$ has genus 00. Since $M$ and $\Sigma$ have genus $g$, $M\setminus\Sigma$ has genus 00, and therefore $(M\setminus\Sigma)\cap W$ has genus 00. Thus it suffices to prove that $\Sigma\cap W$ has genus 00. Suppose to the contrary that $\Sigma\cap W$ has positive genus. Then, by elementary topology (see Lemma 16 below), $\Sigma\cap W$ contains a simple closed curve that does not bound a surface in $K$. Therefore it does not bound a surface in $Q^{c}$ since the map (2) is an isomorphism, and therefore it does not bound a surface in $W$. But that contradicts the assumption that $W$ is simply connected. ∎
Lemma 16.
Let $K$ be a closed region in $\mathbf{R}^{3}$ with smooth boundary $\Sigma$. Let $W$ be an open subset of $\mathbf{R}^{3}$ such that $\Sigma\cap W$ has positive genus. Then $\Sigma\cap W$ contains a simple closed curve that is homologically nontrivial in $K$.
Proof.
Let $\alpha_{1}$ and $\alpha_{2}$ be a pair of simple closed curves in $\Sigma\cap W$ that intersect transversely and at exactly one point. We claim that at most one of $\alpha_{1}$ and $\alpha_{2}$ can be homologically trivial in $K$. For otherwise, there would be smooth, compact embedded surfaces $\sigma_{i}$ in $K$ such that $\partial\sigma_{i}=\alpha_{i}$. We can choose the $\sigma_{i}$ so that they intersect transversely and so that they are never tangent to $\Sigma$. Thus $\sigma_{1}\cap\sigma_{2}$ is a finite collection of curves, and so they have an even number of endpoints. But the set of their endpoints is $\alpha_{1}\cap\alpha_{2}$, which has only one point, a contradiction. ∎
Theorem 17.
Suppose $M\in\mathscr{A}$. Then $T_{Q}\leq T_{\textnormal{pos}}$.
Proof.
Let $0<T<\min\{T_{Q},T_{\textnormal{pos}}\}$. Let $(p,T)$ be a spacetime singular point. Since $T<T_{Q}$, $p\in Q^{c}$. Let $W$ be an open ball in $Q^{c}$ centered at $p$. By Theorems 13, 14, and 15, $\operatorname{genus}(M(t)\cap W)=0$ for all regular times $t<T$. Thus $(p,T)$ must be a genus 00 singularity of the flow. Thus $T<T_{\textnormal{pos}}$, since at time $T_{\textnormal{pos}}$, there must be a singularity of positive genus.
We have proved that $T<T_{\textnormal{pos}}$ for every $T<\min\{T_{Q},T_{\textnormal{pos}}\}$. Hence $T_{Q}\leq T_{\textnormal{pos}}$. ∎
Theorem 18.
Suppose that $M\in\mathscr{A}$ and that $T:=T_{Q}(M)<\infty$. Let $p\in M(T)\cap Q$ and let $\Sigma$ be a shrinker at the spacetime point $(p,T)$.
  1. (1)
    If $p\neq 0$, then $\Sigma$ is a cylinder whose axis is the line $\{rp:r\in\mathbf{R}\}$.
  2. (2)
    If $p=0$, then $\Sigma$ is a connected, $G$-invariant, genus-$g$ surface disjoint from $Q$.
Proof in Case (1).
If $(p,T)$ were a regular point of the flow, then for all $t<T$ sufficiently close to $T$, there would be a unique point $p(t)$ in $M(t)$ closest to $p$. But, since $p(t)\notin Q$, the image of $p(t)$ under rotation by $\pi$ about the line $\{sp:s\in\mathbf{R}\}$ is another point in $M(t)$ closest to $p$, a contradiction. Hence $(p,T)$ is a singular spacetime point of the flow.
Let $\sigma$ be the genus of the singular point $(p,T_{Q})$. Let $S$ be the set consisting of the $g+1$ points in $Q$ at distance $|p|$ from 00, and let
$S^{\prime}=\{(p,T_{Q}):p\in S\}.$
Then all of the $(g+1)$ points in $S^{\prime}$ have the same genus, so
$(g+1)\sigma\leq\operatorname{genus}(M)=g$
by Corollary 5. Thus $\sigma=0$, so $\Sigma$ is a sphere or cylinder. If it were not the cylinder with the indicated axis, then $\Sigma$ would intersect the line $\{sp:s\in\mathbf{R}\}$ transversely in a nonempty set. But then $M(t)$ would intersect $Q$ for all $t<T$ sufficiently close to $T$, which is impossible. This completes the proof in Case (1). ∎
Proof in Case (2).
Recall that $\Sigma$ is a smooth, properly embedded surface. If $0\in\Sigma$, then the oriented tangent plane to $\Sigma$ would be invariant under $G$. But there is no such plane. Hence $0\notin\Sigma$.
If $q$ were a point in $Q\cap\Sigma$, then (as we have just shown) $q\neq 0$. The oriented tangent plane to $\Sigma$ at $q$ is invariant under the stabilizer subgroup $G_{q}$, which implies that $\Sigma$ intersects the ray $\mathbf{R}^{+}p$ orthogonally at $p$. But that is impossible since $\Sigma$ is a smooth limit of surfaces disjoint from $Q$.
We have shown that $\Sigma\subset Q^{c}$. Also, it has genus $\leq g$. It is connected since any smooth embedded shrinker in $\mathbf{R}^{3}$ must be connected. Finally, by the convex hull property, 00 is in the convex hull of $\Sigma$. Thus $\Sigma$ has genus $g$ by Theorem 15. ∎
Theorem 19.
Suppose $M_{n}$ are compact, smoothly embedded surfaces in $Q^{c}$ such that $T_{Q}(M_{n})=\infty$. Suppose the $M_{n}$ converge smoothly to a surface $M$ in $\mathscr{A}$. Then
$M(T_{Q})\cap Q=\{0\},$
where $T_{Q}=T_{Q}(M)$.
Proof.
We can assume that each $M_{n}$ is nonfattening. For let $f_{n}:\mathbf{R}^{3}\to\mathbf{R}$ be a smooth function such that $f_{n}^{-1}(0)=M_{n}$, such that 00 is not a critical value of $f$, and such that $f>0$ at all points of $Q$. Note that there are $\epsilon_{n}\uparrow 0$ such that if $\epsilon_{n}<s_{n}<0$, then the surfaces $M_{n}^{\prime}=f^{-1}(s_{n})$ will converge smoothly to $M$. We can choose the $s_{n}$ so that each $M_{n}^{\prime}$ is nonfattening (since $f^{-1}(s)$ fattens for at most countable set of $s$.) Finally, note that $M_{n}^{\prime}(t)$ never intersects $Q$, since $M_{n}^{\prime}$ lies in the region enclosed by $M_{n}$. Thus we can assume $M_{n}$ is nonfattening (by replacing $M_{n}$ by $M_{n}^{\prime}$, if necessary.)
Suppose the theorem is false. Then there exists a point $p\neq 0$ in $M(T_{Q})\cap Q$.
By Theorem 18,
(3) $(T_{Q}-t)^{-1/2}(M(t)-p)$
converges smoothly and with multiplicity $1$ to a cylinder $\Sigma$ whose axis is the line $\mathbf{R}p:=\{sp:s\in\mathbf{R}\}$. Let $A$ be an Angenent torus that lies in the non-simply connected component of $\Sigma$ and that collapses to the origin.
Then, for all $t<T_{Q}$ sufficiently close to $T_{Q}$, $A$ and 00 will lie in different connected components of the surface (3). Thus the Angenent torus
$\tilde{A}:=(T_{Q}-t)A+p$
and the point $p$ will lie in different components of $M(t)$.
Fix such a time $t$ that is regular for $M(\cdot)$ and also for each of the $M_{n}(\cdot)$.
Then $M_{n}(t)$ converges smoothly to $M(t)$, so, for all sufficiently large $n$, $\tilde{A}$ and $p$ lie in different components of $M_{n}(t)$.
But then, by Lemma 12, $p\in M_{n}(\tau)$ for some $\tau>t$ and thus
$T_{Q}(M_{n})<\infty,$
a contradiction. ∎
Theorem 20.
Suppose $\mathscr{F}$ is a family of compact, connected, smoothly embedded, $G$-invariant surfaces in $\mathbf{R}^{3}\setminus Q$. Suppose also that $\mathscr{F}$ is connected with respect to smooth convergence, and that all the $M\in\mathscr{F}$ lie in some large ball $B(0,R)$. Let $\mathscr{S}$ be the subfamily of $M\in\mathscr{F}$ for which $T_{Q}(M)<\infty$. If $\mathscr{S}$ and $\mathscr{S}^{c}:=\mathscr{F}\setminus\mathscr{S}$ are both nonempty, then there is an $M\in\mathscr{S}$ for which
$M(T_{Q})\cap Q=\{0\}$
(where $T_{Q}=T_{Q}(M)$). If $\Sigma$ is a shrinker at the origin at time $T_{Q}$ to the flow $M(\cdot)$, then $\Sigma$ has genus $g$ and is disjoint from $Q$.
Proof.
Under mean curvature flow, the sphere $\partial B(0,R)$ shrinks to a point in time $(R/2)^{1/2}$. Thus, if $M\in\mathscr{F}$, then $M(t)$ is empty for all $t>(R/2)^{1/2}$. We claim that $\mathscr{S}$ is a closed subset of $\mathscr{F}$. To see that, suppose that $M_{n}\in\mathscr{S}$ converges to $M\in\mathscr{F}$. By hypothesis, there exist $t_{n}$ and $p_{n}$ such that
$p_{n}\in Q\cap M_{n}(t_{n}).$
Since $p_{n}\in B(0,R)$ and since $t_{n}\leq(R/2)^{1/2}$, we can assume that $p_{n}$ and $t_{n}$ converge so to finite limits $p$ and $t$. Thus
$p\in Q\cap M(t),$
so $M\in\mathscr{S}$. Thus $\mathscr{S}$ is a closed subset of $\mathscr{F}$.
By connectedness, $\mathscr{S}^{c}$ cannot be closed: there must be $M_{n}\in\mathscr{S}^{c}$ converging to an $M\in\mathscr{S}$. By Theorem 19, $M(T_{Q})\cap Q=\{0\}$. By Theorem 18, any shrinker at $(0,T_{Q})$ has genus $g$ and is disjoint from $Q$. ∎
Theorem 21.
For each $Q$ as in Theorem 11, there exists a family $\mathscr{F}$ satisfying the hypotheses of Theorem 20.
Proof.
[…]
  1. (1)
    $f^{-1}(0)=E$,
  2. (2)
    $f^{-1}(\infty)=C$,
  3. (3)
[…]
  4. (4)
    For each $s\in(0,1)$, each component of $\{f\geq s\}$ is geodesically convex.
Let
$h:(0,\infty)\to[0,\infty)$
be a smooth function such that be a smooth function such that
  1. (1)
    $h(1)=0$.
  2. (2)
    $h^{\prime}<0$ on $(0,1)$.
  3. (3)
    $h^{\prime}>0$ on $(1,\infty)$.
  4. (4)
    $h(r)\to\infty$ as $r\to 0$ or as $r\to\infty$.
For example, we could let
$h(r)=(\log r)^{2}.$
Extend $h$ to the interval to $[0,\infty)$ by setting $h(0)=\infty$.
Now define
$\Phi:\{p:|p|\in[1/2,2]\}\to[0,\infty]$
by
$\Phi(p)=f(p/|p|)+h(|p|).$
Now for $s\in[0,\infty)$, let $M^{s}$ be the set of $p$ such that $p\in[1/2,2]$ and such that
$\Phi(p)=s.$
Now $M^{0}=E$, which vanishes instantly under level set flow, so $T_{Q}(M^{0})=\infty$ and therefore $T_{Q}(M^{s})=\infty$ for all sufficiently small $s$.
Note that
$\{\Phi<\infty\}=\{p:1/2<|p|<1\}\setminus Q.$
Thus, for all sufficiently large $s$, the region $\Phi<s$ enclosed by $M^{s}$ contains an Angenent torus centered at a point $p\in C\subset Q$. Thus $T_{Q}(M^{s})<\infty$ by Lemma 12. Now fix a $c$ such that $T_{Q}(M_{c})<\infty$. Then the family
=$\mathscr{F}=\{M^{s}:0<s\leq c\}$
satisfies the hypotheses of Theorem 20. This completes the proof in the case of Platonic solids.
Now consider the when $Q$ consists of the rays
$\{(r\cos\theta,r\sin\theta,0):r>0\}$
such that $\theta$ is multiple of $2\pi/(g+1)$. Thus
$Q=\{(r\cos\theta,r\sin\theta,0):r\geq 0,\,\sin\left(\frac{g+1}{2}\theta\right)=0\}.$
[…]
$f(r\cos\theta,r\sin\theta,z)=u\left(\cos\left(\frac{g+1}{2}\theta\right)^{2}\right),$
where $u(s)=\frac{s}{1-s}$.
The rest of the proof is exactly as in the Platonic solid case. ∎
5. A Topological Lemma
In §4, we produced shrinkers that are disjoint from certain collections of rays from the origin. In the remainder of the paper, we will produce shrinkers that contain certain collections of rays from the origin. This section proves basic facts about the topology of such surfaces.
Let $g\geq 1$ be an integer. Let $Q=Q_{g}$ be the union of the lines
$\{(r\cos\theta,r\sin\theta,0):r\in\mathbf{R}\},$
such that $\theta$ is a multiple of $\pi/(g+1)$. For $\theta\in\mathbf{R}$, let $P_{\theta}$ be the vertical plane spanned by $(\cos\theta,\sin\theta,0)$ and $\mathbf{e}_{3}$. Let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ consisting of:
  1. (1)
    Reflections in the planes $P_{\theta}$ for which $\theta$ is an odd multiple of $\pi/(2(g+1))$,
  2. (2)
    Rotations about $Z$ by multiples of $\pi/(g+1)$.
  3. (3)
    Rotation by $\pi$ about each of lines in $Q_{g}$.
In §7, we will produce (for each $g\geq 1$) a $G_{g}$-invariant shrinker $\Sigma$ of genus $g$ such that
(4) $\Sigma\cap\{z=0\}=Q_{g}.$
In §8, we will produce (for all sufficiently large $g$) a $G_{g}$-invariant shrinker $\Sigma$ of genus $2g$ such that (4) holds. In this section, we prove some theorems that apply in both of those cases.
Lemma 22 (Topology Lemma).
Let $U$ be a $G$-invariant region in $\mathbf{R}^{3}$ such that $\partial U$ is smooth and such that $U$ is diffeomorphic to an open ball. Suppose that $M\subset\overline{U}$ is a smoothly embedded, $G$-invariant, compact, connected $2$-manifold-with-boundary such that
$\partial M=M\cap\partial U,$
$M\cap\{z=0\}\cap U=Q\cap U,\,\text{and}$
$\operatorname{genus}(M)\leq 2g.$
Then $M$ has genus 00, $g$, or $2g$.
Let $c$ be the number of boundary components $C$ of $\partial M$ such that $C$ lies in $\{z>0\}$ and such that $C$ winds once around $Z$. Let $d$ be the number of points in
$M\cap Z^{+}=M\cap Z\cap\{z>0\}.$
Then
$\operatorname{genus}(M)=(c+d)g.$
For surfaces of genus $\leq g$, a result similar to Lemma 22 was proved, using similar methods, by Buzano, Nguyen, and Schulz. See Lemma 2.10 and Corollary 2.11 of [buzano].
Proof of Lemma 22.
Let $W$ be the wedge given in cylindrical coordinates by
$\frac{-\pi}{2(g+1)}<\theta<\frac{\pi}{2(g+1)}.$
Note that
$\{x:(x,0,0)\in M\}$
is a compact interval $[-R,R]$.
Consider the component $J$ of
$(\partial M)\cap W=M\cap\partial U\cap W$
that contains the point $(R,0,0)$. Because $J$ is an embedded curve in $\partial U$ and is invariant under rotation by $\pi$ about the $x$-axis, it cannot be a closed curve. Thus, $J$ is a non-closed curve with $\theta=\pi/(g+1)$ on one endpoint and $\theta=-\pi/(g+1)$ on the other endpoint. It follows from the symmetries of $M$ that the component $J^{\prime}$ of $\partial M$ containing $J$ is a simple closed curve that winds once around the $z$-axis and that passes through all the points of $Q\cap\partial U$.
Attach disks symmetrically to each of the other connected components of $\partial M$ to get a surface $M^{\prime}$ with only one boundary component, namely $J^{\prime}$. Note that
(5) $\operatorname{genus}(M^{\prime})=\operatorname{genus}(M).$
Let $k$ be the number of points in $M^{\prime}\cap Z^{+}$. Note that
(6) $k=c+d.$
By symmetry, the number of points of $M^{\prime}\cap Z$ is $12k+$ 1. Let $M^{\prime\prime}$ be the quotient of $M^{\prime}$ under the group consisting of rotations about $Z$ by angles that are multiples of $2\pi/(g+1)$. Then
$\chi(M^{\prime}\setminus Z)=(g+1)\chi(M^{\prime\prime}\setminus Z)$
so
(7) $\chi(M^{\prime})-(2k+1)=(g+1)(\chi(M^{\prime\prime})-(2k+1)).$
Now $M^{\prime}$ is connected and has exactly one boundary component (namely $J^{\prime}$), and likewise for $M^{\prime\prime}$, so
$\chi(M^{\prime})$ $=1-2\operatorname{genus}(M^{\prime}),$
$\chi(M^{\prime\prime})$ $=1-2\operatorname{genus}(M^{\prime\prime}).$
Thus, multiplying (7) by $-1$ gives
$2\operatorname{genus}(M^{\prime})+2k=(g+1)(2\operatorname{genus}(M^{\prime\prime})+2k),$
so
(8) $\operatorname{genus}(M^{\prime})=gk+(g+1)\operatorname{genus}(M^{\prime\prime}).$
The image of $Q\cap U$ under the branched covering map
$\pi:M^{\prime}\to M^{\prime\prime}$
is an embedded curve $\Gamma$ whose endpoints are both on $\partial M^{\prime\prime}$. Now $\Gamma$ divides $M^{\prime\prime}$ into two congruent surfaces
$(M^{\prime\prime})^{+}=M^{\prime\prime}\cap\{z>0\}$
and
$(M^{\prime\prime})^{-}=M^{\prime\prime}\cap\{z<0\}.$
Thus
$\operatorname{genus}(M^{\prime\prime})$ $=\operatorname{genus}((M^{\prime\prime})^{+})+\operatorname{genus}((M^{\prime\prime})^{-})$
$=2\operatorname{genus}((M^{\prime\prime})^{+}).$
Therefore, by (8),
$\operatorname{genus}(M^{\prime})=gk+2(g+1)\operatorname{genus}((M^{\prime\prime})^{+}).$
Since $\operatorname{genus}(M^{\prime})\leq 2g$, we see that $(M^{\prime\prime})^{+}$ has genus 00, that $k$ is 00, $1$, or $2$, and thus that
$\operatorname{genus}(M)$ $=\operatorname{genus}(M^{\prime})$
$=gk$
$=g(c+d).$
by (5) and (6). ∎
Corollary 23.
Suppose $M$ is a $G_{g}$-invariant embedded shrinker in $\mathbf{R}^{3}$ such that
$M\cap\{z=0\}=Q_{g}$
and such that $\operatorname{genus}(M)\leq 2g$. Then $M$ has genus $g$ or $2g$.
Proof.
For $R>0$, let $M_{R}$ be the connected component of $M\cap B(0,R)$ that contains the origin. Since $M$ is a translator, it is connected. Thus
$\operatorname{genus}(M_{R})=\operatorname{genus}(M)$
for all sufficiently large $R$. Choose such an $R$ for which $\partial B(0,R)$ is transverse to $M$. Now apply Lemma 22 to $U=B(0,R)$ and to $\overline{M_{R}}$. ∎
Corollary 24.
Let $M$ be as in Corollary 23. Suppose also that $M\setminus\{z=0\}$ has genus 00.
  1. (1)
    If $M$ has genus $g$, then there is an $r\in(0,\infty)$ such that
[…]
  2. (2)
    If $M$ has genus $2g$ and if
    $\operatorname{genus}(M\setminus\{z=0\})=0,$
    then there exist $r$ and $R$ with $0<r\leq R<\infty$ such that
[…]
Proof.
Let $\Gamma(r)=\operatorname{genus}(M\cap B(0,r))$. Trivially, $\Gamma$ is an increasing function of $r$, $\Gamma(r)$ is 00 for all sufficiently small $r$ and Let $M^{\prime}_{r}$ be the connected component of $M\cap B(0,r)$ containing the origin. The portion of $M\cap B(0,R)$ not in $M^{\prime}_{r}$ lies in
$M\setminus\{z=0\},$
and therefore has genus 00, so
(9) $\Gamma(r)=\operatorname{genus}(M_{r}^{\prime}).$
Now suppose that $M$ has genus $g$. By Lemma 22, The right hand side of (9) is either 00 or $g$. Thus there is an $r$ such that
$\operatorname{genus}(M\cap B(0,\rho))=\begin{cases}0&(\rho\leq r),\\ g&(r<\rho).\end{cases}.$
Thus if $\rho>r$,
$\operatorname{genus}(M\setminus\overline{B(0,\rho)}\leq\operatorname{genus}(M)-\operatorname{genus}(M\cap B(0,\rho)=g-g=0.$
Letting $\rho\to r$ gives $\operatorname{genus}(M\setminus\overline{B(0,r)})=0$. Thus
[…]
$=0$
This completes the proof when $M$ has genus $g$. The case when $M$ has genus $2g$ is essentially the same; the function $\Gamma(\cdot)$ only takes the values $g$ and $2g$. We can let $r$ be the largest $r$ for which $\Gamma(r)<g$ and $R$ be the largest $R$ for which $\Gamma(r)<2g$. ∎
[…]
We add a line at infinity as follows. Let $\mathbf{R}^{2}\cup\{\infty\}$ be the conformal compactification of $\mathbf{R}^{2}$, and let
$N=(\mathbf{R}^{2}\cup\infty)\times\mathbf{R}.$
Thus $N$ is $\mathbf{R}^{3}$ with a vertical line $Z_{\infty}$ at infinity attached:
$Z_{\infty}=\{\infty\}\times\mathbf{R}.$
We let $O$ be the origin, and we let
$O_{\infty}=(\infty,0)=Z_{\infty}\cap\{z=0\}.$
For $r>0$, let $\eta_{r}$ be the metric on $\mathbf{R}^{2}\cup\{\infty\}$ obtained by stereographic projection from the north pole of a sphere of radius $r$ to the tangent plane at the south pole. In this metric, the circle
$(x^{2}+y^{2})^{1/2}=2r$
[…]
Except where otherwise specified, $N$ will refer the manifold $N$ with the Riemannian metric $\gamma:=\gamma_{1}$.
Note that the circle
(10) $C:=\{(x,y,0):(x^{2}+y^{2})^{1/2}=2\}$
and the completed line
$\{(r\cos\theta,r\sin\theta,0):r\in\mathbf{R}\}\cup\{O_{\infty}\}$
are totally geodesic great circles with respect to $\gamma$. Thus the cylinder
(11) $\operatorname{Cyl}:=\{(x,y,z):(x^{2}+y^{2})^{1/2}=2\}$
is also totally geodesic in $N$.
If $g$ is a positive integer, we let $Q=Q_{g}$ be set of horizontal lines and $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ that were defined at the beginning of Section 5. Note that each $\sigma\in G$ extends to an isometry of $N:=\mathbf{R}^{3}\cup Z_{\infty}$.
Definition 25.
If $\Gamma$ is a great circle in $\{z=0\}$, let $\rho_{\Gamma}:N\to N$ denote rotation by $\pi$ about $\Gamma$. Let $\widetilde{G}=\widetilde{G}_{g}$ denote the group of isometries of $N$ generated by $G$ together with $\rho_{C}$. Let
$\widetilde{Q}$ $=\widetilde{Q}_{g}$
$=\overline{Q}\cup C$
$=Q\cup\{O_{\infty}\}\cup C.$
Note that $\sigma(\widetilde{Q}_{g})=\widetilde{Q}_{g}$ for each $\sigma\in\widetilde{G}_{g}$.
The theorems in this section and in the following two sections will all be about the following class of surfaces:
Definition 26.
We let $\mathscr{M}_{g}$ be the collection of surfaces $M$ with the following properties:
  1. (1)
    $M$ is a compact, smoothly embedded, $\widetilde{G}_{g}$-invariant surface in $N$.
  2. (2)
    $M$ intersects the cylinder $\operatorname{Cyl}$ (see (11)) transversely, and the intersection is the circle $C$ (see (10)).
  3. (3)
    $M\cap\{z=0\}=\widetilde{Q}_{g}$.
  4. (4)
    $M\setminus\{z=0\}$ has genus 00.
  5. (5)
    $M$ has genus $\leq 4g$.
  6. (6)
    $\operatorname{area}_{\,\gamma}(M)<2\operatorname{area}_{\,\gamma}(\{z=0\})=8\pi$.
(Recall that $\{z=0\}$ is isometric to the sphere of radius $1$.)
Most of this section is dedicated to establishing that if $M\in\mathscr{M}_{g}$, then the evolving surface $M(t)$ remains in $\mathscr{M}_{g}$ for all regular times $t$ in the interval $[0,T_{\textnormal{pos}}(M)]$. See Theorem 31. We also analyze the behavior of $M(t)$ as $t$ approaches $T_{\textnormal{pos}}(M)$. Specifically, we demonstrate that if $T_{\textnormal{pos}}(M)$ is finite, then the origin is a singular point at time $T_{\textnormal{pos}}(M)$. Additionally, we show that each shrinker $\Sigma$ at $(O,T_{\textnormal{pos}})$ has genus $g$ or $2g$, and that its intersection with $\{z=0\}$ is $Q$. See Lemma 34 and Theorem 35.
Proposition 27.
Suppose that $M$ is a surface in $\mathscr{M}_{g}$ and let $M^{\prime}$ be the connected component of $M$ containing $\widetilde{Q}$. Then
$\operatorname{genus}(M)=\operatorname{genus}(M^{\prime})=2dg,$
where $d$ is the number of points in $M^{\prime}\cap Z^{+}$.
Proof.
Note that
$\operatorname{genus}(M)=\operatorname{genus}(M^{\prime})$
since
$\operatorname{genus}(M\setminus M^{\prime})\leq\operatorname{genus}(M\setminus\widetilde{Q})=\operatorname{genus}(M\setminus\{z=0\})=0.$
Furthermore,
$\operatorname{genus}(M^{\prime})=2\operatorname{genus}(M^{\prime}\cap U),$
where $U=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}.$ Now $M^{\prime}\cap\overline{U}$ has exactly one boundary component, namely the circle $C$. Thus, by Lemma 22,
$\operatorname{genus}(M^{\prime}\cap U)=dg.$
so $M^{\prime}$ has genus $2dg$. ∎
For $\Lambda\geq 0$, let $\operatorname{Cone}(\Lambda)$ be the set of points $p$ in
(12) $\{(x^{2}+y^{2})^{1/2}\leq 2,\,z\geq 0\}$
such that
(13) $|z|\leq\Lambda\operatorname{dist}(\pi(p),C),$
where $\pi$ is the projection $\pi(x,y,z)=(x,y,0)$ and where $\operatorname{dist}$ is geodesic distance with repect to the metric $\gamma$. We let $\operatorname{Cone}^{\prime}(\Lambda)$ be the set of all $p$ satisfying (13) Thus $\operatorname{Cone}(\Lambda)$ is the portion of $\operatorname{Cone}^{\prime}(\Lambda)$ in the set (12). Conversely, $\operatorname{Cone}^{\prime}(\Lambda)$ consists of $\operatorname{Cone}(\Lambda)$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$.
Note that if $\Lambda$ is sufficiently large, then $M\subset\operatorname{Cone}^{\prime}(\Lambda)$.
Lemma 28.
For $t\in[0,\operatorname{T_{\textnormal{fat}}}]$,
$M(t)\cap\{z=0\}=\widetilde{Q}.$
Also, if $M\in\operatorname{Cone}^{\prime}(\Lambda)$, then $M(t)\in\operatorname{Cone}^{\prime}(\Lambda)$ for all $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
Proof.
Let $\Sigma_{0}$ be the closure of
$M\cap\{(x^{2}+y^{2})^{1/2}<2\}\cap\{z>0\},$
Note that
(14) $\Gamma$ $:=\partial\,\Sigma_{0}$
$=\Sigma_{0}\cap\{z=0\}$
$=(Q\cap\{(x^{2}+y^{2})^{1/2}<2\})\cup S,$
where $S$ consists of alternate components of $C\setminus Q$ (see Fig. 4.)
Figure 4. The curve $\Gamma$ in the case $g=3$.
We prove the lemma by elliptic regularization. We work in $N\times\mathbf{R}$. We let $\widetilde{G}$ (and therefore $\widetilde{G}^{+}$) act on $N\times\mathbf{R}$ by acting on the first factor.
Let $w:N\times\mathbf{R}\to\mathbf{R}$ be the projection onto the second factor. For $\lambda\geq 0$, let $g_{\lambda}$ be $e^{-(2/3)\lambda w}$ times the product metric on $N\times\mathbf{R}$.
Let $A_{\lambda}$ be a $3$-dimensional surface (flat chain mod $2$) that minimizes $g_{\lambda}$-area among surfaces in
$\operatorname{Cone}(\Lambda)\times[0,\infty)$
whose boundary consists of the surface $\Sigma_{0}\times\{0\}$ together with the surface $\Gamma\times[0,\infty)$.
Standard cut-and-paste arguments (as in [morgan]*Lemma 7.3) show that $A_{\lambda}$ is $G^{+}$-invariant. Alternatively, if one does not wish to invoke those arguments, one can simply let $A_{\lambda}$ minimize $g_{\lambda}$-area in the class of $G^{+}$-invariant surfaces.
We remark that, by standard GMT regularity, $A_{\lambda}$ is smooth except possibly at those points where its boundary is not smooth. However, we do not need those regularity results here.
Now let $A^{\prime}_{\lambda}$ be the surface consisting of $A_{\lambda}$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$. Thus $A_{\lambda}^{\prime}$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)\times\mathbf{R}$. Note that the boundary of $A^{\prime}_{\lambda}$ is $M\times\{0\}$, and that the integral varifold associated to $A^{\prime}_{\lambda}$ is stationary in the set $\{w>0\}$ with respect to the metric $g_{\lambda}$.
It follows that
(15) $t\in(0,\infty)\mapsto A_{\lambda}(t):=A_{\lambda}-t\lambda\mathbf{e}_{4},$
$t\in(0,\infty)\mapsto A^{\prime}_{\lambda}(t):=A^{\prime}_{\lambda}-t\lambda\mathbf{e}_{4}$
are mean curvature flows (with moving boundaries).
By the theory of elliptic regularization, there is a sequence of $\lambda$ tending to infinity for which the flows (15) converge to integral Brakke flows
(16) $t\in(0,\infty)\mapsto A(t),$
$t\in(0,\infty)\mapsto A^{\prime}(t)$
Furthermore, these flows have a special form: there exist mean curvature flows
$t\in[0,\infty)\mapsto\Sigma(t),$
$t\in[0,\infty)\mapsto\Sigma^{\prime}(t).$
in $N$ with the following properties:
  1. (1)
    $A(t)=\Sigma(t)\times\mathbf{R}$ and $A^{\prime}(t)=\Sigma^{\prime}(t)\times\mathbf{R}$ for all $t>0$.
  2. (2)
    $\Sigma(0)=\Sigma_{0}$ and $\Sigma^{\prime}(0)=M$.
  3. (3)
    $\Sigma(\cdot)$ is an mean curvature flow with (fixed) boundary, the boundary being $\Gamma=\partial\Sigma_{0}$.
  4. (4)
    $\Sigma^{\prime}(\cdot)$ is an integral Brakke flow (without boundary).
By construction,
$\Sigma(t)\subset\operatorname{Cone}(\Lambda)\quad(t\geq 0)$
and
(17) $\Sigma^{\prime}(t)\subset\operatorname{Cone}^{\prime}(\Lambda)\quad(t\geq 0).$
Let $U$ be a connected component of $\{z=0\}\setminus\Gamma$. If $\Sigma(t)$ contained a point of $U$, then, by the strong maximum principle, $\Sigma(\tau)$ would contain $U$ for all $\tau\in[0,t]$. But $\Sigma(0)=\Sigma_{0}$ is disjoint from $U$. Thus
$\Sigma(t)\cap U=\emptyset\quad(t\geq 0).$
Since $U$ can be any component of $\{z=0\}\setminus\Gamma$,
$\Sigma(t)\cap\{z=0\}=\Gamma.$
for all $t\geq 0$. Now $\Sigma^{\prime}(t)$ consists of $\Sigma$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$. Thus
(18) $\Sigma^{\prime}(t)\cap\{z=0\}=\widetilde{Q}\qquad(t\geq 0).$
Now the flow $\Sigma^{\prime}(\cdot)$ is unit-regular and cyclic mod $2$. (This is true for any flow constructed by elliptic regularization. The unit-regularity follows from the fact that the flows (15) are unit-regular, which is an immediate consequence of Allard’s Regularity Theorem.)
Thus
(19) $\Sigma^{\prime}(t)=M(t)\quad\text{for all $t\in[0,\operatorname{T_{\textnormal{fat}}}(M))$}.$
(See [cms]*Lemma 9.3.)
The assertions of the lemma now follow immediately from (17), (18), and (19). ∎
Lemma 29.
The points in $C$ are regular points of the flow at all times $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
Proof.
Let $\Sigma$ be a shrinker to the flow at a spacetime point $(p,t)$, where $p\in C$. By Lemma 28, $\Sigma$ (after a rotation) is contained in the set
$\{(x,y,z):|z|\leq|x|\}.$
By [white-boundary]*Theorem 15.1, $\Sigma$ is a union of halfplanes bounded by the $y$-axis. But $\Sigma$ is smooth with multiplicity $1$, so it is a single multiplicty $1$ plane, and therefore $(p,T)$ is a regular point. ∎
Lemma 30.
Suppose $M\in\mathscr{M}_{g}$. If $t<T_{\textnormal{pos}}(M)$ is a regular time, then
$\operatorname{genus}(M(t))\leq 4g,$
$\operatorname{genus}(M(t)\setminus\widetilde{Q})=0.$
for all regular times $t<T_{\textnormal{pos}}(M)$.
Proof.
This is a special case of the following theorem (see [white-genus]): if $M$ is a smoothly embedded, compact surface in a complete $3$-manifold of bounded Ricci curvature and if $K$ is a closed set such that $K\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}}]$, then for every regular time $t<T_{\textnormal{pos}}$,
$\operatorname{genus}(M(t)\setminus K)\leq\operatorname{genus}(M\setminus K).$
Theorem 31.
Suppose $M\in\mathscr{M}_{g}$. Then $M(t)\in\mathscr{M}_{g}$ for every regular time $t<T_{\textnormal{pos}}(M)$.
Proof.
This is an immediate consequence of Lemmas 2830, and the fact that $\operatorname{area}(M(t))$ is a decreasing function of $t$. ∎
Recall that $\{z=0\}$ is isometric to a sphere of radius $1$, so it has area $4\pi$.
Theorem 32.
If $\operatorname{T_{\textnormal{fat}}}=\infty$, then $M(t)$ converges smoothly to $\{z=0\}$ with multiplicity $1$. In particular, there is a time $T<\infty$ such that $M(t)$ is a smooth graph (i.e., such that $M(t)$ projects diffeomorphically onto $\{z=0\}$) for all $t\geq T$.
Proof.
Let $t_{n}\to\infty$. After passing to a subsequence, the flows
$t\in[-t_{n},\infty)\mapsto M_{n}(t):=M(t+t_{n})$
converge to a static integral Brakke flow. That is, there is a stationary integral varifold $V$ such that for every $t$, $M_{n}(t)$ converges (as Radon measures) to the Radon measure associated to $V$.
Now $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$, so, by the maximum principle, $V$ is contained in $\{z=0\}$. Thus $V$ is $\{z=0\}$ with some integer multiplicity $k$. Since $M_{n}(t)$ contains $\widetilde{Q}$ for all $t$, we see that $\widetilde{Q}$ is in the support of $V$, and thus that
$k\geq 1.$
Also,
$k\operatorname{area}(\{z=0\})=\operatorname{area}(V)\leq\operatorname{area}(M)<2\operatorname{area}(\{z=0\}).$
Thus $k=1$.
We have shown that the flows $M_{n}(\cdot)$ converge to the constant flow given by $\{z=0\}$ with multiplicity $1$. By the Local Regularity Theorem (Theorem 7), the convergence is smooth with multiplicity $1$. ∎
Theorem 33.
Let $p$ be a point in $\widetilde{Q}\setminus\{O,O_{\infty}\}$. Then for all $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, $(p,T)$ is a regular point of the flow $M(\cdot)$.
Proof.
[…]
$4(g+1)\tilde{g}\leq\operatorname{genus}(M)\leq 4g,$
so $\tilde{g}=0$. Thus if $(p,T)$ were a singular point, the shrinker would be sphere or a cylinder, which is impossible since $p\in M(t)$ for all $t\in[0,T]$. Thus $(p,T)$ is a regular point. ∎
Lemma 34.
The origin is a regular point of the flow $M(\cdot)$ at all times $<T_{\textnormal{pos}}$. If $T_{\textnormal{pos}}<\infty$, then the origin is singular at time $T_{\textnormal{pos}}$, and if $\Sigma$ is a shrinker at $(O,T_{\textnormal{pos}})$, then $\Sigma$ has genus $>0$.
Proof.
Let $0<T<T_{\textnormal{pos}}$ and let $\Sigma$ be a shrinker at the spacetime point $(0,T)$. Then $\Sigma$ is a plane, a sphere, or a cylinder. Since $\Sigma$ contains $Q$, it must be a plane, and thus $(O,T)$ is a regular point.
Now suppose $T=T_{\textnormal{pos}}<\infty$. Then there is a spacetime point $(p,T)$ of positive genus. Now $p$ must lie in $\{z=0\}$, since
$\operatorname{genus}(M(t)\setminus\{z=0\})=0$
for all regular times $t\in[0,T_{\textnormal{pos}}]$. Thus $p\in\widetilde{Q}$. By Theorem 33, $p$ is either the origin or the point $O_{\infty}$. By symmetry, we can assume it is the origin. ∎
Theorem 35.
Suppose that $T_{\textnormal{pos}}<\infty$ and that $\Sigma$ is a shrinker at the origin time $T_{\textnormal{pos}}$. Then $\Sigma\cap\{z=0\}=Q$. Furthermore,
$\operatorname{genus}(\Sigma)\leq\operatorname{genus}(M)/2,$
and $\Sigma$ has genus $g$ or genus $2g$.
Proof.
Let $t_{i}\uparrow T=T_{\textnormal{pos}}$ be such that
$S_{i}:=(T-t_{i})^{-1/2}M(t_{i})$
converges smoothly to $\Sigma$. Since $\widetilde{Q}\subset M(t)$ for each $t$, $Q\subset\Sigma$. Since $(0,T)$ is not a regular point (by Lemma 34), $\Sigma$ is not a plane. Thus, as $\Sigma$ and $\{z=0\}$ are both minimal with respect to the shrinker metric, $\Sigma\cap\{z=0\}$ is a network of curves, the vertices of which are the points of tangency. Hence if $\Sigma\cap\{z=0\}$ contained a point not in $Q$, it would also contain such point $q$ where $\Sigma$ and $\{z=0\}$ intersect transversely. But then $q$ would be a limit of points where $S_{i}$ and $\{z=0\}$ intersect, which is impossible since $M(t_{i})\cap\{z=0\}$ is $\widetilde{Q}$.
Because the singularities at $O$ and $O_{\infty}$ at time $T$ have the same genus, namely, $\operatorname{genus}(\Sigma)$,
$2\operatorname{genus}(\Sigma)\leq\operatorname{genus}(M).$
Finally, $\Sigma$ has genus $g$ or $2g$ by Lemma 22. ∎
7. Shrinkers with one end
[…]
$2>d_{1}>d_{2}.$
Figure 5. A numerical simulation of the shrinker $\Sigma_{5}$ in Theorem 36. Figure courtesy of Mario Schulz.
Theorem 36.
[…]
  1. (1)
    $\Sigma_{g}$ is $G_{g}$-invariant and has genus $g$.
  2. (2)
    $\Sigma_{g}\cap\{z=0\}=Q_{g}$.
  3. (3)
    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
  4. (4)
    $\Sigma_{g}$ has entropy $\leq\delta$.
Furthermore, for large $g$, $\Sigma_{g}$ has exactly one end, and, as $g\to\infty$, $\Sigma_{g}$ converges to the union of the plane $\{z=0\}$ and the sphere $\partial B(0,2)$. The convergence is smooth with multiplicity $1$ away from the intersection $\{z=0\}\cap\partial B(0,2)$.
See Figure 5.
Existence of shrinkers with properties (1)–(4) was proved (by surgery), for all sufficiently large $g$, by Kapouleas, Kleene, and Moeller [kkm], and, independently, by X. H. Nguyen [nguyen]. It was subsquently proved for all $g$ by Buzano, Huy The Nguyen, and Schulz using minimax [buzano]. However, none of those works proved that any of the shrinkers arose as blowups of the mean curvature flow of a compact, smoothly embedded initial surface.
Figure 6. The piecewise smooth surface
$\Delta_{R}\cup(\{z>0\}\cap\partial B(0,R))$
in the case $g=1$. Its boundary consists of a pair of diameters, and a pair of circular arcs.
Proof of Theorem 36.
As in §6, we work in
[…]
we let
$\widetilde{Q}=Q\cup C\cup\{O_{\infty}\},$
and we let $\widetilde{G}=\widetilde{G}_{g}$ be the group of isometries defined in §6. Recall that $C$ is the circle
$C=\{(x,y,0):(x^{2}+y^{2})^{1/2}=2\}.$
For $20\leq R<$ 2, let
$\Delta_{R}$ =$=\{(r\cos\theta,r\sin\theta,0):\text{$0\leq r\leq R$ and $\sin((g+1)\theta)\geq 0$}\}$
$\qquad\cup\{(r\cos\theta,r\sin\theta,0):\text{$R\leq r\leq 2$ and $\sin((g+1)\theta)\leq 0$}\}.$
See Figure 6. Note that $\Delta_{0}$ is the union of $g+1$ wedge-shaped regions in the unit disk of the $xy$-plane. Thus $\partial\Delta_{0}$ consists of $g+1$ diameters of the unit disk together with $g+1$ arcs in the unit circle. Note that the surface
(*) $\Delta_{R}\cup(\{z>0\}\cap\partial B(0,R))$
[…]
Now reflect $S^{+}$ about the $x$-axis to get $S^{-}$. Then $S:=S^{+}\cup S^{-}$ will be a $2$-manifold whose boundary is the circle $C$.
Now let
$M=S\cup\rho_{C}S.$
Note that we can choose $S^{+}$ so that $M$ is smooth and $\widetilde{G}$-invariant.
Note that $S$ can be obtained by taking the union of the disk $D$ and the sphere $\partial B(0,R)$ and then doing surgery along the circle of intersection. The Euler characteristic of the disjoint union is $1+2=3$. The surgery does not change the number of edges (in a suitable triangulation); it merely changes the way that faces are attached to edges. But the surgery does reduce the number of vertices by $2(g+1)$, since $2(g+1)$ vertices on the equator of the sphere get identified with the corresponding points in $D$. Thus
$\chi(S)=3-2(g+1)=1-2g.$
On the other hand, since $S$ is connected and has one boundary component, $\chi(S)=1-2\operatorname{genus}(S)$. Thus $\operatorname{genus}(S)=g$, and so (since the simple closed curve $C$ disconnects $M$),
$\operatorname{genus}(M)$ $=\operatorname{genus}(S)\cup\operatorname{genus}(\rho_{C}S)$
$=2\operatorname{genus}(S)$
$=2g.$
[…]
$\operatorname{area}(M)<2\operatorname{area}(\{z=0\}).$
To summarize, the surface $M$ we constructed has the following properties:
  1. (1)
    $M$ is compact, connected, smoothly embedded, and $\widetilde{G}_{g}$ invariant.
  2. (2)
    $M\cap\{z=0\}=\widetilde{Q}_{g}$.
  3. (3)
    $M$ has genus $2g$.
  4. (4)
    $M\setminus\{z=0\}$ has genus 00.
  5. (5)
    $M$ intersects $\{x^{2}+y^{2}=2\}$ transversely, and the intersection is $C$.
  6. (6)
    $\operatorname{area}(M)<2\operatorname{area}(\{z=0\})$.
We let $\mathscr{M}_{g}^{\prime}$ be the class of surfaces having properties (1)–(6). Thus $\mathscr{M}_{g}^{\prime}$ is the set of $M\in\mathscr{M}_{g}$ such that $M$ is connected and has genus $2g$, where $\mathscr{M}_{g}$ is the class of surfaces specified in Definition 26. In particular, all the theorems in §6 apply to the surfaces in $\mathscr{M}_{g}^{\prime}$.
Now let $M$ be any surface in $\mathscr{M}_{g}^{\prime}$. We orient $M$ so that the unit normal at $(2,0,0)$ is $\mathbf{e}_{3}$. It follows that the unit normal at the origin is $-\mathbf{e}_{3}$:
(20) $\nu_{M}(2,0,0)$ $=\mathbf{e}_{3},$
$\nu_{M}(0,0,0)$ $=-\mathbf{e}_{3}.$
Let $M(t)$ denote the result of letting $M$ flow for time $t$ by the level set flow. Let $T_{\textnormal{pos}}=T_{\textnormal{pos}}(M)$ be the first time that there is singularity of positive genus; if there is no such time, let $T_{\textnormal{pos}}=\infty$.
Claim 1.
For all $t\in[0,T_{\textnormal{pos}})$,
(21) $\nu_{M(t)}(2,0,0)$ $=\mathbf{e}_{3},$
$\nu_{M(t)}(0,0,0)$ $=-\mathbf{e}_{3}.$
Proof.
By Theorems 33 and 34, the flow is regular at all points of $\tilde{Q}$ for all times $t\in[0,T_{\textnormal{pos}})$. In particular, the points $(0,0,0)$ and $(2,0,0)$ are such points. At each of those two points, two or more great circles in $\widetilde{Q}$ intersect transversely. Since $\widetilde{Q}\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}})$, we see that the tangent plane to $M(t)$ at each of those two points is horizontal. Thus since (21) holds for $t=0$ (by (20)), it holds for all $t\in[0,T_{\textnormal{pos}})$. ∎
Next, we claim that
(22) $T_{\textnormal{pos}}<\infty.$
For, if not, then (by Theorem 32) there would be a $t<\infty$ for which $M(t)$ is a graph. But that is impossible by Claim 1.
By Theorem 35, the flow has a singularity at the origin at time $T_{\textnormal{pos}}$ of genus $\tilde{g}$, where $\tilde{g}$ is either $g$ or $2g$ and
$\tilde{g}\leq\operatorname{genus}(M)/2=g.$
Thus $\tilde{g}=g$.
Let $\Sigma_{g}$ be a shrinker at the spacetime point $(0,T_{\textnormal{pos}})$. By Theorem 35, it has all the properties asserted in Theorem 36 except for the property
$\operatorname{entropy}(\Sigma)<\delta,$
and the behavior of $\Sigma_{g}$ as $g\to\infty$.
By Theorem 37 below, we can choose $\Sigma_{g}$ to have entropy $<\delta$. By Theorem 38, given that entropy bound, $\Sigma_{g}$ has the asserted behavior as $g\to\infty$. ∎
Theorem 37.
In the proof of Theorem 36, we can choose the initial surface $M$ so that $\operatorname{entropy}(\Sigma)<\delta$.
Proof.
Let $\mathscr{M}^{\prime}_{g}$ be as in the proof of Theorem 36.
[…]
Note that there exist $M\in\mathscr{M}^{\prime}_{g}$ that are arbitrarily close to $\Psi_{r}$ in the weak sense. That is, there exists a sequence $M_{i}\in\mathscr{M}^{\prime}_{g}$ such that $M_{i}$ converges weakly to $\Psi_{r}$.
[…]
  1. (1)
    $M_{i}$ converges weakly with with multiplicity $1$ to $\{z=0\}$
  2. (2)
[…]
(Here $SS\SS(\infty)$ is the empty set.)
Let $T_{i}=T_{\textnormal{pos}}(M_{i})$ and, as usual, let
$t\in[0,T_{i}]\mapsto M_{i}(t)$
be the mean curvature flow with $M_{i}(0)=M_{i}$. Let $\Sigma_{i}$ be a shrinker to the flow at the origin at time $T_{i}$.
For $t\in[0,1)$, let
$M_{i}^{\prime}(t)=\frac{1}{\sqrt{T_{i}}}M_{i}(T_{i}t).$
Thus
$t\in[0,1)\mapsto M_{i}^{\prime}(t)$
is a mean curvature flow in $\mathbf{R}^{2}\cup Z_{\infty}$, but with the dilated metric $\gamma_{\lambda_{i}}$, where $\lambda_{i}:=(T_{i})^{-1/2}$.
Let $\operatorname{area}_{s}$ denote area with respect to the shrinker metric. (The shrinker metric is the Euclidean metric on $\mathbf{R}^{3}$ multiplied by the conformal factor $(4\pi)^{-1}\exp(-|p|^{2}/4)$.) By monotonicity,
$\operatorname{area}_{s}(\Sigma_{i})\leq\operatorname{area}_{s}(M_{i}^{\prime}(0))+\epsilon_{i},$
where $\epsilon_{i}\to 0$. (The $\epsilon_{i}$ is there because $M_{i}^{\prime}(\cdot)$ is mean curvature flow for a non-flat metric; and $\epsilon_{i}\to 0$ because the metric is converging to smoothly to the flat metric.)
[…]
$\operatorname{area}_{s}(\Sigma_{i})$ $\to\operatorname{area}_{s}(P)+\operatorname{area}_{s}(S)$
$\leq 1+d_{2}$
$<\delta.$
Hence
$\operatorname{entropy}(\Sigma_{i})=\operatorname{area}_{s}(\Sigma_{i})<\delta$
for all sufficiently large $i$. ∎
Theorem 38.
[…]
Theorem 38 was proved by Buzano, Nguyen, and Schulz [buzano]*Theorem 1.3. For the reader’s convenience, we include a somewhat different proof here.
Proof.
By Theorem 57, we can assume (after passing to a subsequence) that the $\Sigma_{g}$ converge to a limit $\Sigma$ consisting of the plane $\{z=0\}$ with multiplicity one together with one of the following:
  1. (1)
    The cylinder $\{x^{2}+y^{2}=2\}$ with multiplicity one, or
  2. (2)
[…]
In the cylindrical case (Case 1), $\Sigma$ would have entropy $1+d_{1}$, which is impossible since
$\operatorname{entropy}(\Sigma_{g})<\delta<1+d_{1},$
by Theorem 37.
[…]
Finally, for large $g$, $\Sigma_{g}$ has exactly one end by Lemma 60. ∎
8. Shrinkers with three ends
Let $g\geq 1$ be an integer, let $Q_{g}$ be the set of $g+1$ horizontal lines through the origin defined in §5, and let $G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ defined in §5. The goal of this section is to prove the following theorem.
Theorem 39.
[…]
  1. (1)
    $\Sigma_{g}$ has genus $2g$ and $3$ ends.
  2. (2)
    $\Sigma$ is $G_{g}$-invariant and $\Sigma_{g}\cap\{z=0\}=Q_{g}$.
  3. (3)
    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
  4. (4)
    The entropy of $\Sigma_{g}$ is bounded above by a constant independent of $g$.
Furthermore, as $g$ tends to infinity, $\Sigma_{g}$ converges to the plane $\{z=0\}$ with multiplicity $3$. After passing to a subsequence, the convergence is smooth away from one circle.
See Figure 7.
Figure 7. Conjectured qualitative appearance of the shrinker $\Sigma_{g}$ in case $g=1$. The surface shown has genus $2=2g$ and it seems to have $3$ ends. However, we know that $\Sigma_{g}$ has genus $2g$ and $3$ ends only when $g$ is large. Figure courtesy of Matthias Weber. (This figure is not based on a numerical approximation of a shrinker.)
Theorem 39 is a combination of the following results:
  1. (1)
    In Theorem 49, we prove, for each $g$, existence of a initial surface $M_{g}$ that has various nice properties, including a property called criticality.
  2. (2)
    Let $\Sigma_{g}$ be a shrinker to $M_{g}(\cdot)$ at the origin at time $T_{\textnormal{pos}}$. In Theorem 51, we show that $\Sigma_{g}$ has most of the properties asserted in Theorem 39.
  3. (3)
    In Theorem 55 and Proposition 58, we show, for all sufficiently large $g$, that $\Sigma_{g}$ has the remaining properties asserted in Theorem 39: it has genus $2g$, it has exactly $3$ ends, and it converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$.
To prove these theorems, we work in
[…]
with the metric $\gamma=\gamma_{1}$, as defined at the beginning of Section 6. We let $\mathscr{M}_{g}^{\prime\prime}$ be the set of $M\in\mathscr{M}_{g}$ such that $M$ is a compact, connected surface of genus $4g$, where $\mathscr{M}_{g}$ is as in Definition 26.
Proposition 40.
The class $\mathscr{M}_{g}^{\prime\prime}$ is nonempty.
Proof.
We prove Proposition 40 by defining a certain one-parameter family of surfaces in $\mathscr{M}_{g}^{\prime\prime}$. We will use the one-parameter family later.
First, we fix a small $R\in(0,2)$. (How small will be specified later. The choice of $R$ will not depend on $g$.)
Figure 8. This figure presents a quarter of the surface $M[s]$ for $g=1$. The complete surface is generated by successively rotating this piece by $180^{\circ}$ around the $x$-axis and the geodesic $C$. The resulting surface has genus $4$.
Let
$D$ =$=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq 2\},$
$D(R)$ =$=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq R\},$
$A^{+}$ $=\partial(B(0,R)\cap\{z>0\}),$
$A^{-}$ $=\partial(B(0,R)\cap\{z<0\}).$
[…]
Let $k$ be an integer, and consider the $2(g+1)$ wedge regions
$W_{k}:=\frac{\pi k}{g+1}<\theta<\frac{\pi(k+1)}{g+1}.$
Let $\Omega(s)$, $10\leq s<$ 1, be a smooth, $1$-parameter family of nested, closed convex regions in $W_{0}\cap D(R)$ such that $\Omega(0)$ is a single point and such that $\cup_{s}\Omega(s)$ is all of $W_{0}\cap D(R)$.
For $s\in[0,1)$, define $h(s)$ by
(23) $h(s)=\frac{\operatorname{dist}(0,\Omega(s))^{2}}{(g+1)}.$
Translate $A^{+}$ upward by $h(s)$ and translate $A^{-}$ downward by $h(s)$ to get
$A^{+}_{s}$ $:=A^{+}+(0,0,h(s)),$
$A^{-}_{s}$ $:=A^{-}+(0,0,-h(s)).$
For $10<s<$ 1, we do surgery on $D\cup A^{+}_{s}\cup A^{-}_{s}$ as follows. For each $W_{k}$ with $k$ even, we connect $D$ to $A^{+}_{s}$ be a neck in the region $W_{k}$. Likewise, for each $W_{k}$ with $k$ odd, we connected $D$ to $A^{-}_{s}$ by a neck in $W_{k}$. We do this in a way that preserves the $G_{g}$ symmetry, so it suffices to describe the surgery in $W_{0}$. The surgery in $W_{0}$ is the following: remove $\Omega(s)$ from $D$, remove the corresponding region
$\Omega(s)+(0,0,h(s))$
from $A^{+}_{s}$, and then attach the annular surface
[…]
The resulting surface has $(g+1)$ necks in $\{0\leq z\leq h(s)\}$ and $(g+1)$ necks in $\{-h(s)\leq z\leq 0\}$. Now let $M[s]$ be the union of that surface and its image under rotation by $\pi$ about $C=\partial D$. The resulting surface is $\widetilde{G}$-invariant.
Let
$\operatorname{neck}[s]:=(\partial\Omega(s))\times[0,h(s)]$
We have:
(24) $\operatorname{area}(\operatorname{neck}[s])$ $=h(s)\operatorname{length}(\partial\Omega(s))$
$\leq h(s)\operatorname{length}(\partial(D(R)\cap W_{0}))$
$=h(s)(2R+(\pi R/(g+1))$
$\leq h(s)(2\pi R)$
$\leq(g+1)^{-1}\operatorname{dist}(0,\Omega(s))^{2}2\pi R$
$\leq(g+1)^{-1}2\pi R^{3},$
where the second-to-last inequality follows from (23). The same bound holds for $\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$ since $\gamma_{ij}\leq\delta_{ij}$ everywhere. Consequently, since $M[s]$ has $4(g+1)$ necks,
$\operatorname{area}_{\,\gamma}(M[s])$ $\leq\operatorname{area}_{\,\gamma}(\{z=0\})+4\operatorname{area}_{\,\gamma}(A^{+}_{s})+4(g+1)\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$
$\leq 4\pi+4\operatorname{area}(A^{+}_{s})+4(g+1)(g+1)^{-1}2\pi R^{3}$
$=4\pi+12\pi R^{2}+8\pi R^{3}.$
We fix an $R$ for which this last expression is $<5\pi$. Thus
$\operatorname{area}_{\,\gamma}(M[s])<5\pi$
for all $s\in(0,1)$.
Now the surface $M[s]$ is only piecewise smooth. But we can smooth it (preserving the symmetries) so that the resulting surface $M^{\prime}[s]$ is smooth, and so that
$M^{\prime}[s]\cap\{z=0\}=\widetilde{Q}.$
We can do the smoothing in such a way that $M^{\prime}[s]$ depends continuously on $s$, and also so that
$\operatorname{area}_{\,\gamma}(M^{\prime}[s])<5\pi.$
From the construction, it is straightforward to check that $M^{\prime}[s]$ has genus $4g$ and that $M^{\prime}[s]\setminus\{z=0\}$ has genus 00. Thus $M^{\prime}[s]\in\mathscr{M}_{g}^{\prime\prime}$. ∎
Remark 41.
[…]
(25) $D\cup A_{0}^{+}\cup A_{0}^{-}\cup I,$
where $I$ is the union of $2(g+1)$ vertical line segments. (The necks collapse to the segments $I$.) Recall that $A_{0}^{+}$ is $A^{+}$ translated upward by $h(0)$, and that $h(0)>0$. (See (23).) We can do the smoothing in such a way that the surfaces $M^{\prime}[s]\cap U$ converge to the same limit set (25) as $s\to 0$.
[…]
Lemma 42.
There is a constant $\Lambda<\infty$ (independent of $g$) with the following property:
$\frac{\operatorname{area}(M[s]\cap B(0,r))}{\pi r^{2}}<\Lambda$
for all $s\in(0,1)$ and all $20<r<$ 2.
Proof.
Note that
$\operatorname{area}(B(0,r)\cap(D\cup A^{+}_{s}\cup A^{-}_{s})$ $=\pi r^{2}+2\operatorname{area}(B(0,r)\cap A^{+}_{s})$
$\leq\pi r^{2}+2\operatorname{area}(B(0,r)\cap A^{+})$
=$=\begin{cases}3\pi r^{2}&\text{if $r<R$}\\ \pi r^{2}+6\pi R^{2}&\text{if $r>R$}.\end{cases}$
$\leq 7\pi r^{2}.$
Thus, if $\mathscr{R}$ is the union of the necks we attached in the surgery to produce $M[s]$, then
(26) $\operatorname{area}(M[s]\cap B(0,r))<7\pi r^{2}+\operatorname{area}(B(0,r)\cap\mathscr{R})).$
Now we need to bound the area of $B(0,r)\cap\mathscr{R}$. We may assume
(27) $r>\operatorname{dist}(0,\Omega(s)),$
as otherwise $B(0,r)\cap\mathscr{R}$ is empty. Since $B(0,r)$ touches at most $2(g+1)$ necks, we have
$\operatorname{area}(B(0,r)\cap\mathscr{R})$ $=2(g+1)\operatorname{area}(B(0,r)\cap\operatorname{neck}[s])$
$\leq 2(g+1)\operatorname{area}(\operatorname{neck}[s])$
$\leq(\operatorname{dist}(0,\partial\Omega(s))^{2}4\pi R$
$\leq 4\pi Rr^{2}$
by (24) and (27). Thus, by (26),
$\operatorname{area}(M[s]\cap B(0,r))\leq(7+4R)\pi r^{2},$
so we have proved Lemma 42 with $\Lambda=7+4R$. ∎
Remark 43.
Recall that the piecewise smooth surface $M[s]$ was smoothed to make the smooth surface $M^{\prime}[s]$. It follows from Lemma 42 that we can do the smoothing in such a way that $M^{\prime}[s]$ satisfies
$\frac{\operatorname{area}(B(0,r)\cap M^{\prime}[s])}{\pi r^{2}}<\Lambda$
for all $r\in(0,2)$.
Definition 44.
Let $\mathscr{F}=\mathscr{F}_{g}$ be the family of smooth surfaces $M^{\prime}[s]$ constructed in the proof of Proposition 40, where the smoothing has been done in accord with Remarks 41 and 43.
Theorem 45.
Suppose that $M\in\mathscr{M}_{g}^{\prime\prime}$. If $t<T_{\textnormal{pos}}(M)$ is a regular time, then $M(t)\in\mathscr{M}_{g}$, and the number of points in $M(t)\cap Z^{+}$ is even. Furthermore, if we give $M$ the orientation such that $\nu_{M}(2,0,0)=\mathbf{e}_{3}$, then
$\nu_{M(t)}(0,0,0)=-\mathbf{e}_{3}$
for all $t\in[0,T_{\textnormal{pos}}(M))$, where the orientation is the standard induced orientation as defined in Section 3.
Proof.
By Theorem 31, $M(t)\in\mathscr{M}_{g}$.
Give $M$ and $M(\cdot)$ the indicated orientations. By Lemma 29 and Lemma 34, $(0,0,0)$ and $(2,0,0)$ are regular points at all times $t<T_{\textnormal{pos}}$. Since $\widetilde{Q}\subset M(t)$ (by Lemma 28), we see that the tangent plane at each of those points is horizontal. Thus
(28) $\nu_{M(t)}(2,0,0)=\nu_{M}(2,0,0)=\mathbf{e}_{3},$
(29) $\nu_{M(t)}(0,0,0)=\nu_{M}(0,0,0)$
for all $t\in[0,T_{\textnormal{pos}})$.
Let $t\in[0,T_{\textnormal{pos}})$ be a regular time, let $n(t)$ be the number of points of $M(t)\cap Z^{+}$, and let $K(t)$ be the closed region bounded by $M(t)$ such that $\nu_{M(t)}(\cdot)$ is the unit normal that points out of $K(t)$.
Since $(2,0,z)\in M(t)$ only for $z=0$, we see that
$(1,0,z)\in K(t)\iff z\leq 0.$
Thus if $p$ is any point for which
$z(p)>\max_{q\in M(t)}z(\cdot),$
then $p\notin K(t)$. Consequently, at the highest point of $M(t)\cap Z$, $\nu_{M(t)}=\mathbf{e}_{3}$, and the next highest point, $\nu_{M(t)}=-\mathbf{e}_{3}$, and so on. Thus
$\nu_{M(t)}(0,0,0)=\begin{cases}\mathbf{e}_{3}&\text{if $n(t)$ is even,}\\ -\mathbf{e}_{3}&\text{if $n(t)$ is odd.}\end{cases}$
Thus, by (29), the parity of $n(t)$ is the same for all regular times $t\in[0,T_{\textnormal{pos}})$.
Now at time 00,
$4g$ $=\operatorname{genus}(M)$
$=2\operatorname{genus}(M\cap\{(x^{2}+y^{2})^{1/2}<2\})$
$=2n(0)$
by Lemma 22. Thus $n(t)$ is even for all regular times $t\in[0,T_{\textnormal{pos}})$. ∎
Definition 46.
We let
$\mathscr{U}_{g}=\{M\in\mathscr{M}_{g}^{\prime\prime}:T_{\textnormal{pos}}(M)=\infty\}.$
Lemma 47.
The set $\mathscr{U}_{g}$ is an open subset of $\mathscr{M}_{g}^{\prime\prime}$ with respect to smooth convergence.
Proof.
Suppose that $M\in\mathscr{U}_{g}$ and that $M_{n}\in\mathscr{M}_{g}^{\prime\prime}$ converges smoothly to $M$. Let $T_{n}=T_{\textnormal{pos}}(M_{n})$. We must show that $T_{n}=\infty$ for all sufficiently large $n$.
First, note that $T_{n}\to\infty$. For if not, we can assume (after passing to a subsequence) that $T_{n}$ converges to a finite limit $T$. By Lemma 34, the flow $M_{n}(\cdot)$ has a singularity at the origin at time $T_{n}$. Thus the flow $M(\cdot)$ also has a singularity at the origin at time $T$ (by the Local Regularity Theorem, Theorem 7). But that implies (by Lemma 34) that $T_{\textnormal{pos}}(M)=T$, a contradiction. Thus $T_{n}\to\infty$.
By Theorem 32, there is a time $\tau<\infty$ such that $M(\tau)$ is a smooth graph. Since $T_{n}\to\infty$, we can assume that $T_{n}>\tau$ for all $n$. By the Local Regularity Theorem (Theorem 7), $M_{n}(\tau)$ converges smoothly to $M(\tau)$. Thus $M_{n}(\tau)$ is a smooth graph if $n$ is sufficiently large. But smooth graphs never develop singularities, and thus $T_{\textnormal{pos}}(M_{n})=\infty$ for such $n$. ∎
Definition 48.
A surface $M\in\mathscr{M}_{g}^{\prime\prime}$ is called critical if
$M\in\overline{\mathscr{U}_{g}}\setminus\mathscr{U}_{g},$
i.e., provided $M\notin\mathscr{U}_{g}$ and there exist $M_{n}\in\mathscr{U}_{g}$ that converges smoothly to $M$.
Theorem 49.
Let $\mathscr{F}$ be the family of surfaces in $\mathscr{M}_{g}^{\prime\prime}$ constructed in the proof of Proposition 40. Then $\mathscr{F}$ contains a critical surface $M$.
Proof.
Note that $\mathscr{F}$ is a connected family. It suffices to show that $\mathscr{F}\cap\mathscr{U}_{g}$ and $\mathscr{F}\setminus\mathscr{U}_{g}$ are both nonempty. For, in that case, since $\mathscr{F}\cap\mathscr{U}_{g}$ is an open subset of $\mathscr{F}$, it cannot be also be closed (by connectedness of $\mathscr{F}$). Thus there is a surface $M$ in $\mathscr{F}\setminus\mathscr{U}_{g}$ that is a limit of surfaces in $\mathscr{F}\cap\mathscr{U}_{g}$. By definition, such an $M$ is critical.
Claim 1.
The family $\mathscr{F}\,\cap\,\mathscr{U}_{g}$ is nonempty: there exist $M\in\mathscr{F}$ for which $T_{\textnormal{pos}}(M)=\infty$.
Let
(30) $U:=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}.$
To prove Claim 1, consider a sequence $M_{n}=M^{\prime}[s_{n}]$ in $\mathscr{F}$ for which $s_{n}\to 0$. Thus the $M_{n}\cap\overline{U}$ converge as sets to
$L:=D\cup A_{0}^{+}\cup A_{0}^{-}\cup I,$
where $I$ is the union of $2(g+1)$ vertical line segments. (See Remark 41.) In particular, there is an $\epsilon>0$ such that
(31) $L\cap\{0<|z|<\epsilon\}=I.$
By the Initial Regularity Theorem (Theorem 8), there is a time $\tau>0$ such that for all sufficiently large $n$, the flow $M_{n}(\cdot)$ has no singularities on $Z$ during the time interval $[0,\tau]$. In particular, for $t\in[0,\tau]$, $M_{n}(t)\cap Z$ consists of exactly $5$ points, and those points depend smoothly on $t$. Thus, by replacing $\tau$ by a smaller $\tau>0$, we can assume that
(32) $M_{n}(t)\cap Z\cap\{|z|\leq\epsilon/2\}=\{O\}$ for all $t\in[0,\tau]$.
Using Angenent Tori as barriers, or by Brakke’s Clearing Out Lemma, it follows from (31) that
(33) $M_{n}(t)\cap\{|z|=\epsilon/2\}=\emptyset$
for $t=\tau/2$ and for all sufficiently large $n$. It follows (for such $n$) that (33) holds for all $t\geq\tau/2$.
Let $t\in(\tau/2,\tau)$ be a regular time for $M_{n}(\cdot)$.
Let $M_{n}^{c}(t)$ be the connected component of $M_{n}(t)$ that contains $\widetilde{Q}$. By (33), $M_{n}^{c}(t)$ lies in $\{|z|<\epsilon/2\}$ Thus, by (32), $M^{c}_{n}(t)$ is disjoint from $Z^{+}$. Hence
$\operatorname{genus}(M_{n}(t))=0$
by Proposition 27. Therefore $T_{\textnormal{pos}}(M_{n})=\infty$. This completes the proof of Claim 1.
Claim 2.
The family $\mathscr{F}\setminus\mathscr{U}_{g}$ is nonempty: there exist $M\in\mathscr{F}$ for which $T_{\textnormal{pos}}(M)<\infty$.
To prove Claim 2, let $M_{n}=M[s_{n}]$ be a sequence $\mathscr{F}$ for which $s_{n}\to 1$. We claim that there are $n$ for which $T_{\textnormal{pos}}(M_{n})<\infty$. Suppose, to the contrary, that $T_{\textnormal{pos}}(M_{n})=\infty$ for each $n$. By passing to a subsequence, we can assume that the $M_{n}(\cdot)$ converge weakly to a mean curvature flow $t\in[0,\infty)\mapsto M(t)$.
By Remark 41, the surfaces
$M_{n}\cap U,$
[…]
$M_{n}(\tau)\cap\{(x^{2}+y^{2})^{1/2}<R/2\}$
is the union of three smooth graphs. Note we can choose $\tau$ to be a regular time for all the $M_{n}(\tau)$. Then for $n\geq n_{\tau}$,
$M_{n}(\tau)\cap Z^{+}$
contains exactly one point, which is impossible by Theorem 45. This completes the proof of Claim 2, and therefore the proof of Theorem 49. ∎
By Lemma 34, we know that at time $T_{\textnormal{pos}}(M)<\infty$, the flow $M(t)$ has a singularity at the origin with positive genus.
Definition 50.
Let $\mathscr{S}_{g}$ be the collection of shrinkers $\Sigma$ that occur at the spacetime point $(O,T_{\textnormal{pos}}(M))$ for a surface $M\in\mathscr{F}_{g}$ that is critical in the sense of Definition 48.
Theorem 51.
Suppose $\Sigma\in\mathscr{S}_{g}$. Then
  1. (1)
    $\Sigma$ is $G_{g}$-invariant,
  2. (2)
    $\Sigma\setminus\{z=0\}$ has genus 00,
  3. (3)
    $\Sigma\cap\{z=0\}=Q_{g}$,
  4. (4)
    $\operatorname{genus}(\Sigma)$ is $g$ or $2g$.
  5. (5)
    The entropy of $\Sigma$ is bounded above by a constant independent of $g$.
Proof.
Trivially, $\Sigma_{g}$ is $G_{g}$-invariant. Note that $M(t)\setminus\{z=0\}$ has genus 00 for all regular $t<T_{\textnormal{pos}}$ by Theorem 45. Hence $\Sigma\setminus\{z=0\}$ has genus 00. By Theorem 35, $\Sigma\cap\{z=0\}=Q_{g}$ and $\Sigma$ has genus $g$ or $2g$. Finally, the uniform entropy bound follows from Remark 43 and Huisken monotonicity. ∎
Conjecture 52.
If $\Sigma$ is as in Theorem 51, then $\Sigma$ has genus $2g$.
The conjecture is true if $g$ is sufficiently large by Theorem 55 below.
Definition 53.
A special Brakke flow is a unit-regular, integral Brakke flow
$t\in I\mapsto M(t)$
that admits a strong orientation $\nu$ with the following property: for each $t$, if the origin is a regular point at time $t$, then $\nu(O,t)=-\mathbf{e}_{3}$.
Lemma 54.
Suppose $\Sigma\in\mathscr{S}_{g}$. Then there is a $G_{g}$-invariant, special Brakke flow
$t\in\mathbf{R}\mapsto M^{\prime}(t)$
in $\mathbf{R}^{3}$ such that
$M^{\prime}(t)=|t|^{-1/2}\Sigma\quad\text{for all $t<0$}.$
Proof.
Let $T=T_{\textnormal{pos}}(M)$. Let $M\in\mathscr{M}_{g}^{\prime\prime}$ be a critical surface such that $\Sigma$ is a shrinker to $M(\cdot)$ at the spacetime point $(O,T)$. Since $M$ is critical, there exist $M_{n}\in\mathscr{M}_{g}^{\prime\prime}$ such that $T_{\textnormal{pos}}(M_{n})=\infty$ and such that $M_{n}$ converges smoothly to $M$.
Note that $M_{n}(\cdot)$ is special by Theorem 45.
By passing to a subsequence, we can assume that the flow $M_{n}(\cdot)$ converge to a unit-regular Brakke flow
$t\in[0,\infty)\mapsto\tilde{M}(t).$
Now
$\tilde{M}(0)=M,$
and thus
$\tilde{M}(t)=M(t)\quad\text{for $0\leq t\leq\operatorname{T_{\textnormal{fat}}}(M)$}.$
By the Local Regularity Theorem (Theorem 7), $\tilde{M}(\cdot)$ is special since each $M_{n}(\cdot)$ is.
Now
(34) $t\in(-\infty,0)\mapsto|t|^{-1/2}\Sigma$
occurs as a tangent flow to $M(\cdot)$ at the spacetime point $(0,T)$, so it also occurs as a tangent flow to $\tilde{M}(\cdot)$ at that point, since the two flows coincide up to time $T$. Thus there is a sequence $\lambda_{n}\to\infty$ such that the parabolically dilated flows
(35) $t\in[-\lambda_{n}T,\infty)\mapsto\lambda_{n}^{1/2}\tilde{M}(T+\lambda_{n}^{-1}t)$
converge for $t<0$ to the flow (34).
By passing to a subsequence (if necessary) we can assume that the flows converge for all $t$ a flow
(36) $t\in\mathbf{R}\mapsto M^{\prime}(t)$
Since the $\tilde{M}(\cdot)$ is special, so are the flows (35), and therefore (by the Local Regularity Theorem, Theorem 7) so is the flow (36). ∎
Theorem 55.
Suppose, for each $g$, that $\Sigma_{g}\in\mathscr{S}_{g}$. Then, for all sufficiently large $g$, the genus of $\Sigma_{g}$ is $2g$. Moreover, $\Sigma_{g}$ converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$, and the convergence is smooth except along a circle $C(r)$ in the plane $\{z=0\}$.
Proof.
Suppose the theorem is false. In Theorem 57 in following section, we analyze all possible subsequential limits as $g\to\infty$ of surfaces like the $\Sigma_{g}$ in Theorem 51. According to that theorem, if the conclusion of Theorem 55 is false, then (by passing to a subsequence) we can assume that the $\Sigma_{g}$ converge with multiplicity $1$ to a limit
$\Sigma=P+S$
[…]
By Lemma 54, for each $g$, there is a $G_{g}$-invariant, special Brakke flow
$t\in\mathbf{R}\mapsto M_{g}^{\prime}(t)$
such that
$M_{g}^{\prime}(t)=|t|^{-1/2}\Sigma_{g}\quad\text{for all $t<0$}.$
By passing to a further subsequence, we can assume that the flows $M_{g}^{\prime}(\cdot)$ converge to a flow
$t\in\mathbf{R}\mapsto M^{\prime}(t).$
Note that
$M^{\prime}(g)=|t|^{-1/2}\Sigma\quad(t<0).$
Note that $M^{\prime}(0)$, the limit of $|t|^{-1/2}\Sigma$ as $t\to 0$ is the $P$, the plane $\{z=0\}$ with multipicity $1$. Hence $M^{\prime}(t)=P$ for all $t\geq 0$. That is, $M^{\prime}(\cdot)$ is the flow
(37) $t\in\mathbf{R}\mapsto\begin{cases}|t|^{-1/2}\Sigma&(t<0),\\ \Sigma&(t\geq 0).\end{cases}$
Since $M^{\prime}(\cdot)$ is a limit of special flows, it is also special.
Note that at each time $t=(-1)$, the singular set is the circle $J:=S\cap P$. Thus, by specialness, there is a closed set $K$ such that $K\setminus J$ is a smooth manifold-with-boundary. At the origin, $-\mathbf{e}_{3}$ points out of $K$. It follows that $K$ consists of the closure of the region above $P$ and inside $S$, together with the closure of region the region below $P$ and outside $S$. See Fig. 9.
Figure 9. The region $K$ (in the case that $S$ is a sphere)
By specialness, at time $t=1$,
Thus $\nu(p,-1)=\mathbf{e}_{3}$ at all points $p$ in $\{z=0\}$ that lie outside $S$.
By the same argument (or simply by continuity)
$\nu(p,t)=\mathbf{e}_{3}$
for all $t<0$ and $p$ in $\{z=0\}$ outside $|t|^{-1/2}S$.
Hence (by continuity)
$\nu(p,0)=\mathbf{e}_{3}$
for all $p\neq 0$ in $\{z=0\}$. Therefore, also by continuity,
$\nu(p,t)=\mathbf{e}_{3}$
for all $p$ in $\{z=0\}$ and $t>0$. In particular, $\nu(0,t)=\mathbf{e}_{3}$ for all $t>0$, contrary to specialnesss of $M^{\prime}(\cdot)$. ∎
9. Limits of Shrinkers
In this section, $\mathscr{G}$ is an infinite set of positive integers, and for each $g\in\mathscr{G}$, $\Sigma_{g}$ is a connected, properly embedded, $G_{g}$-invariant shrinker such that
  1. (1)
    $\Sigma_{g}\cap\{z=0\}=Q_{g}$.
  2. (2)
    $0<\operatorname{genus}(\Sigma_{g})\leq 2g$.
  3. (3)
    $\operatorname{genus}(\Sigma\cap\{z>0\})=0$.
  4. (4)
    $\sup_{g\in\mathscr{G}}\operatorname{entropy}(\Sigma_{g})<\infty$.
(The notation is as in Section 5.) Examples of such shrinkers were produced in §7 and §8. We will analyze the behavior of $\Sigma_{g}$ as $g\to\infty$. In this section, “passing to a subsequence” means replacing the set $\mathscr{G}$ by an infinite subset.
Lemma 56.
Let $G_{g}^{+}$ be the subgroup of $G_{g}$ consisting of $\sigma$ such that $\sigma(0,0,z)=(0,0,z)$ for all $z$. Let $\Gamma_{g}$ be a smooth, $G^{+}_{g}$-invariant, simple closed curve in $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$. Then $\Gamma_{g}$ divides $\Sigma_{g}$ into two components. Let $U_{g}$ be the component that does not contain $Q_{g}$. Then
$E_{g}:=U_{g}\cup\Gamma_{g}$
lies in $\{z>0\}$ and is properly embedded in $\mathbf{R}^{3}$.
Proof.
The proof of [brendle]*Theorem 2 shows that $\Sigma_{g}^{+}$ is connected. Since $\Sigma_{g}^{+}$ has genus 00, $\Gamma_{g}$ divides $\Sigma_{g}^{+}$ into two connected components $U_{g}$ and $U_{g}^{\prime}$. Note that $U_{g}$ and $U_{g}^{\prime}$ are each $G_{g}^{+}$-invariant. Since $\Sigma_{g}$ is connected, the closure of at least one component, which we may assume is $U_{g}^{\prime}$, contains points in $Q_{g}$. It follows that $\overline{U_{g}^{\prime}}$ contains an entire ray of $Q$. By $G_{g}^{+}$ invariance, $\overline{U_{g}^{\prime}}$ contains all of $Q$. It follows that $\Gamma_{g}$ divides $\Sigma_{g}$ into two components, namely $U_{g}$ and
$U_{g}^{\prime\prime}:=U_{g}^{\prime}\cup(\Sigma_{g}\cap\{z\leq 0\}).$
Now let $E_{g}=U_{g}\cup\Gamma_{g}$. Since $E_{g}$ contains $\Gamma_{g}$, and is disjoint from $Q_{g}=\Sigma_{g}\cap\{z=0\}$, we see that $E_{g}$ lies in $\{z>0\}$. As $\Sigma_{g}$ is properly embedded in $\mathbf{R}^{3}$ and as $\Gamma_{g}$ is a curve that divides $\Sigma_{g}$ into two components, one of which is $U_{g}$, then $E_{g}$ is properly embedded in $\mathbf{R}^{3}.$
Theorem 57.
After passing to a subsequence, the $\Sigma_{g}$ converge to a shrinker $\Sigma$, and the convergence is smooth away from a circle
$C(r):=\{(x,y,0):x^{2}+y^{2}=r^{2}\}.$
Furthermore, one of the following holds:
  1. (1)
    $\Sigma$ is the plane $\{z=0\}$ with multiplicity $3$, and $\Sigma_{g}$ has genus $2g$ for all sufficiently large $g$.
  2. (2)
    $\Sigma$ is the plane $\{z=0\}$ together with the sphere $\partial B(0,2)$, each with multiplicity $1$.
  3. (3)
    $\Sigma$ is the plane $\{z=0\}$ together with the cylinder $\{x^{2}+y^{2}=2\}$, each with multiplicity $1$.
Proof.
By passing to a subsequence, we can assume that the $\Sigma_{g}$ converge as sets to a limit set $\Sigma$. We can also assume that there is a closed set $K$ (the curvature blowup set) such the convergence $\Sigma_{g}$ to $\Sigma$ is smooth (perhaps with multiplicity) in compact subsets of $\mathbf{R}^{3}\setminus K$ and such that
$\sup_{B(p,\epsilon)}|A(\Sigma_{g},\cdot)|\to\infty$
for every $p\in K$ and every $\epsilon>0$, where $A(\Sigma_{g},\cdot)$ is the second fundamental form of $\Sigma_{g}$.
Since $Q_{g}\subset\Sigma_{g}$, it follows that $\{z=0\}$ is contained in $\Sigma$, and thus that
(38) $\Sigma$ =$=\{z=0\}\cup\Sigma^{+}\cup\Sigma^{-}$
=$=\{z=0\}\cup\Sigma^{+}\cup\rho_{X}\Sigma^{+},$
where $\Sigma^{+}=\Sigma\cap\{z>0\}$, $\Sigma^{-}=\Sigma\cap\{z<0\}$, and $\rho_{X}$ denotes rotation by $\pi$ about the $x$-axis, $X$.
By assumption, $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$ has genus 00. By the compactness theorem [white-compact]*Theorem 1.1 for minimal surfaces of locally bounded area and genus, $\Sigma^{+}$ is smooth and $K\cap\{z>0\}$ is a discrete subset of $\{z>0\}$. The proof of [brendle]*Theorem 2 shows that $\Sigma^{+}$ is unstable (as a minimal surface for the shrinker metric) and connected. Thus (again by the compactness theorem [white-compact]*Theorem 1.1), the convergence of $\Sigma_{g}^{+}$ to $\Sigma^{+}$ is smooth and with multiplicity $1$, and therefore $K\cap\{z>0\}$ is empty. Therefore, by symmetry,
(39) $K\subset\{z=0\}.$
Claim 1.
One of the following holds:
  1. (1)
    $\Sigma^{+}$ is empty.
  2. (2)
    $\overline{\Sigma^{+}}$ is a disk bounded by a circle $C(r)$ in the plane $\{z=0\}$.
  3. (3)
    $\overline{\Sigma^{+}}$ is an annulus whose boundary is a circle $C(r)$ in the plane $\{z=0\}$.
Proof of Claim 1.
Suppose $\Sigma^{+}$ is nonempty. let $C^{\prime}$ be a horizontal circle in $\Sigma^{+}$. Let $C_{q}^{\prime}$ be smooth, simple closed curves in $\Sigma_{q}$ that converge smoothly to $C^{\prime}$. By Lemma 56, $C_{q}^{\prime}$ bounds a region $E_{n}$ in $\Sigma_{n}$ that is either a disk or an annulus, that lies in $\{z>0\}$, and that is properly embedded in $\mathbf{R}^{3}$. Passing to the limit, we see that $C^{\prime}$ bounds a region $E$ in $\Sigma^{+}$ that is either a disk or an annulus, and that is properly embedded in $\mathbf{R}^{3}$.
The other end $E^{\prime}$ of $\Sigma^{+}$ must diverge to infinity, converge to the origin, or converge to a circle $C(r)$ in the plane $\{z=0\}$. The end $E^{\prime}$ cannot diverge to infinity, because then $\Sigma^{+}$ would be shrinker (without boundary) in $\{z>0\}$. The end $E^{\prime}$ cannot converge the origin, for if it did, then $\Sigma^{+}\cup\{0\}$ and $\{z=0\}$ would violate the strong maximum principle at the origin. (One can invoke various versions of the maximum principle here. For example, the classical strong maximum principle suffices, because $\Sigma^{+}\cup\{0\}$ is a smooth minimal surface with respect to the shrinker metric by [structure]*Corollary 2.7.)
Thus the end $E^{\prime}$ converges to a circle $C(r)$ in the plane $\{z=0\}$. This completes the proof of Claim 1. ∎
Recall (Corollary 24) that there exist $r_{g}\in(0,\infty)$ and $R_{g}\in[r_{g},\infty]$ such that
[…]
By the compactness theorem [white-compact]*Theorem 1.1 for minimal surfaces of locally bounded area and genus,
[…]
Therefore, since $K\subset\{z=0\}$ (see (39)),
(40) $K\subset\{0\}\cup C(r_{\infty})\cup C(R_{\infty}).$
where $C(r)$ is the circle of radius $r$ about the origin in the plane $\{z=0\}$.
Claim 2.
Suppose that $B$ is a ball centered at a point $p$ in the $x$-axis, $X$, and that $\Sigma_{g}\cap B$ converges smoothly with some multiplicity $m$ to $\{z=0\}\cap B$. Then $m$ is odd.
Proof.
By rotational symmetry by $\pi$ about $X$, If $L$ is a closed vertical line segment centered at $p$, then $L\cap\Sigma_{g}$ has an odd number of points (namely, the point $p$ and equal numbers of points above and below $p$.) ∎
Claim 3.
The multiplicity of $\Sigma$ is the same in each component of
(*) $\{z=0\}\setminus(\{0\}\cup C(r_{\infty})\cup C(R_{\infty})).$
Furthermore, if $\Sigma^{+}$ is nonempty, then it meets $\{z=0\}$ orthogonally.
Proof.
Note that $\Sigma$ with its multiplicities is a stationary varifold with respect to the shrinker metric. Let $\Omega$ and $\Omega^{\prime}$ be two components of (*3) with a common boundary circle $C(r)$, where $\Omega$ is the inner component. (Thus $r$ is $r_{\infty}$ or $R_{\infty}$.) Let $m$ be the multiplicity in $\Omega$ and $m^{\prime}$ be the multiplicity in $\Omega^{\prime}$. It suffice to show that $m=m^{\prime}$.
Case 1: Either $\Sigma^{+}$ is empty, or $\Sigma^{+}$ is nonempty and $r$ is not equal to the radius of $\partial\Sigma^{+}$. Then $m=m^{\prime}$ follows immediately from stationarity.
Case 2: $\Sigma^{+}$ is nonempty and $r$ is the radius of $\partial\Sigma^{+}$.
Consider the point $p=(r,0,0)$. Let
$\nu^{+}=(\cos\theta,0,\sin\theta)$
be the unit vector that is normal to $C(r)$, tangent to $\Sigma^{+}$ and that points into $\Sigma^{+}$. Thus we can choose $\theta\in(0,\pi)$. Then
$\nu^{-}=(\cos\theta,0,-\sin\theta)$
is the corresponding vector that points into $\Sigma^{-}$. By stationarity,
$-m\mathbf{e}_{1}+m^{\prime}\mathbf{e}_{1}+\nu^{+}+\nu^{-}=0.$
Taking the dot product with $\mathbf{e}_{1}$ gives
$m-m^{\prime}=2\cos\theta.$
and thus
(41) $|m-m^{\prime}|=2|\cos\theta|<2.$
Since $m$ and $m^{\prime}$ are both odd, $m-m^{\prime}$ is even. Therefore, by (41), $m=m^{\prime}$ and $\theta=\pi/2$. ∎
Claim 4.
The convergence $\Sigma_{g}\to\Sigma$ is smooth at the origin.
Proof.
Choose $r>0$, $r\neq r_{\infty}$, $r\neq R_{\infty}$, sufficiently small that $\overline{B(0,r)}$ is disjoint from $\Sigma^{+}$ and such that $\Sigma_{g}\setminus B(0,r)$ is connected for all $g\in\mathscr{G}$. (Such an $r$ exists by Lemma 61.)
Now $\Sigma_{g}\cap\overline{B(0,r)}$ converges smoothly and with multiplicity $m$ away from the origin to $\overline{D(r)}$.
If $m=1$, the convergence is smooth by Allard’s Regularity Theorem. Thus we assume that $m>1$.
For all sufficiently large $g$, $\Sigma_{g}\cap\partial B(0,r)$ consists of $m$ simple closed curves. Let $\Gamma_{g}^{0}$ be the curve that passes through the points of $Q_{g}\cap\partial B(0,r)$. Let $\Gamma_{g}$ be one of the other components. By symmetry, we may assume that $\Gamma_{g}$ lies in $\{z>0\}$. Let $\Sigma_{g}$ be the component of $\Sigma_{g}\setminus\Gamma_{g}$ that lies in $\{z>0\}$. (The component exists by Lemma 56). Since $\Sigma^{g}\setminus B(0,r)$ is connected and contains points in $\{z\leq 0\}$, we see that $\Sigma_{g}$ is contained in $B(0,r)$.
We have shown: each component of $\Sigma_{g}\cap\partial B(0,r)$ other than $\Gamma_{g}^{0}$ bounds a connected component of $\Sigma_{g}\cap B(0,r)$. Hence the same is also true for $\Gamma_{g}^{0}$. Now those $m$ components together converge weakly to the disk $\overline{D(0,r)}$ with multiplicity $m$. Thus each one converges weakly to $\overline{D(0,r)}$ with multiplicity $1$. By the Allard Regularity Theorem, the convergence is smooth. ∎
We can now complete the proof of Theorem 57.
Case 1: $\Sigma^{+}$ is empty. Then (as a varifold) $\Sigma$ is the plane $\{z=0\}$ with some odd multiplicity $m=2c+1$, by Claims 2 and 3.
If $\Sigma$ were a multiplicity $1$ plane, then $\Sigma_{g}$ would be a multiplicity $1$ plane for all sufficiently large $g$ (by Lemma 59), contradicting the assumption that $\Sigma_{g}$ has positive genus. Thus $m>1$.
Let
$R\in(0,\infty)\setminus\{r_{\infty},R_{\infty}\}$
be such that $\Sigma_{g}\cap B(0,R)$ is connected for all sufficiently large $g\in\mathscr{G}$. (Such an $R$ exists by Lemma 61.) Note that $\Sigma_{g}\cap\partial B(0,R)$ consists of $m$ simple closed curves, $c$ of which lie in $\{z>0\}$ and wind around $Z$. Also, by the smooth convergence of $\Sigma_{g}$ to $\Sigma$ at the origin (Claim 4), $\Sigma_{g}\cap Z$ has $m$ points, $c$ of which lie in $Z^{+}$.
By Lemma 22,
$\operatorname{genus}(\Sigma_{g}\cap B(0,R))=2cg.$
Since $c>0$ and $\operatorname{genus}(\Sigma_{g})\leq 2g$, we see that $c=1$ and that $\operatorname{genus}(\Sigma_{g})=2g$.
It remains (in Case 1) only to show that $R_{\infty}=r_{\infty}$ (so that the convergence is smooth except along a single circle.) Suppose to the contrary that $r_{\infty}<R_{\infty}$. Let $r_{\infty}<R<R_{\infty}$. Thus we may assume that
$r_{g}<R<R_{g}$
for all $g\in\mathscr{G}$.
For all sufficiently large $g$, we have shown that, if $\Sigma_{g}\cap B(0,R)$ is connected, then it has genus $2g$, which is impossible since $R<R_{g}$. Thus we may assume, for all $g$, that $\Sigma_{g}\cap B(0,r)$ is not connected. Now it has at least $2$ components and at most $3$ (since $\Sigma_{g}\cap\partial B(0,R)$ has $3$ components.) The number of components is odd (by symmetry) and thus it is $3$. But now (exactly as in the proof of Claim 4) each of those components converges smoothly to the disk $\overline{D(0,R)}$, so $\Sigma_{g}\cap B(0,R)$ has genus 00, contradicting $r_{g}<R$. This completes the proof in Case 1.
Case 2: $\Sigma^{+}$ is nonempty. By Claim 1, its boundary is a circle $C(\rho)$. By Claim 3, $\Sigma^{+}$ meets $\{z=0\}$ orthogonally, so
$\Sigma:=\Sigma^{+}\cup\Sigma^{-}\cup C(\rho)$
is a smoothly embedded, rotationally invariant shrinker $\Sigma$. Thus $\Sigma$ is a sphere or a cylinder.
Now choose $R>0$ such that $R$ is not equal to $r_{\infty}$ or $R_{\infty}$, and large enough that $\Sigma_{g}\cap B(0,R)$ is connected for all $g\in\mathscr{G}$. (Such an $R$ exists by Lemma 61.)
If $\Sigma$ is a sphere, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m=2c+1$ simple closed curves, $c$ of which lie in $\{z>0\}$. Also, $\Sigma_{g}\cap Z^{+}$ consists $c+1$ points, $c$ of which are close to the origin and one of which is close to $\Sigma^{+}\cap Z$. Thus
$\operatorname{genus}(\Sigma_{g}\cap B(0,R))=(2c+1)g$
If, on the other, $\Sigma$ is a cylinder, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m+2$ simple closed curves, $m$ of which lie near the circle $C(R)$ and two of which are near the circles of intersection of the cylinder $\Sigma$ and the sphere $\partial B(0,R)$. Note that $c+1$ of those curves lie in $\{z>0\}$. Furthermore, $\Sigma_{g}\cap Z$ consist of $m$ points, $c$ of which lie in $Z^{+}$.
Thus, again,
(42) $\operatorname{genus}(\Sigma_{g}\cap B(0,R))=(2c+1)g.$
We have shown that (42) holds whether $\Sigma$ is a sphere or a cylinder. Since $\Sigma_{g}$ has genus $\leq 2g$, we see that $c=0$. This completes the proof of Theorem 57 in Case 2. ∎
Proposition 58.
Suppose, in Theorem 57, that the $\Sigma_{g}$ converge to the plane $\{z=0\}$ with multiplicity $3$. Then all sufficiently large $g$, $\Sigma_{g}$ has exactly $3$ ends.
Proof.
Let $R>r$, where $C(r)$ is the circle away from which the convergence $\Sigma_{g}\to\Sigma$ is smooth. In the proof of Theorem 57, we showed that for large $g$, $\Sigma_{g}\cap B(0,R)$ is connected, has genus $2g$, and three boundary components.
Since $\Sigma_{g}$ is also connected and has genus $2g$, it it follows (by elementary topology) that $\Sigma_{g}\setminus B(0,R)$ has three components, $E_{g}^{1}$, $E_{g}^{2}$, and $E_{g}^{3}$. Note that (by hypothesis) $E_{g}^{i}$ converges smoothly (with multiplicity one) to $\{z=0\}\setminus B(0,R)$. It follows (see Lemma 60) that (for large $g$), each $E_{g}^{i}$ is topologically an annulus. ∎
The following is well-known, but we include a proof for the reader’s convenience.
Lemma 59.
Suppose $\Sigma_{i}$ are shrinkers that converge weakly to a multiplicity $1$ plane $\Sigma$. Then $\Sigma_{i}$ is a multiplicity $1$ plane for all sufficiently large $i$.
Proof.
The flows
(*) $t\in(-\infty,0]\mapsto|t|^{1/2}\Sigma_{i}$
converge weakly to the flow
$t\in(-\infty,0]\mapsto|t|^{1/2}\Sigma=\Sigma.$
The latter flow is smooth and multiplicity $1$, so the convergence is smooth on compact subsets of spacetime (by the Local Regularity Theorem, Theorem 7). In particular, for all sufficiently large $i$, the flow (*9) is smooth near the spacetime origin.
But, for each $i$, (trivially) the curvature of (*9) blows up at the spacetime origin unless $\Sigma_{i}$ is a plane. ∎
Lemma 60 (Ends Lemma).
Suppose $0<R<\infty$. Suppose that $C$ is a cone that is smooth except at the origin, and suppose that $M$ is a smooth shrinker with boundary in $B(0,R)$ and that $M$ is smoothly asymptotic (with multiplicity one) to $C$. Suppose for each $i$ that $M_{i}$ is a smooth shrinker with boundary in $B(0,R)$. Suppose that $M_{i}$ converges smoothly (and with multiplicity one) to $M$ on compact subsets of $\mathbf{R}^{3}\setminus\overline{B(0,R)}$. Then the convergence $M_{i}\to M$ is uniform at infinity in the following sense: if $\lambda_{i}\to 0$, then $\lambda_{i}\Sigma_{i}$ converges smoothly (with multiplicity $1$) on compact subsets of $\mathbf{R}^{3}\setminus\{0\}$ to $C\setminus\{0\}$.
Proof.
Consider the following open subset of spacetime:
$\Omega=\{(x,y,z,t):x^{2}+y^{2}\neq 0\}.$
Consider also the following mean curvature flows (with moving boundary):
$t\in(-\infty,0]\mapsto M_{i}(t)=|t|^{1/2}M_{i},$
$t\in(-\infty,0]\mapsto M(t)=|t|^{1/2}M.$
Note that the flow $M(\cdot)$ is smooth except at the spacetime origin.
By hypothesis, the flows $M_{i}(\cdot)$ converge smoothly to the flows $M(\cdot)$ on compact (spacetime) subsets of $\mathbf{R}^{3}\times(-\infty,0)$.
Since the flow $M(\cdot)$ is smooth away from the spacetime origin, it follows (by the Local Regularity Theorem, Theorem 7) that the convergence $M_{i}(\cdot)$ to $M(\cdot)$ is smooth in a spacetime neighborhood of each point $(p,0)$ with $p\neq O$. The result follows immediately, since
$\lambda_{i}M_{i}=M_{i}(-\lambda_{i}^{2}).$
10. Connectivity
Lemma 61.
Given $\Lambda<\infty$, there exist $\epsilon>0$ and $R<\infty$ with the following property. If $\Sigma$ is an embedded shrinker in $\mathbf{R}^{3}$ with $\operatorname{entropy}(\Sigma)\leq\Lambda$, and if $0<r\leq\epsilon$, then
$\Sigma\setminus B(0,r)$
is connected, and if $r>R$, then
$\Sigma\cap B(0,r)$
is connected.
Proof.
Suppose there is no such $\epsilon$. Then there is a sequence a sequence $\Sigma_{n}$ of embedded shrinkers and a sequence $r_{n}\to 0$ such that
$\operatorname{entropy}(\Sigma_{n})\leq\Lambda$
and such that $\Sigma_{n}\setminus B(0,r_{n})$ contains two connected components $C_{n}$ and $C_{n}^{\prime}$. (It may have more components.) Let $\Sigma_{n}^{\prime}$ be a surface in $\mathbf{R}^{3}\setminus B(0,r_{n})$ of least shrinker area that separates $C_{n}$ and $C_{n}^{\prime}$.
Then $\Sigma_{n}^{\prime}$ is a surface whose boundary is in $\partial B(0,r_{n})$. It is smoothly embedded away from the boundary, and its shrinker area is at most
$\operatorname{area}_{s}(\Sigma_{n})=\operatorname{entropy}(\Sigma_{n})\leq\Lambda.$
Now the $\Sigma_{n}^{\prime}$ are stable, so (after passing to a subsequence, they converge smoothly in $\mathbf{R}^{3}\setminus\{0\}$ to a stable shrinker $\Sigma^{\prime}$ (possibly with multiplicity).
By [gulliver-lawson], $\Sigma^{\prime}$ is smoothly embedded everywhere. But then $\Sigma^{\prime}$ is unstable, a contradiction.
The proof for $R$ is essentially the same. ∎
11. Genus of Shrinkers
Chodosh, Choi, and Schulze proved that that all shrinkers at a spacetime point have the same genus [cms]*Proposition H.8. Here, we give a different proof that is perhaps somewhat simpler.
Lemma 62.
Suppose that $T\leq\operatorname{T_{\textnormal{fat}}}$ and that $p\in M(T)$. Let $\mathscr{S}$ be the collection of all shrinkers at $(p,T)$, i.e., the the set of all subsequential limits of
$(T-t)^{1/2}(M(t)-p)$
as $t\uparrow T$.
There is an $R>0$ with the following property. If $\Sigma\in\mathscr{S}$, then the function
$q\in\Sigma\setminus\overline{B(0,R)}\mapsto|q|$
is a Morse function, and the only critical points are local maxima.
(In other words, at each critical point $q$, both principal curvatures with respect to the inward pointing normal are $>1/|q|$.)
Proof.
Suppose not. Then there exist $\Sigma_{i}\in\mathscr{S}$, and critical points $q_{i}\in\Sigma_{i}$ such that $R_{i}:=|q_{i}|\to\infty$ and such that one of the principal curvatures at $q_{i}$ (with respect to the inward pointing normal) is $\leq 1/|q_{i}|$.
Consider the tangent flow:
(*) $t\in(-\infty,0)\mapsto|t|^{1/2}\Sigma_{i}.$
The mean curvature of $\Sigma_{i}$ at $q_{i}$ is $-q_{i}/2$. Thus the mean curvature of $R_{i}\Sigma_{i}$ at $R_{i}q_{i}$ is $-q_{i}/(2|q_{i}|)$.
Now translate the flow (*11) in spacetime by $(-R_{i}q_{i},R_{i}^{2})$ to get
$t\in(-\infty,R_{i}^{2})\mapsto$ $|R_{i}^{2}-t|^{1/2}\Sigma_{i}-R_{i}q_{i}$
$=|1-(t/R_{i}^{2})|^{1/2}R_{i}\Sigma_{i}-R_{i}q_{i}$
$=|(1-(t/R_{i}^{2})|^{1/2}R_{i}(\Sigma_{i}-q_{i})+|(1-(t/R_{i}^{2})|^{1/2}R_{i}q_{i}-R_{i}q_{i}$
$=|(1-(t/R_{i}^{2})|^{1/2}R_{i}(\Sigma_{i}-q_{i})-(t+o(t))\frac{q_{i}}{|q_{i}|}$
By [bamler-kleiner], this converges (perhaps after passing to a subsequence) to a flow that is regular at almost all times, and the convergence is smooth for almost all $t$. Thus we see that $R_{i}(\Sigma_{i}-q_{i})$ converges smoothly to a limit $\Sigma^{\prime}$ such that
(43) $t\in\mathbf{R}\mapsto\Sigma^{\prime}-tq$
is mcf, i.e., such that $\Sigma^{\prime}$ is a translator with velocity $-q$. Thus the mean curvature at each point of $\Sigma^{\prime}$ is $-q^{\perp}$.
Let
$t\in(-\infty,0)\mapsto|t|^{-1/2}\Sigma^{\prime\prime}$
be tangent flow at $\infty$ to the flow (43). Thus, there are $t_{n}\to-\infty$ such that
$\Sigma^{\prime}_{n}:=|t_{n}|^{-1/2}\Sigma^{\prime}(t_{n})=|t_{n}|^{-1/2}(\Sigma^{\prime}-t_{n}q)$
converges to $\Sigma^{\prime\prime}$. By [bamler-kleiner], the convergence is smooth with multiplicity one. Note that the mean curvature vector $H_{n}$ at each point of $\Sigma^{\prime}_{n}$ is $(-t_{n}q)^{\perp}$. Thus
$|H_{n}|=|\nu\cdot H_{n}|=|t_{n}|\,|q^{\perp}|,$
or
$|q^{\perp}|=|t_{n}|^{-1}|H_{n}|.$
Letting $n\to\infty$, we see that all points $q$ of $\Sigma^{\prime\prime}$, $|q^{\perp}|=0$. That is, $\Sigma^{\prime}$ is translation-invariant in the $q$ direction. Thus $\Sigma^{\prime}$ is the shrinking cylinder, and therefore $\Sigma^{\prime}$ is the bowl soliton [chh].
But that is a contradiction. ∎
Corollary 63.
If $\Sigma\in\mathscr{S}$, if $r>R$, and if $\partial B(0,r)$ intersects $\Sigma$ transversely, then $\operatorname{genus}(\Sigma\cap B(0,r))=\operatorname{genus}(\Sigma)$.
Proof.
By the lemma and by standard Morse Theory, each component of $M\setminus B(0,r)$ is diffeomorphic to a closed disk or to $S^{1}\times[0,1)$. (Here $B(0,r)$ is the open ball.) ∎
Theorem 64.
All shrinkers at $(p,T)$ have the same genus.
Proof.
The set $\mathscr{S}$ is connected (with respect to smooth convergence), so it suffices to show that if $\Sigma_{i}\in\mathscr{S}$ converges smoothly to $\Sigma\in\mathscr{S}$, then
$\operatorname{genus}(\Sigma_{i})=\operatorname{genus}(\Sigma)$
for all sufficiently large $i$. Let $R$ be as in the Lemma, and let $r>R$ be such that $\partial B(0,r)$ intersects $\Sigma$ transversely.
Thus $\partial B(0,r)$ intersects $\Sigma_{i}$ transversely for all sufficiently large $i$, say $i\geq I$.
Also, by smooth convergence,
$\operatorname{genus}(M_{i}\cap B(0,r))=\operatorname{genus}(M\cap B(0,r))$
for all sufficently large $i$.
The result follows immediately, since, by Corollary 63, $\operatorname{genus}(M\cap B(0,r))=\operatorname{genus}(M)$ and $\operatorname{genus}(\Sigma_{i}\cap B(0,r))=\operatorname{genus}(M_{i})$ for all $i\geq I$. ∎
Conflict of interest statement: On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statement: This manuscript has no associated data.

### v3

@@
 F.
-Martín’s research is partially supported by the framework of IMAG-María de Maeztu grant CEX2020-001105-M funded by MCIN/AEI/ 10.13039/50110001103.
+Martín’s research is partially supported by the framework of IMAG-María de Maeztu grant CEX2020-001105-M and by the grant PID2024-156031NB-I00 both funded by MCIU/AEI/ 10.13039/501100011033.
 Some of this work was carried out while the authors were visitors at the Simons Laufer Mathematical Sciences Institute (SLMath, formerly MSRI) in Berkeley, CA during the Fall 2024 semester, in a program supported by National Science Foundation Grant No.
@@
 Their examples presumably coincide with the Angenent Tori, whose existence was established in [angenent] by a shooting method.
-Figure 1. A cross section of a numerical simulation of one of the shrinkers constructed in Section 7. Figure courtesy of Mario Schulz.
+Figure 1. A numerical simulation of the shrinkers constructed in Section 7, for genus $g=1,\ldots,6$. Figures courtesy of Francisco Torralbo (UGR).
 2. Preliminaries
 […]
-In this paper, “mean curvature flow” means “unit-regular Brakke flow”.
-Roughly speaking, a unit-regular Brakke flow is an integral Brakke flow that is smooth (without sudden vanishing) in a spacetime neighborhood of each point of Gauss density $1$.
-(See [white-boundary]*Definition 11.1 for the precise definition.) Any weak limit of unit-regular flows is also unit-regular [white-local].
-If $K$ is a closed subset of $\mathbf{R}^{3}$, or, more generally, of a smooth Riemannian $3$-manifold $N$, we let $K(t)$ denote the result of flowing $K$ for time $t$ by the level set flow.
-Let $\operatorname{T_{\textnormal{fat}}}:=\operatorname{T_{\textnormal{fat}}}(K)$ be the fattening time, i.e, the infimum of times $t$ such that $K(t)$ has nonempty interior.
-Definition 1.
-If $M$ is smooth, a regular time for $M$ is a time $t\in[0,\operatorname{T_{\textnormal{fat}}}(M)]$ such that $M(\tau)$ is smooth for all $\tau$ in some neighborhood $[0,\operatorname{T_{\textnormal{fat}}}]\cap(t-\epsilon,t+\epsilon)$ of $t$.
-If $M$ is smoothly embedded and compact, then 00 is a regular time. By [white-genus], if $t_{1}<t_{2}<\operatorname{T_{\textnormal{fat}}}(M)$ are regular times, then
+If $K$ is a closed subset of $\mathbf{R}^{3}$, or, more generally, of a smooth Riemannian $3$-manifold NN, we let $K(t)$ denote the result of flowing $K$ for time $t$ by the level set flow.
+Let $\operatorname{T_{\textnormal{fat}}}(K)$ be the fattening time, i.e, the infimum of times $t$ such that $K(t)$ has nonempty interior.
+If (as we are assuming in this section) $M$ is compact and smoothly embedded, then $\operatorname{T_{\textnormal{fat}}}(M)$ is greater than or equal to the first time when classical mean curvature starting from $M$ develops a singularity.
+In particular, $\operatorname{T_{\textnormal{fat}}}(M)>0$.
+By [bamler-kleiner], $M(t)$ is a compact, smoothly embedded surface for almost every $t\in[0,\operatorname{T_{\textnormal{fat}}}]$, and the set of such times is relatively open in $0,\operatorname{T_{\textnormal{fat}}}]$.
+Furthermore,
+($\star$) $t\in[0,\operatorname{T_{\textnormal{fat}}}]\mapsto\mu(t):=\mathscr{H}^{2}\llcorner M(t)$
+is a unit-regular Brakke flow.
+(Roughly speaking, a unit-regular Brakke flow is an integral Brakke flow that is smooth (without sudden vanishing) in a spacetime neighborhood of each point of Gauss density $1$.
+See [white-boundary]*Definition 11.1????
+for the precise definition.
+Weak limits of unit-regular flows are also unit regular [white-local].)
+By slight abuse of terminology, we will refer to
+$t\in[0,\operatorname{T_{\textnormal{fat}}}]\mapsto M(t)$
+as a Brakke flow, when we mean the Brakke flow ($\star$2).
+By [bamler-kleiner], if $0\leq t_{1}<t_{2}\leq\operatorname{T_{\textnormal{fat}}}(M)$ are times at which $M(t)$ is smooth, then
 (1) $\operatorname{genus}(M(t_{2}))\leq\operatorname{genus}(M(t_{1})).$
-By [bamler-kleiner], almost every time in $[0,\operatorname{T_{\textnormal{fat}}}(M)]$ is regular, and all tangent flows at times $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ are given by smoothly embedded shrinkers of multiplicity $1$.
+By [bamler-kleiner], all tangent flows at times $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ are given by smoothly embedded shrinkers of multiplicity $1$.
 Thus if $p\in N$ and $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, then every sequence $t_{i}\uparrow T$ has a subsequence such that
 (2) $(T-t_{i})^{-1/2}(M(t_{i})-p)$
-converges smoothly and with multiplicity $1$ on compact subsets to a shrinker $\Sigma$. If $\tau_{i}/t_{i}$ is bounded away from 00 and away from $\infty$, then
+converges smoothly and with multiplicity $1$ on compact subsets to a shrinker $\Sigma$. If $\tau_{i}/t_{i}$ is bounded away from 0 and away from $\infty$, then
 $(T-\tau_{i})^{-1/2}(M(\tau_{i})-p)$
@@
 We expect there to be only one shrinker at each spacetime point, but that is not known in all cases.
-(See [chodosh-schulze] for cases in which uniqueness is known.) However, all the shrinkers at $(p,T)$ have the same genus (by Theorem 64 in the appendix), so the following definition makes sense.
+(See [chodosh-schulze] for some cases in which uniqueness is known.) However, all the shrinkers at $(p,T)$ have the same genus (by Theorem 68 in the appendix), so the following definition makes sense.
 Definition 2.
 The genus of a spacetime point $(p,t)$ with $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ is the number $g$ such that each shrinker at $(p,t)$ has genus $g$.
-At a singular point of genus 00, the shrinker is either a sphere or a cylinder [brendle].
-(Earlier, Kleene and Møller proved the result assuming an axis of rotational symmetry [kleene-moller].) Furthermore, the shrinker is unique, by [huisken] (in the spherical case) and [colding-m] (in the cylindrical case).
-See [zhu] for a simpler proof in the cylindrical case.
+At a singular point of genus 0, the shrinker is either a sphere or a cylinder [brendle].
+(Earlier, Kleene and Møller proved the result assuming an axis of rotational symmetry [kleene-moller].) Furthermore, the shrinker is unique by [colding-m].
+(See also [zhu].)
 The following result is trivial, but important.
@@
 Suppose $T\leq\operatorname{T_{\textnormal{fat}}}$. Then the sum of the genuses of the singularities at time $T$ is less than or equal to the genus of $M$.
-If $\operatorname{T_{\textnormal{fat}}}<\infty$, then there is a spacetime singular point $(p,\operatorname{T_{\textnormal{fat}}})$ of positive genus.
+Theorem 5.
+If $\operatorname{T_{\textnormal{fat}}}<\infty$, then there is shrinker at time $\operatorname{T_{\textnormal{fat}}}$ with genus $\geq 1$.
+Proof.
+By [chh]*Theorem 1.9, there is a shrinker $\Sigma$ at time $\operatorname{T_{\textnormal{disc}}}\leq\operatorname{T_{\textnormal{fat}}}$ that is not a plane, a shrinking sphere, or a cylinder.
+(See [hw]*Definition 2 or [chh]*p.
+227 for the definition of the discrepancy time $\operatorname{T_{\textnormal{disc}}}=\operatorname{T_{\textnormal{disc}}}(M)$.) Since planes, spheres, and cylinders are the only smooth genus 0 shrinkers, $\Sigma$ must have genus $\geq 1$.
+Finally, by [bamler-kleiner]*Theorem 1.8, $\operatorname{T_{\textnormal{disc}}}=\operatorname{T_{\textnormal{fat}}}$.
+∎
 Definition 6.
-We let $T_{\textnormal{pos}}(M)$ be the infimum of times $t$ at which there is a singularity of positive genus.
-The set of singular spacetime points of genus $>0$ is a closed subset of spacetime.
-Thus if $T_{\textnormal{pos}}(M)<\infty$, then there is spacetime singularity $(p,T_{\textnormal{pos}}(M))$ of positive genus.
-(In [ccms]*§11, $T_{\textnormal{pos}}(M)$ is written $T_{\textnormal{gen}}(M)$.)
+The time $T_{\textnormal{pos}}(M)$ is the infimum of times $t\leq\operatorname{T_{\textnormal{fat}}}$ at which there is a singularity of positive genus.
+In some papers, $T_{\textnormal{pos}}(M)$ is written $T_{\textnormal{gen}}(M)$.
+Theorem 7.
+If $T_{\textnormal{pos}}<\infty$, then there is a shrinker at time $T_{\textnormal{pos}}$ of genus $\geq 1$.
+Proof.
+Note that, by definition,
+$T_{\textnormal{pos}}(M)\leq\operatorname{T_{\textnormal{fat}}}(M).$
+If $T_{\textnormal{pos}}=\operatorname{T_{\textnormal{fat}}}$, then there is a positive-genus shrinker at time $T_{\textnormal{pos}}$ by Theorem 5.
+If $T_{\textnormal{pos}}<\operatorname{T_{\textnormal{fat}}}$, then there is a positive-genus shrinker at time $T_{\textnormal{pos}}$ because the set of spacetime points at which the genus is positive is a closed subset of spacetime.
+See, for example, [white-genus]*Theorem 25.
+∎
+Remark.
+In this paper, we only need to flow up to time $T_{\textnormal{pos}}$.
 Theorem 8 (Local Regularity Theorem).
@@
 Theorem 9 (Initial Regularity Theorem).
-Suppose $t\in[0,T)\mapsto M(t)$ is a mean curvature flow in an open set $U$ such that $M(0)$ is a smooth, propertly embedded, multiplicity $1$ submanifold of $U$.
+Suppose $t\in[0,T)\mapsto M(t)$ is a mean curvature flow in an open set $U$ such that $M(0)$ is a smooth, properly embedded, multiplicity $1$ submanifold of $U$.
 Let $K$ be a compact region in $U$ with smooth boundary transverse to $M(0)$.
@@
 For notational simplicity, we give the proof in Euclidean space. Let
-$K^{\prime}$ =$=\{\operatorname{dist}(\cdot,K)\leq\delta\},$
+$K^{\prime}$ $=\{\operatorname{dist}(\cdot,K)\leq\delta\},$
 where $\delta>0$ is small enough that $K^{\prime}$ is compact subset of $U$ and that $\partial K^{\prime}$ is transverse to $M(0)$.
 Claim 1.
-Suppose $p_{i}$ converges to $p\in M(0)$ and that $t_{i}>0$ converges to 00. Let $q_{i}$ be the point in $M(0)$ closest to $p_{i}$. Let
+Suppose $p_{i}$ converges to $p\in M(0)$ and that $t_{i}>0$ converges to 0. Let $q_{i}$ be the point in $M(0)$ closest to $p_{i}$. Let
 $r_{i}=\max\{|p_{i}-q_{i}|,t_{i}^{1/2}\}.$
@@
 It follows from Claim 1 and from [sw-local]*Corollary A.3 that there is an $\epsilon>0$ such that
-(*) The flow $t\in[0,2\epsilon]\mapsto M(t)\cap K^{\prime}$ is smooth.
-Assertion (1) follows immediately. Assertion (2) follows from (*2) and the Local Regularity Theorem (Theorem 7).
+($\dagger$) The flow $t\in[0,2\epsilon]\mapsto M(t)\cap K^{\prime}$ is smooth.
+Assertion (1) follows immediately. Assertion (2) follows from ($\dagger$2) and the Local Regularity Theorem (Theorem 8).
 Now let
@@
 Equivalently, $\operatorname{Reg}(t)$ is the set of points $p$ such that the Gauss Density at $(p,t)$ is $1$.
-We let $\operatorname{Sing}(t)$ be the set of $p$ such that the Gauss Density of at $(p,t)$ is $>1$.
+We let $\operatorname{Sing}(t)$ be the set of $p$ such that the Gauss Density of at $(p,t)$ is >1>1.
 Definition 10.
@@
   3. (3)
-    $\Sigma$ occurs as a shrinker in a mean curvature flow $M(\cdot)$ for which $M(0)$ is a compact, smoothly embedded surface in $Q^{c}$.
+    $\Sigma$ occurs as a shrinker in a mean curvature flow $M(\cdot)$ in $\mathbf{R}^{3}$ for which $M(0)$ is a compact, smoothly embedded surface in $Q^{c}$.
 We do not know whether any of the shrinkers in Theorem 12 are compact.
@@
 The proof of this lemma is an immediate consequence of Theorem 5.2 in [white-95]. ∎
+Definition 14.
 Let $\mathscr{A}$ be the class of compact, smoothly embedded surfaces $M$ in $Q^{c}$ such that
@@
 Theorem 15.
-If $M\in\mathscr{A}$ and if $t<\min\{T_{Q},\operatorname{T_{\textnormal{fat}}}\}$ is a regular time of the flow $M(\cdot)$, then $M(t)\in\mathscr{A}$.
+If $M\in\mathscr{A}$ and if $t<\min\{T_{Q},T_{\textnormal{pos}}\}$ is a regular time of the flow $M(\cdot)$, then $M(t)\in\mathscr{A}$.
+Later we will see that $T_{Q}=T_{\textnormal{pos}}$.
 Proof.
-The theorem is a trivial consequence of the facts that (1) level set flow preserves symmetries, and (2) for any compact smooth embedded surface, $\operatorname{genus}(M(t))$ is a decreasing function of regular times in $[0,\operatorname{T_{\textnormal{fat}}}]$.
+The theorem is a trivial consequence of the facts that (1) level set flow preserves symmetries, and (2) for any compact smooth embedded surface, $\operatorname{genus}(M(t))$ is a decreasing function of regular times in $0,T_{\textnormal{pos}}]$.
 (See (1).) ∎
 Theorem 16.
-If $M\in\mathscr{A}$, then there is a connected component $\Sigma$ of $M$ such that the convex hull of $\Sigma$ contains points of $Q$,
+If $M\in\mathscr{A}$, then there is a connected component $\Sigma$ of $M$ such that the convex hull of $\Sigma$ contains points of $Q$.
 Proof.
@@
 Theorem 17.
-Let $M$ be a $G$-invariant surface properly embedded in $\mathbf{R}^{3}$, disjoint from $Q$, and of genus $\leq g$.
+Let $M$ be a compact, $G$-invariant surface embedded in $\mathbf{R}^{3}$, disjoint from $Q$, and of genus $\leq g$.
 Suppose $M$ has a connected component $\Sigma$ whose convex hull contains points of $Q$.
-Then $\Sigma$ has genus $g$, and if $\Omega$ is a simply connected open subset of $Q^{c}$, then $\operatorname{genus}(M\cap\Omega)=0$.
-Proof.
+Then $\Sigma$ is a $G$-invariant, genus $g$ surface, and if $\Omega$ is a simply connected open subset of $Q^{c}$, then $\operatorname{genus}(M\cap\Omega)=0$.
+Proof of Theorem 17.
 […]
@@
 Claim 3.
-$\Psi$ contains points in $F_{12}$ and points in $F_{13}$
+$\Psi$ contains points in $F_{12}$ and points in $F_{13}$.
 Suppose not.
-Then (by relabelling) we may assume that it contains no points in the face $F_{12}$.
+Then (by relabelling) we may assume that $\Psi$ contains no points in the face $F_{12}$.
 It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{3}$, that $\Sigma$ lies in $\operatorname{interior}(K_{3})$, a convex open set disjoint from $Q$.
-But that contradicts the hypothesis on $\Sigma$.
+But that contradicts the hypothesis that the convex hull of $\Sigma$ contains points of $Q$.
 Thus the claim is proved.
-Let $M^{\prime}$ be the union of $\Sigma$ and its images under $G$.
-Let $U^{\prime}$ be the region bounded by $M^{\prime}$ that is disjoint from $Q$.
-(Thus, a point in $(M^{\prime})^{c}$ is in $U^{\prime}$ if and only if there is a path from that point to a point in $Q$ that is transverse to $M^{\prime}$ and that intersects $M^{\prime}$ in an odd number of points.) Let $K=M^{\prime}\cup U^{\prime}$.
-Thus $K$ is smooth $3$-manifold-with-boundary, the boundary being $M^{\prime}$.
 Let $\gamma$ be an embedded curve in $\Psi$ such that one of its endpoints is in $F_{12}$ and the other endpoint is in $F_{13}$.
-Iterated reflections in the planes of $\mathscr{P}$ containing $E_{1}$ produces a simple closed curve in $M^{\prime}\cap K_{1}$ that winds once around the line containing $E_{1}$.
-Of course, there is such a curve for each ray of $Q$.
-Let $\mathscr{G}$ be the set of those $g+1$ curves.
-Note that any $g$ of those $g+1$ curves form a basis for the vector space $H_{1}(Q^{c};\mathbf{Z}_{2})$.
-Thus if $\iota$ is the inclusion of $K$ into $Q^{c}$, then the map
-(2) $\iota_{\#}:H_{1}(K;\mathbf{Z}_{2})\to H_{1}(Q^{c};\mathbf{Z}_{2}).$
-is surjective.
-Now
-$g\geq\operatorname{genus}(M^{\prime})=\dim H_{1}(K;\mathbf{Z}_{2})\geq\dim H_{1}(Q^{c};\mathbf{Z}_{2})=g.$
-Thus all the inequalities are equalities and the map (2) is an isomorphism.
-Now the $g+1$ curves of $\mathscr{G}$ taken together are homologically trivial (mod $2$) in $Q^{c}$ and thus also in $K$.
-Thus they bound a surface in $K$.
-That surface must be connected since no subcollection of those curves bounds a surface in $Q^{c}$.
-Hence $K$ is connected.
-Therefore $M^{\prime}=\partial K$ is connected, so $\Sigma=M^{\prime}$.
-It remains to show that if $W$ is a simply connected open subset of $Q^{c}$, then $M\cap W$ has genus 00.
-Since $M$ and $\Sigma$ have genus $g$, $M\setminus\Sigma$ has genus 00, and therefore $(M\setminus\Sigma)\cap W$ has genus 00.
-Thus it suffices to prove that $\Sigma\cap W$ has genus 00.
-Suppose to the contrary that $\Sigma\cap W$ has positive genus.
-Then, by elementary topology (see Lemma 16 below), $\Sigma\cap W$ contains a simple closed curve that does not bound a surface in $K$.
-Therefore it does not bound a surface in $Q^{c}$ since the map (2) is an isomorphism, and therefore it does not bound a surface in $W$.
-But that contradicts the assumption that $W$ is simply connected.
+Note that $\gamma$ is disjoint from $E_{1}$ since $M\subset Q^{c}$.
+Iterated reflections in the planes of $\mathscr{P}$ containing $E_{1}$ produces a simple closed curve $C$ in $\Sigma^{*}\cap\operatorname{interior}(K_{1})$ that winds once around $E_{1}$ in $K_{1}\setminus E_{1}$.
+In particular,
+[…]
+Since $C\subset\Sigma$ is homologically nontrivial in $Q^{c}$, it is homologically nontrivial in $\Sigma$, and thus
+$\operatorname{genus}(\Sigma)\geq 1.$
+Claim 4.
+$\Psi$ contains points in $F_{23}$.
+Suppose not.
+It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{1}$, that $\Sigma$ lies in $K_{1}$.
+For each ray $E_{i}^{\prime}$ of $Q$, $M$ contains a copy of $\Sigma$.
+There are $g+1$ such rays, so $M$ contains $g+1$ such copies.
+Since each copy has positive genus, $M$ has genus $\geq g+1$, a contradiction.
+Thus the claim is proved.
+Since $\Sigma$ contains points in each $F_{ij}$, it follows that $\Sigma$ is invariant under reflection in each of the corresponding three planes.
+Those reflections generate $G$, so $\Sigma$ is invariant under $G$.
+Now $E_{1}$ is one of the rays of $Q$.
+Let $E_{1}^{\prime}$ be another ray of $Q$, and let $\sigma\in G$ map $E_{1}$ to $E_{1}^{\prime}$.
+Then $C^{\prime}:=\sigma(C)$ is a simple closed curve in $\Sigma\cap\operatorname{interior}(K_{1}^{\prime})$, where $K_{1}^{\prime}=\sigma(K_{1})$, and $C^{\prime}$ winds once around $E_{1}^{\prime}$ in $K_{1}^{\prime}\setminus E_{1}^{\prime}$.
+Note that $C$ and $C^{\prime}$ are disjoint since $\operatorname{interior}(K_{1})$ and $\operatorname{interior}(K_{1}^{\prime})$ are disjoint.
+[…]
+Since $\Sigma\setminus\cup\mathscr{G}^{\prime}$ is connected,
+$0\leq\operatorname{genus}(\Sigma\setminus\cup\mathscr{G}^{\prime})=\operatorname{genus}(\Sigma)-g\leq\operatorname{genus}(M)-g\leq 0.$
+Thus
+(3) $\operatorname{genus}(\Sigma)=\operatorname{genus}(M)=g,$
+and
+$\operatorname{genus}(\Sigma\setminus\mathscr{G}^{\prime})=0.$
+By (3), $M\setminus\Sigma$ is a union of spheres, so
+(4) $\operatorname{genus}(M\setminus\mathscr{G}^{\prime})=\operatorname{genus}(\Sigma\setminus\mathscr{G}^{\prime})=0.$
+We have shown that $\Sigma$ is a $G$-invariant surface of genus $g$, and also that the genus of $M$ is equal to $g$. It remains to prove the last assertion of the theorem.
+Let $U$ be the region bounded by $M$ that is disjoint from $Q$.
+(Thus, a point in $M^{c}$ is in $U$ if and only if there is a path from that point to a point in $Q$ that is transverse to $M$ and that intersects $M$ in an odd number of points.) Let $K=\overline{U}=U\cup M$.
+Thus $K$ is smooth 3-manifold-with-boundary, the boundary being $M$.
+Since $K$ is a smooth, compact $3$-manifold-with-boundary in $\mathbf{R}^{3}$, we can use [hatcher]*Lemma 3.5 and Poincaré duality to get that
+$\dim H_{1}(K)\geq\frac{1}{2}\dim H_{1}(M)=g.$
+On the other hand, one can easily get the reverse inequality using the Mayer-Vietoris Theorem. Hence, we deduce that:
+$\dim H_{1}(K)=\frac{1}{2}\dim H_{1}(M)=g.$
+Now consider the inclusions
+$\mathscr{G}^{\prime}\subset K\subset Q^{c}.$
+Since $\mathscr{G}^{\prime}$ is a basis for $H_{1}(Q^{c})$, and since $H_{1}(\mathscr{G})$, $H_{1}(K)$, and $H_{1}(Q^{c})$ each have dimension $g$, the inclusion $\iota$ of $K$ into $Q^{c}$ induces an isomorphism on $H_{1}$:
+(5) $\iota_{\#}:H_{1}(K)\xrightarrow{\cong}H_{1}(Q^{c}).$
+It remains to show that if $\Omega$ is a simply connected open subset of $Q^{c}$, then $M\cap\Omega$ has genus 0.
+Suppose, to the contrary, that $M\cap\Omega$ has positive genus.
+Then, by elementary topology (see Lemma 19 below), $M\cap\Omega$ contains a simple closed curve $\alpha$ that does not bound a surface in $K$.
+Therefore $\alpha$ does not bound a surface in $Q^{c}$ since the map (5) is an isomorphism, and therefore $\alpha$ does not bound a surface in $\Omega$.
+But that contradicts the assumption that $\Omega$ is simply connected.
 ∎
+The following theorem gives additional information about the surface in Theorem 17.
+Theorem 18.
+Let $M$ be as in Theorem 17 and let $R>0$. Suppose the convex hull of a connected component of $M\cap B(0,R)$ contains points in $Q$. Then for every $r>R$, $M\cap B(0,r)$ has genus 0 or genus $g$.
+Proof.
+It suffices to prove Theorem 18 in the case when $\partial B(0,r)$ intersects $M$ transversely.
+Case 1: The convex hull of one of the components of $M\cap\partial B(0,r)$ contains a point of $Q$.
+Let $C$ be such a component, and let $\mathscr{G}$ be the union of its images under the group $G$.
+Exactly as in the proof of Theorem 17 (see (4)), $M\setminus\cup\mathscr{G}$ has genus 0.
+Thus $M\cap B(0,r)$ has genus 0.
+Case 2: The convex hull of each component of $M\cap\partial B(0,r)$ is disjoint from $Q$. Then each such component bounds an embedded disk in
+$\mathbf{R}^{3}\setminus(B(0,r)\cup Q).$
+We can choose those disks so that the union $M^{\prime}$ of those disks and $M\cap B(0,r)$ is a smoothly embedded, $G$-invariant surface. By Theorem 17, $M^{\prime}$ has genus $g$. Thus
+$\operatorname{genus}(M\cap B(0,r))=\operatorname{genus}(M^{\prime})=g.$
 Lemma 19.
-Let $K$ be a closed region in $\mathbf{R}^{3}$ with smooth boundary $\Sigma$.
-Let $W$ be an open subset of $\mathbf{R}^{3}$ such that $\Sigma\cap W$ has positive genus.
-Then $\Sigma\cap W$ contains a simple closed curve that is homologically nontrivial in $K$.
+Suppose that $K$ is a smooth, $3$-dimensional manifold-with-boundary.
+Let $M=\partial K$.
+Suppose that $W$ is a relatively open subset of $K$ such that $M\cap W$ has positive genus.
+Then $M\cap W$ contains a simple closed curve that is homologically nontrivial in $K$.
 Proof.
-Let $\alpha_{1}$ and $\alpha_{2}$ be a pair of simple closed curves in $\Sigma\cap W$ that intersect transversely and at exactly one point.
+Let $\alpha_{1}$ and $\alpha_{2}$ be a pair of simple closed curves in $M\cap W$ that intersect transversely and at exactly one point.
 We claim that at most one of $\alpha_{1}$ and $\alpha_{2}$ can be homologically trivial in $K$.
 For otherwise, there would be smooth, compact embedded surfaces $\sigma_{i}$ in $K$ such that $\partial\sigma_{i}=\alpha_{i}$.
-We can choose the $\sigma_{i}$ so that they intersect transversely and so that they are never tangent to $\Sigma$.
+We can choose the $\sigma_{i}$ so that they intersect transversely and so that they are never tangent to $M$.
 Thus $\sigma_{1}\cap\sigma_{2}$ is a finite collection of curves, and so they have an even number of endpoints.
@@
 Theorem 20.
-Suppose $M\in\mathscr{A}$. Then $T_{Q}\leq T_{\textnormal{pos}}$.
+Suppose $M\in\mathscr{A}$. (See Definition 14.) Then $T_{Q}\leq T_{\textnormal{pos}}$.
 Proof.
@@
 Let $W$ be an open ball in $Q^{c}$ centered at $p$.
-By Theorems 13, 14, and 15, $\operatorname{genus}(M(t)\cap W)=0$ for all regular times $t<T$.
-Thus $(p,T)$ must be a genus 00 singularity of the flow.
+By Theorems 15, 16, and 17, $\operatorname{genus}(M(t)\cap W)=0$ for all regular times $t<T$.
+Thus $(p,T)$ must be a genus 0 singularity of the flow.
 Thus $T<T_{\textnormal{pos}}$, since at time $T_{\textnormal{pos}}$, there must be a singularity of positive genus.
@@
 Theorem 21.
-Suppose that $M\in\mathscr{A}$ and that $T:=T_{Q}(M)<\infty$. Let $p\in M(T)\cap Q$ and let $\Sigma$ be a shrinker at the spacetime point $(p,T)$.
+Suppose that $M\in\mathscr{A}$, and, therefore, that $T:=T_{Q}(M)<\infty$. Let $p\in M(T)\cap Q$ and let $\Sigma$ be a shrinker at the spacetime point $(p,T)$.
   1. (1)
@@
 If $(p,T)$ were a regular point of the flow, then for all $t<T$ sufficiently close to $T$, there would be a unique point $p(t)$ in $M(t)$ closest to $p$.
-But, since $p(t)\notin Q$, the image of $p(t)$ under rotation by $\pi$ about the line $\{sp:s\in\mathbf{R}\}$ is another point in $M(t)$ closest to $p$, a contradiction.
+Now the group $G$ contains a nontrivial group of rotations about the line $\{sp:s\in\mathbf{R}\}$.
+Since $p(t)\notin Q$, the image of $p(t)$ under those rotations would be other points in $M(t)$ closest to $p$, a contradiction.
 Hence $(p,T)$ is a singular spacetime point of the flow.
-Let $\sigma$ be the genus of the singular point $(p,T_{Q})$. Let $S$ be the set consisting of the $g+1$ points in $Q$ at distance $|p|$ from 00, and let
+Let $\sigma$ be the genus of the singular point $(p,T_{Q})$. Let $S$ be the set consisting of the $g+1$ points in $Q$ at distance $|p|$ from 0, and let
 $S^{\prime}=\{(p,T_{Q}):p\in S\}.$
@@
 Proof in Case (2).
-Recall that $\Sigma$ is a smooth, properly embedded surface.
-If $0\in\Sigma$, then the oriented tangent plane to $\Sigma$ would be invariant under $G$.
-But there is no such plane.
-Hence $0\notin\Sigma$.
-If $q$ were a point in $Q\cap\Sigma$, then (as we have just shown) $q\neq 0$.
-The oriented tangent plane to $\Sigma$ at $q$ is invariant under the stabilizer subgroup $G_{q}$, which implies that $\Sigma$ intersects the ray $\mathbf{R}^{+}p$ orthogonally at $p$.
-But that is impossible since $\Sigma$ is a smooth limit of surfaces disjoint from $Q$.
-We have shown that $\Sigma\subset Q^{c}$.
-Also, it has genus $\leq g$.
-It is connected since any smooth embedded shrinker in $\mathbf{R}^{3}$ must be connected.
-Finally, by the convex hull property, 00 is in the convex hull of $\Sigma$.
-Thus $\Sigma$ has genus $g$ by Theorem 15.
+First, we show that $\Sigma$ is disjoint from $Q$. Note that there exist $t_{i}\uparrow T$ such that the surfaces
+$M_{i}^{\prime}:=(T-t_{i})^{-1/2}M(t_{i})$
+converge smoothly and with multiplicity one to $\Sigma$.
+Thus if $q\in\Sigma$, then (for large $i$) there is a unique point $q_{i}$ in $M_{i}^{\prime}$ closest to 0.
+But if $q$ were in $Q$, then, since $M_{i}^{\prime}$ is disjoint from $Q$, the symmetries of $M_{i}^{\prime}$ imply that there are multiple points in $M_{i}^{\prime}$ closest to $q$, a contradiction.
+Hence $q\notin Q$.
+Thus $Q$ and $\Sigma$ are disjoint.
+By Theorem 18, for each $i$ and each $r>0$,
+$M_{i}^{\prime}\cap B(0,r)$
+has genus 0 or genus $g$. Thus, by smooth convergence of $M_{i}^{\prime}$ to $\Sigma$, $\Sigma$ has genus 0 or genus $g$.
+If $\Sigma$ had genus 0, then, by Brendle’s Theorem [brendle].
+it would be a plane through the origin, a cylinder whose axis contains the origin, or a sphere centered at the origin, all of which intersect $Q$.
+But $\Sigma$ is disjoint from $Q$.
+Thus, $\Sigma$ has genus $g$.
 ∎
 Theorem 22.
-Suppose $M_{n}$ are compact, smoothly embedded surfaces in $Q^{c}$ such that $T_{Q}(M_{n})=\infty$. Suppose the $M_{n}$ converge smoothly to a surface $M$ in $\mathscr{A}$. Then
+Suppose $M_{n}$ are compact, smoothly embedded surfaces in $Q^{c}$ such that $T_{Q}(M_{n})=\infty$. Suppose the $M_{n}$ converge smoothly to a surface $M$ in $\mathscr{A}$. (See Definition 14.) Then
 $M(T_{Q})\cap Q=\{0\},$
@@
 We can assume that each $M_{n}$ is nonfattening.
-For let $f_{n}:\mathbf{R}^{3}\to\mathbf{R}$ be a smooth function such that $f_{n}^{-1}(0)=M_{n}$, such that 00 is not a critical value of $f$, and such that $f>0$ at all points of $Q$.
+For let $f_{n}:\mathbf{R}^{3}\to\mathbf{R}$ be a smooth function such that $f_{n}^{-1}(0)=M_{n}$, such that 0 is not a critical value of $f$, and such that $f>0$ at all points of $Q$.
 Note that there are $\epsilon_{n}\uparrow 0$ such that if $\epsilon_{n}<s_{n}<0$, then the surfaces $M_{n}^{\prime}=f^{-1}(s_{n})$ will converge smoothly to $M$.
@@
 converges smoothly and with multiplicity $1$ to a cylinder $\Sigma$ whose axis is the line $\mathbf{R}p:=\{sp:s\in\mathbf{R}\}$.
-Let $A$ be an Angenent torus that lies in the non-simply connected component of $\Sigma$ and that collapses to the origin.
-Then, for all $t<T_{Q}$ sufficiently close to $T_{Q}$, $A$ and 00 will lie in different connected components of the surface (3). Thus the Angenent torus
-$\tilde{A}:=(T_{Q}-t)A+p$
-and the point $p$ will lie in different components of $M(t)$.
-Fix such a time $t$ that is regular for $M(\cdot)$ and also for each of the $M_{n}(\cdot)$.
-Then $M_{n}(t)$ converges smoothly to $M(t)$, so, for all sufficiently large $n$, $\tilde{A}$ and $p$ lie in different components of $M_{n}(t)$.
+Let $A$ be a scaled Angenent torus that lies in the non-simply connected component of the complement of $\Sigma$, and that collapses to the origin.
+Then, for all $t<T_{Q}$ sufficiently close to $T_{Q}$, $A$ and 0 will lie in different connected components of the complement of the surface (6).
+Fix such a time $t$ that is regular for flow $M(\cdot)$ and also for each of the flows $M_{n}(\cdot)$.
+Note that the scaled Angenent torus
+(7) $\tilde{A}:=(T_{Q}-t)^{1/2}A+p$
+and the point $p$ lie in different components of the complement of $M(t)$ in $\mathbf{R}^{3}$.
+(The reader may notice that the right-hand side of (7) would not be a mean curvature flow if $t$ were varying.
+But that is irrelevant: we have fixed a time $t$ and thus are defining a single surface $\tilde{A}$, not a one-parameter family of surfaces.)
+By the smooth convergence of $M_{n}(t)$ to $M(t)$, $\tilde{A}$ and $p$ lie in different components of $M_{n}(t){{}^{c}}$ for all sufficiently large $n$.
 But then, by Lemma 13, $p\in M_{n}(\tau)$ for some $\tau>t$ and thus
@@
 Theorem 23.
-Suppose $\mathscr{F}$ is a family of compact, connected, smoothly embedded, $G$-invariant surfaces in $\mathbf{R}^{3}\setminus Q$.
+Suppose $\mathscr{F}$ is a family of compact, connected, smoothly embedded, $G$-invariant surfaces in $\mathbf{R}^{3}\setminus Q$ of genus $g$.
 Suppose also that $\mathscr{F}$ is connected with respect to smooth convergence, and that all the $M\in\mathscr{F}$ lie in some large ball $B(0,R)$.
 Let $\mathscr{S}$ be the subfamily of $M\in\mathscr{F}$ for which $T_{Q}(M)<\infty$.
-If $\mathscr{S}$ and $\mathscr{S}^{c}:=\mathscr{F}\setminus\mathscr{S}$ are both nonempty, then there is an $M\in\mathscr{S}$ for which
+In other words, let $\mathscr{S}=\mathscr{F}\cap\mathscr{A}$.
+(See Definition 14.) If $\mathscr{S}$ and $\mathscr{S}^{c}:=\mathscr{F}\setminus\mathscr{S}$ are both nonempty, then there is an $M\in\mathscr{S}$ for which
 $M(T_{Q})\cap Q=\{0\}$
@@
 Proof.
-Under mean curvature flow, the sphere $\partial B(0,R)$ shrinks to a point in time $(R/2)^{1/2}$.
-Thus, if $M\in\mathscr{F}$, then $M(t)$ is empty for all $t>(R/2)^{1/2}$.
+Under mean curvature flow, the sphere $\partial B(0,R)$ shrinks to a point in time $R^{2}/4$.
+Thus, if $M\in\mathscr{F}$, then $M(t)$ is empty for all $t>R^{2}/4$.
 We claim that $\mathscr{S}$ is a closed subset of $\mathscr{F}$.
@@
 $p_{n}\in Q\cap M_{n}(t_{n}).$
-Since $p_{n}\in B(0,R)$ and since $t_{n}\leq(R/2)^{1/2}$, we can assume that $p_{n}$ and $t_{n}$ converge so to finite limits $p$ and $t$. Thus
+Since $p_{n}\in B(0,R)$ and since $t_{n}\leq R^{2}/4$, we can assume that $p_{n}$ and $t_{n}$ converge so to finite limits $p$ and $t$. Thus
 $p\in Q\cap M(t),$
@@
 For each $Q$ as in Theorem 12, there exists a family $\mathscr{F}$ satisfying the hypotheses of Theorem 23.
+The reader may regard Theorem 24 as obviously true. For completeness, we give a proof.
 Proof.
+[…]
 […]
@@
   4. (4)
-    For each $s\in(0,1)$, each component of $\{f\geq s\}$ is geodesically convex.
+    For each $s\in(0,\infty)$, each component of $\{f\geq s\}$ is geodesically convex. (There is one component in each face of $U$.)
 Let
-$h:(0,\infty)\to[0,\infty)$
-be a smooth function such that be a smooth function such that
+$h:(1/2,2)\to[0,\infty)$
+be a smooth function such that
   1. (1)
@@
   2. (2)
-    $h^{\prime}<0$ on $(0,1)$.
+    $h^{\prime}<0$ on $(1/2,1)$.
   3. (3)
-    $h^{\prime}>0$ on $(1,\infty)$.
+    $h^{\prime}>0$ on $(1,2)$.
   4. (4)
-    $h(r)\to\infty$ as $r\to 0$ or as $r\to\infty$.
+    $h(r)\to\infty$ as $r\to 1/2$ or as $r\to 2$.
 For example, we could let
-$h(r)=(\log r)^{2}.$
-Extend $h$ to the interval to $[0,\infty)$ by setting $h(0)=\infty$.
-Now define
-$\Phi:\{p:|p|\in[1/2,2]\}\to[0,\infty]$
-by
+$h(r)=\left(\log\frac{(2r-1)}{2-r}\right)^{2}.$
+Extend $h$ to the closed interval by setting $h(1/2)=h(2)=\infty$.
+Now let
+${\bf W}:=\{p\in\mathbf{R}^{3}:1/2\leq|p|\leq 2\},$
+and define $\Phi:{\bf W}\to[0,\infty]$ by
 $\Phi(p)=f(p/|p|)+h(|p|).$
-Now for $s\in[0,\infty)$, let $M^{s}$ be the set of $p$ such that $p\in[1/2,2]$ and such that
-$\Phi(p)=s.$
-Now $M^{0}=E$, which vanishes instantly under level set flow, so $T_{Q}(M^{0})=\infty$ and therefore $T_{Q}(M^{s})=\infty$ for all sufficiently small $s$.
-Note that
-$\{\Phi<\infty\}=\{p:1/2<|p|<1\}\setminus Q.$
-Thus, for all sufficiently large $s$, the region $\Phi<s$ enclosed by $M^{s}$ contains an Angenent torus centered at a point $p\in C\subset Q$.
-Thus $T_{Q}(M^{s})<\infty$ by Lemma 12.
-Now fix a $c$ such that $T_{Q}(M_{c})<\infty$.
-Then the family
-=$\mathscr{F}=\{M^{s}:0<s\leq c\}$
+For $s\in[0,\infty]$, let $M^{s}:=\Phi^{-1}(s)$. By construction, $\Phi$ is $G$-invariant and has nonvanishing gradient on
+$\operatorname{interior}({\bf W})\setminus(Q\cup E)=\bigcup_{0<s<\infty}M^{s}.$
+Hence the surfaces $M^{s}$ are $G$-invariant, smooth, and vary smoothly for $s\in(0,\infty)$.
+Also, for $s>0$ small, the genus of $M^{s}$ is the same as the genus of the boundary of a tubular neighborhood of $E=M^{0}$, which is equal to $g$.
+It follows that the genus of $M^{s}$ equals $g$ for all $s\in(0,\infty)$.
+Now, $M^{0}=E$ vanishes instantly under level set flow, so $T_{Q}(M^{0})=\infty$ and therefore $T_{Q}(M^{s})=\infty$ for all sufficiently small $s$.
+Moreover, as $s\rightarrow\infty$,
+[…]
+Thus, for all sufficiently large $s$, the region $\{\Phi<s\}\subset({\bf W}\setminus Q)$ enclosed by $M^{s}$ contains a suitably scaled Angenent torus centered at a point $p\in{\bf W}\cap Q$.
+It follows from Lemma 13 that $T_{Q}(M^{s})<\infty$.
+Therefore, the family
+$\mathscr{F}=\{M^{s}:0<s<\infty\}$
 satisfies the hypotheses of Theorem 23. This completes the proof in the case of Platonic solids.
-Now consider the when $Q$ consists of the rays
-$\{(r\cos\theta,r\sin\theta,0):r>0\}$
-such that $\theta$ is multiple of $2\pi/(g+1)$. Thus
-$Q=\{(r\cos\theta,r\sin\theta,0):r\geq 0,\,\sin\left(\frac{g+1}{2}\theta\right)=0\}.$
+Now consider the case when $Q$ consists of the rays
+$\{(r\cos\theta,r\sin\theta,0):r\geq 0\}$
 […]
-$f(r\cos\theta,r\sin\theta,z)=u\left(\cos\left(\frac{g+1}{2}\theta\right)^{2}\right),$
-where $u(s)=\frac{s}{1-s}$.
-The rest of the proof is exactly as in the Platonic solid case. ∎
 5. A Topological Lemma
 In §4, we produced shrinkers that are disjoint from certain collections of rays from the origin.
-In the remainder of the paper, we will produce shrinkers that contain certain collections of rays from the origin.
+In the remainder of the paper, we will produce shrinkers that contain certain collections of lines through the origin.
 This section proves basic facts about the topology of such surfaces.
-Let $g\geq 1$ be an integer. Let $Q=Q_{g}$ be the union of the lines
+Let $g\geq 1$ be an integer. Let $Q=Q_{g}$ be the union of the $g+1$ lines
 $\{(r\cos\theta,r\sin\theta,0):r\in\mathbf{R}\},$
@@
 For $\theta\in\mathbf{R}$, let $P_{\theta}$ be the vertical plane spanned by $(\cos\theta,\sin\theta,0)$ and $\mathbf{e}_{3}$.
-Let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ consisting of:
+Let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ generated by:
   1. (1)
@@
   2. (2)
-    Rotations about $Z$ by multiples of $\pi/(g+1)$.
+    Rotations about $Z$ by multiples of $2\pi/(g+1)$.
   3. (3)
     Rotation by $\pi$ about each of lines in $Q_{g}$.
+(Equivalently, $G_{g}$ is the group of symmetries of the surface given in cylindrical coordinates by
+$z=r^{g+1}\sin((g+1)\theta).$
+The group $G_{g}$ is also the group of symmetries of the Costa-Hoffman-Meeks surface of genus $g$.)
 In §7, we will produce (for each $g\geq 1$) a $G_{g}$-invariant shrinker $\Sigma$ of genus $g$ such that
@@
 $\operatorname{genus}(M)\leq 2g.$
-Then $M$ has genus 00, $g$, or $2g$.
-Let $c$ be the number of boundary components $C$ of $\partial M$ such that $C$ lies in $\{z>0\}$ and such that $C$ winds once around $Z$. Let $d$ be the number of points in
+Then $M$ has genus 0, $g$, or $2g$.
+Let $c$ be the number of components of $\partial M$ that lie in $\{z>0\}$ and that wind once around $Z$. Let $d$ be the number of points in
 $M\cap Z^{+}=M\cap Z\cap\{z>0\}.$
@@
 $\operatorname{genus}(M)=(c+d)g.$
+Remark 26.
 For surfaces of genus $\leq g$, a result similar to Lemma 25 was proved, using similar methods, by Buzano, Nguyen, and Schulz. See Lemma 2.10 and Corollary 2.11 of [buzano].
@@
 $\{x:(x,0,0)\in M\}$
-is a compact interval $[-R,R]$.
+is a compact interval [R,R][-R,R].
 Consider the component $J$ of
@@
 (10) $k=c+d.$
-By symmetry, the number of points of $M^{\prime}\cap Z$ is $12k+$ 1.
+By symmetry, the number of points of $M^{\prime}\cap Z$ is $2k+1$.
 Let $M^{\prime\prime}$ be the quotient of $M^{\prime}$ under the group consisting of rotations about $Z$ by angles that are multiples of $2\pi/(g+1)$.
@@
 $\chi(M^{\prime\prime})$ $=1-2\operatorname{genus}(M^{\prime\prime}).$
-Thus, multiplying (7) by $-1$ gives
+Thus, multiplying (11) by 1-1 gives
 $2\operatorname{genus}(M^{\prime})+2k=(g+1)(2\operatorname{genus}(M^{\prime\prime})+2k),$
@@
 $\operatorname{genus}(M^{\prime})=gk+2(g+1)\operatorname{genus}((M^{\prime\prime})^{+}).$
-Since $\operatorname{genus}(M^{\prime})\leq 2g$, we see that $(M^{\prime\prime})^{+}$ has genus 00, that $k$ is 00, $1$, or $2$, and thus that
+Since $\operatorname{genus}(M^{\prime})\leq 2g$, we see that $(M^{\prime\prime})^{+}$ has genus 0, that $k$ is 0, 11, or $2$, and thus that
 $\operatorname{genus}(M)$ $=\operatorname{genus}(M^{\prime})$
@@
 Proof.
-For $R>0$, let $M_{R}$ be the connected component of $M\cap B(0,R)$ that contains the origin. Since $M$ is a translator, it is connected. Thus
+For $R>0$, let $M_{R}$ be the connected component of $M\cap B(0,R)$ that contains the origin. Since $M$ is a shrinker, it is connected. Thus
 $\operatorname{genus}(M_{R})=\operatorname{genus}(M)$
@@
 Corollary 28.
-Let $M$ be as in Corollary 23. Suppose also that $M\setminus\{z=0\}$ has genus 00.
+Let $M$ be as in Corollary 27. Suppose also that $M\setminus\{z=0\}$ has genus 0.
   1. (1)
@@
 Let $\Gamma(r)=\operatorname{genus}(M\cap B(0,r))$.
-Trivially, $\Gamma$ is an increasing function of $r$, $\Gamma(r)$ is 00 for all sufficiently small $r$ and Let $M^{\prime}_{r}$ be the connected component of $M\cap B(0,r)$ containing the origin.
+Trivially, $\Gamma$ is an increasing function of $r$, and $\Gamma(r)=0$ for all sufficiently small $r$.
+Let $M^{\prime}_{r}$ be the connected component of $M\cap B(0,r)$ containing the origin.
 The portion of $M\cap B(0,R)$ not in $M^{\prime}_{r}$ lies in
 $M\setminus\{z=0\},$
-and therefore has genus 00, so
+and therefore has genus 0, so
 (13) $\Gamma(r)=\operatorname{genus}(M_{r}^{\prime}).$
-Now suppose that $M$ has genus $g$. By Lemma 22, The right hand side of (9) is either 00 or $g$. Thus there is an $r$ such that
+Now suppose that $M$ has genus $g$. By Lemma 25, The right hand side of (13) is either 0 or $g$. Thus there is an $r$ such that
 $\operatorname{genus}(M\cap B(0,\rho))=\begin{cases}0&(\rho\leq r),\\ g&(r<\rho).\end{cases}.$
@@
 […]
-$=0$
-This completes the proof when $M$ has genus $g$.
-The case when $M$ has genus $2g$ is essentially the same; the function $\Gamma(\cdot)$ only takes the values $g$ and $2g$.
-We can let $r$ be the largest $r$ for which $\Gamma(r)<g$ and $R$ be the largest $R$ for which $\Gamma(r)<2g$.
-∎
+$=0,$
+[…]
 […]
@@
 […]
-Except where otherwise specified, $N$ will refer the manifold $N$ with the Riemannian metric $\gamma:=\gamma_{1}$.
+Except where otherwise specified, $N$ will refer the manifold $N$ with the Riemannian metric $\gamma:=\gamma_{1}$. One easily checks that
+(14) $\gamma_{ij}\leq\delta_{ij}$
+at all points.
 Note that the circle
@@
 is also totally geodesic in $N$.
-If $g$ is a positive integer, we let $Q=Q_{g}$ be set of horizontal lines and $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ that were defined at the beginning of Section 5.
+If $g$ is a positive integer, we let $Q=Q_{g}$ be the set of $g+1$ horizontal lines and we let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ that were defined at the beginning of Section 5.
 Note that each $\sigma\in G$ extends to an isometry of $N:=\mathbf{R}^{3}\cup Z_{\infty}$.
@@
   4. (4)
-    $M\setminus\{z=0\}$ has genus 00.
+    $M\setminus\{z=0\}$ has genus 0.
   5. (5)
@@
 (Recall that $\{z=0\}$ is isometric to the sphere of radius $1$.)
-Most of this section is dedicated to establishing that if $M\in\mathscr{M}_{g}$, then the evolving surface $M(t)$ remains in $\mathscr{M}_{g}$ for all regular times $t$ in the interval $[0,T_{\textnormal{pos}}(M)]$.
+Most of this section is dedicated to establishing that if $M\in\mathscr{M}_{g}$, then the evolving surface $M(t)$ remains in $\mathscr{M}_{g}$ for all regular times $t$ in the interval $0,T_{\textnormal{pos}}(M)]$.
 See Theorem 35.
@@
 $\operatorname{genus}(M\setminus M^{\prime})\leq\operatorname{genus}(M\setminus\widetilde{Q})=\operatorname{genus}(M\setminus\{z=0\})=0.$
-Furthermore,
-$\operatorname{genus}(M^{\prime})=2\operatorname{genus}(M^{\prime}\cap U),$
-where $U=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}.$ Now $M^{\prime}\cap\overline{U}$ has exactly one boundary component, namely the circle $C$. Thus, by Lemma 22,
+Let $U=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}$, and note that $\overline{U}=U\cup\operatorname{Cyl}$, where $\operatorname{Cyl}$ is the cylinder defined in (16).
+Now $M^{\prime}\cap\overline{U}$ has exactly one boundary component, namely the circle $C$ $\subset\{z=0\}$.
+Thus, by Lemma 25,
 $\operatorname{genus}(M^{\prime}\cap U)=dg.$
-so $M^{\prime}$ has genus $2dg$. ∎
+But $M^{\prime}\setminus C$ is the disjoint union of $M^{\prime}\cap{U}$ and $\rho_{C}(M^{\prime}\cap{U})$. Hence
+$\operatorname{genus}(M^{\prime})=2\operatorname{genus}(M^{\prime}\cap U).$
+From the three displayed equations above, we conclude that
+$\operatorname{genus}(M)=\operatorname{genus}(M^{\prime})=2dg.$
 For $\Lambda\geq 0$, let $\operatorname{Cone}(\Lambda)$ be the set of points $p$ in
@@
 where $\pi$ is the projection $\pi(x,y,z)=(x,y,0)$ and where $\operatorname{dist}$ is geodesic distance with repect to the metric $\gamma$.
-We let $\operatorname{Cone}^{\prime}(\Lambda)$ be the set of all $p$ satisfying (13) Thus $\operatorname{Cone}(\Lambda)$ is the portion of $\operatorname{Cone}^{\prime}(\Lambda)$ in the set (12).
+We let $\operatorname{Cone}^{\prime}(\Lambda)$ be the set of all $p$ satisfying (18).
+Thus $\operatorname{Cone}(\Lambda)$ is the portion of $\operatorname{Cone}^{\prime}(\Lambda)$ in the set (17).
 Conversely, $\operatorname{Cone}^{\prime}(\Lambda)$ consists of $\operatorname{Cone}(\Lambda)$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$.
-Note that if $\Lambda$ is sufficiently large, then $M\subset\operatorname{Cone}^{\prime}(\Lambda)$.
+Note that if $\Lambda$ is sufficiently large, then
+(19) $M\subset\operatorname{Cone}^{\prime}(\Lambda).$
 Lemma 32.
-For $t\in[0,\operatorname{T_{\textnormal{fat}}}]$,
+For $t\in[0,T_{\textnormal{pos}}]$,
 $M(t)\cap\{z=0\}=\widetilde{Q}$.
-Also, if $M\in\operatorname{Cone}^{\prime}(\Lambda)$, then $M(t)\in\operatorname{Cone}^{\prime}(\Lambda)$ for all $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
+Also, if $M\in\operatorname{Cone}^{\prime}(\Lambda)$, then $M(t)\in\operatorname{Cone}^{\prime}(\Lambda)$ for all $t\in[0,T_{\textnormal{pos}}]$.
 Proof.
@@
 Note that
-(14) $\Gamma$ $:=\partial\,\Sigma_{0}$
+(20) $\Gamma$ $=\partial\,\Sigma_{0}$
 $=\Sigma_{0}\cap\{z=0\}$
@@
 where $S$ consists of alternate components of $C\setminus Q$ (see Fig. 4.)
-Figure 4. The curve $\Gamma$ in the case $g=3$.
-We prove the lemma by elliptic regularization. We work in $N\times\mathbf{R}$. We let $\widetilde{G}$ (and therefore $\widetilde{G}^{+}$) act on $N\times\mathbf{R}$ by acting on the first factor.
-Let $w:N\times\mathbf{R}\to\mathbf{R}$ be the projection onto the second factor. For $\lambda\geq 0$, let $g_{\lambda}$ be $e^{-(2/3)\lambda w}$ times the product metric on $N\times\mathbf{R}$.
+Figure 4. The curve $\Gamma$ in the case $g=4$.
+We prove the lemma by elliptic regularization.
+We work in $N\times\mathbf{R}$.
+We let $\widetilde{G}$ act on $N\times\mathbf{R}$ by acting on the first factor.
+Recall that $\widetilde{G}^{+}$ is the subgroup of $\widetilde{G}$ that fixes the positive $z$-axis of $N$.
+Let $w:N\times\mathbf{R}\to\mathbf{R}$ be the projection onto the second factor.
+For $\lambda\geq 0$, let $g_{\lambda}$ be $e^{-(2/3)\lambda w}$ times the product metric on $N\times\mathbf{R}$.
 Let $A_{\lambda}$ be a $3$-dimensional surface (flat chain mod $2$) that minimizes $g_{\lambda}$-area among surfaces in
@@
 whose boundary consists of the surface $\Sigma_{0}\times\{0\}$ together with the surface $\Gamma\times[0,\infty)$.
-Standard cut-and-paste arguments (as in [morgan]*Lemma 7.3) show that $A_{\lambda}$ is $G^{+}$-invariant.
-Alternatively, if one does not wish to invoke those arguments, one can simply let $A_{\lambda}$ minimize $g_{\lambda}$-area in the class of $G^{+}$-invariant surfaces.
+Standard cut-and-paste arguments (as in [morgan]*Lemma 7.3) show that $A_{\lambda}$ is $\widetilde{G}^{+}-$invariant.
+Alternatively, if one does not wish to invoke those arguments, one can simply let $A_{\lambda}$ minimize $g_{\lambda}$-area in the class of $\widetilde{G}^{+}-$invariant surfaces.
 We remark that, by standard GMT regularity, $A_{\lambda}$ is smooth except possibly at those points where its boundary is not smooth. However, we do not need those regularity results here.
@@
 It follows that
-(15) $t\in(0,\infty)\mapsto A_{\lambda}(t):=A_{\lambda}-t\lambda\mathbf{e}_{4},$
-$t\in(0,\infty)\mapsto A^{\prime}_{\lambda}(t):=A^{\prime}_{\lambda}-t\lambda\mathbf{e}_{4}$
+(21) $t\in(0,\infty)\mapsto A_{\lambda}(t)=A_{\lambda}-t\lambda\mathbf{e}_{4},$
+$t\in(0,\infty)\mapsto A^{\prime}_{\lambda}(t)=A^{\prime}_{\lambda}-t\lambda\mathbf{e}_{4}$
 are mean curvature flows (with moving boundaries).
@@
 (25) $\Sigma^{\prime}(t)=M(t)\quad\text{for all $t\in[0,\operatorname{T_{\textnormal{fat}}}(M))$}.$
-(See [cms]*Lemma 9.3.)
+(See [ccs]*Lemma 9.3.)
 The assertions of the lemma now follow immediately from (23), (24), and (25). ∎
 Lemma 33.
-The points in $C$ are regular points of the flow at all times $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
+The points in $C$ are regular points of the flow at all times $t\in[0,T_{\textnormal{pos}}]$.
 Proof.
-Let $\Sigma$ be a shrinker to the flow at a spacetime point $(p,t)$, where $p\in C$. By Lemma 28, $\Sigma$ (after a rotation) is contained in the set
-$\{(x,y,z):|z|\leq|x|\}.$
+Let $\Sigma$ be a shrinker to the flow at a spacetime point $(p,t)$, where $p\in C$. By (19) and Lemma 32, $\Sigma$ (after a rotation) is contained in the set
+$\{(x,y,z):|z|\leq\Lambda|x|\}.$
 By [white-boundary]*Theorem 15.1, $\Sigma$ is a union of halfplanes bounded by the $y$-axis.
@@
 Proof.
-This is a special case of the following theorem (see [white-genus]): if $M$ is a smoothly embedded, compact surface in a complete $3$-manifold of bounded Ricci curvature and if $K$ is a closed set such that $K\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}}]$, then for every regular time $t<T_{\textnormal{pos}}$,
+This is a special case of the following fact (see [white-genus]*Corollary 22): if $M$ is a smoothly embedded, compact surface, in a complete $3$-manifold NN of bounded Ricci curvature, if $K\subset N$ is the union of finitely many smooth, simple closed curves, no two of which meet tangentially, and if $K\subset\operatorname{Reg}M(t)$ for all $t\in[0,T_{\textnormal{pos}})$, then for every regular time $t<T_{\textnormal{pos}}$,
 $\operatorname{genus}(M(t)\setminus K)\leq\operatorname{genus}(M\setminus K).$
@@
 This is an immediate consequence of Lemmas 3234, and the fact that $\operatorname{area}(M(t))$ is a decreasing function of $t$. ∎
-Recall that $\{z=0\}$ is isometric to a sphere of radius $1$, so it has area $4\pi$.
 Theorem 36.
-If $\operatorname{T_{\textnormal{fat}}}=\infty$, then $M(t)$ converges smoothly to $\{z=0\}$ with multiplicity $1$.
+Suppose $M\in\mathscr{M}_{g}$.
+If $T_{\textnormal{pos}}=\infty$, then $M(t)$ converges smoothly to $\{z=0\}$ with multiplicity $1$.
 In particular, there is a time $T<\infty$ such that $M(t)$ is a smooth graph (i.e., such that $M(t)$ projects diffeomorphically onto $\{z=0\}$) for all $t\geq T$.
 Proof.
-Let $t_{n}\to\infty$. After passing to a subsequence, the flows
+Since the area of $M(t)$ is a positive, decreasing function of $t$, it has a limit $A$ as $t\to\infty$. Let $t_{n}\to\infty$. After passing to a subsequence, the flows
 $t\in[-t_{n},\infty)\mapsto M_{n}(t):=M(t+t_{n})$
-converge to a static integral Brakke flow.
-That is, there is a stationary integral varifold $V$ such that for every $t$, $M_{n}(t)$ converges (as Radon measures) to the Radon measure associated to $V$.
-Now $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$, so, by the maximum principle, $V$ is contained in $\{z=0\}$.
+converge to an integral Brakke flow $t\in\mathbf{R}\mapsto M^{\prime}(t)$.
+Note that the area of $M^{\prime}(t)$ is $A$ for every $t$.
+Since the area is independent of $t$, the flow is static: there is a stationary integral varifold $V$ such that $M^{\prime}(t)$ is the Radon measure associated to $V$ for $t$.
+Now $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$ by (19) and Lemma 32, so the support of $V$ is compact.
+In particular, $z(\cdot)$ attains a maximum $z_{\textnormal{max}}$ on the support of $V$.
+It follows from the maximum principle ([sw], for example) that $\operatorname{spt}(V)$ contains all of $\{z=z_{\textnormal{max}}\}$.
+Since $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$, it follows that $z_{\textnormal{max}}=0$.
+Likewise, the minimum of $z(\cdot)$ on $\operatorname{spt}(V)$ is 0.
 Thus $V$ is $\{z=0\}$ with some integer multiplicity $k$.
@@
 Theorem 37.
-Let $p$ be a point in $\widetilde{Q}\setminus\{O,O_{\infty}\}$. Then for all $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, $(p,T)$ is a regular point of the flow $M(\cdot)$.
+Suppose $M\in\mathscr{M}_{g}$. Let $p$ be a point in $\widetilde{Q}\setminus\{O,O_{\infty}\}$. Then for all $T\in(0,T_{\textnormal{pos}}(M)]$, $(p,T)$ is a regular point of the flow $M(\cdot)$.
 Proof.
@@
 Lemma 38.
-The origin is a regular point of the flow $M(\cdot)$ at all times $<T_{\textnormal{pos}}$.
-If $T_{\textnormal{pos}}<\infty$, then the origin is singular at time $T_{\textnormal{pos}}$, and if $\Sigma$ is a shrinker at $(O,T_{\textnormal{pos}})$, then $\Sigma$ has genus $>0$.
+The origin is a regular point of the flow $M(\cdot)$ at all times $T_{\textnormal{pos}}$.
+If $T_{\textnormal{pos}}<\infty$, then the origin is singular at time $T_{\textnormal{pos}}$, and if $\Sigma$ is a shrinker at $(O,T_{\textnormal{pos}})$, then $\Sigma$ has genus >0>0.
 Proof.
@@
 Theorem 39.
-Suppose that $T_{\textnormal{pos}}<\infty$ and that $\Sigma$ is a shrinker at the origin time $T_{\textnormal{pos}}$. Then $\Sigma\cap\{z=0\}=Q$. Furthermore,
+Suppose that $M\in\mathscr{M}_{g}$ and that $T_{\textnormal{pos}}<\infty$. Let $\Sigma$ be a shrinker at the origin at time $T_{\textnormal{pos}}$. Then $\Sigma\cap\{z=0\}=Q$. Furthermore,
 $\operatorname{genus}(\Sigma)\leq\operatorname{genus}(M)/2,$
@@
 Since $\widetilde{Q}\subset M(t)$ for each $t$, $Q\subset\Sigma$.
-Since $(0,T)$ is not a regular point (by Lemma 34), $\Sigma$ is not a plane.
+Since $(O,T)$ is not a regular point (by Lemma 38), $\Sigma$ is not a plane.
 Thus, as $\Sigma$ and $\{z=0\}$ are both minimal with respect to the shrinker metric, $\Sigma\cap\{z=0\}$ is a network of curves, the vertices of which are the points of tangency.
@@
   3. (3)
-    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
+    $\Sigma_{g}\setminus\{z=0\}$ has genus 0.
   4. (4)
@@
 It was subsquently proved for all $g$ by Buzano, Huy The Nguyen, and Schulz using minimax [buzano].
-However, none of those works proved that any of the shrinkers arose as blowups of the mean curvature flow of a compact, smoothly embedded initial surface.
-Figure 6. The piecewise smooth surface
-$\Delta_{R}\cup(\{z>0\}\cap\partial B(0,R))$
-in the case $g=1$. Its boundary consists of a pair of diameters, and a pair of circular arcs.
+Those papers also proved that, as $g\to\infty$, the shrinkers have the behavior stated in Theorem 40.
+However, none of those papers proved that any of the shrinkers arise as blowups of the mean curvature flow of a compact, smoothly embedded initial surface.
+Figure 6. The piecewise smooth surface Δ_R∪({z¿0}∩∂B(0,R)) in the case $g=1$. Its boundary consists of a pair of diameters, and a pair of circular arcs.
 Proof of Theorem 40.
@@
 $C=\{(x,y,0):(x^{2}+y^{2})^{1/2}=2\}.$
-For $20\leq R<$ 2, let
-$\Delta_{R}$ =$=\{(r\cos\theta,r\sin\theta,0):\text{$0\leq r\leq R$ and $\sin((g+1)\theta)\geq 0$}\}$
+For $0\leq R<2$, let
+$\Delta_{R}$ $=\{(r\cos\theta,r\sin\theta,0):\text{$0\leq r\leq R$ and $\sin((g+1)\theta)\geq 0$}\}$
 $\qquad\cup\{(r\cos\theta,r\sin\theta,0):\text{$R\leq r\leq 2$ and $\sin((g+1)\theta)\leq 0$}\}.$
@@
 Note that we can choose $S^{+}$ so that $M$ is smooth and $\widetilde{G}-$invariant.
-Note that $S$ can be obtained by taking the union of the disk $D$ and the sphere $\partial B(0,R)$ and then doing surgery along the circle of intersection.
+Note that $S$ can be obtained by taking the union of the disk $D$ of Euclidean radius 2 in the plane $\{z=0\}$ and the sphere $\partial B(0,R)$ and then doing surgery along the circle of intersection.
 The Euler characteristic of the disjoint union is $1+2=3$.
@@
   4. (4)
-    $M\setminus\{z=0\}$ has genus 00.
+    $M\setminus\{z=0\}$ has genus 0.
   5. (5)
@@
 In particular, all the theorems in §6 apply to the surfaces in $\mathscr{M}_{g}^{\prime}$.
-Now let $M$ be any surface in $\mathscr{M}_{g}^{\prime}$. We orient $M$ so that the unit normal at $(2,0,0)$ is $\mathbf{e}_{3}$. It follows that the unit normal at the origin is $-\mathbf{e}_{3}$:
+Now let $M$ be any surface in $\mathscr{M}_{g}^{\prime}$. We orient $M$ so that the unit normal at $(2,0,0)$ is $\mathbf{e}_{3}$. It follows that the unit normal at the origin is $\mathbf{e}_{3}$:
 (26) $\nu_{M}(2,0,0)$ $=\mathbf{e}_{3},$
@@
 Proof.
-By Theorems 33 and 34, the flow is regular at all points of $\tilde{Q}$ for all times $t\in[0,T_{\textnormal{pos}})$.
+By Theorems 37 and 38, the flow is regular at all points of $\tilde{Q}$ for all times $t\in[0,T_{\textnormal{pos}})$.
 In particular, the points $(0,0,0)$ and $(2,0,0)$ are such points.
@@
 and the behavior of $\Sigma_{g}$ as $g\to\infty$.
-By Theorem 37 below, we can choose $\Sigma_{g}$ to have entropy $<\delta$. By Theorem 38, given that entropy bound, $\Sigma_{g}$ has the asserted behavior as $g\to\infty$. ∎
+By Theorem 41 below, we can choose $\Sigma_{g}$ to have entropy $\delta$.
+By Theorem 42, given that entropy bound, $\Sigma_{g}$ has the behavior asserted in the statement of Theorem 40, as $g\to\infty$.
+∎
 Theorem 41.
@@
 […]
-Note that there exist $M\in\mathscr{M}^{\prime}_{g}$ that are arbitrarily close to $\Psi_{r}$ in the weak sense.
-That is, there exists a sequence $M_{i}\in\mathscr{M}^{\prime}_{g}$ such that $M_{i}$ converges weakly to $\Psi_{r}$.
+Note that there exist $M\in\mathscr{M}^{\prime}_{g}$ that are arbitrarily close to $\Psi(r)$ in the weak sense.
+That is, there exists a sequence $M_{i}\in\mathscr{M}^{\prime}_{g}$ such that $M_{i}$ converges weakly to $\Psi(r)$.
 […]
   1. (1)
-    $M_{i}$ converges weakly with with multiplicity $1$ to $\{z=0\}$
+    $M_{i}$ converges weakly with multiplicity $1$ to $\{z=0\}$
   2. (2)
 […]
-(Here $SS\SS(\infty)$ is the empty set.)
+[…]
 Let $T_{i}=T_{\textnormal{pos}}(M_{i})$ and, as usual, let
@@
 be the mean curvature flow with $M_{i}(0)=M_{i}$. Let $\Sigma_{i}$ be a shrinker to the flow at the origin at time $T_{i}$.
+We claim that
+(29) $T_{i}\to 0.$
+For suppose not.
+Then (after passing to a subsequence) $T_{i}$ converges to a limit $T\in(0,\infty]$.
+Since $M_{i}$ converges weakly with multiplicity $1$ to $\{z=0\}$, the flow $t\in[0,T_{i}]\mapsto M_{i}(t)$ converges weakly to the multiplicity-one flow
+$t\in[0,T]\mapsto\{z=0\}.$
+By local regularity theory [white-local], the convergence is smooth on $\epsilon,T]$ for every $\epsilon>0$.
+In particular, $M_{i}(\epsilon)$ is a smooth graph for all sufficiently large $i$.
+But that contradicts Claim 1 in the proof of Theorem 40, thus proving (29).
 For $t\in[0,1)$, let
@@
 $t\in[0,1)\mapsto M_{i}^{\prime}(t)$
-is a mean curvature flow in $\mathbf{R}^{2}\cup Z_{\infty}$, but with the dilated metric $\gamma_{\lambda_{i}}$, where $\lambda_{i}:=(T_{i})^{-1/2}$.
+is a mean curvature flow in $\mathbf{R}^{3}\cup Z_{\infty}$, but with the dilated metric $\gamma_{\lambda_{i}}$, where $\lambda_{i}:=(T_{i})^{-1/2}$.
 Let $\operatorname{area}_{s}$ denote area with respect to the shrinker metric.
@@
 where $\epsilon_{i}\to 0$.
-(The $\epsilon_{i}$ is there because $M_{i}^{\prime}(\cdot)$ is mean curvature flow for a non-flat metric; and $\epsilon_{i}\to 0$ because the metric is converging to smoothly to the flat metric.)
+(The $\epsilon_{i}$ is there because $M_{i}^{\prime}(\cdot)$ is mean curvature flow for a non-flat metric; and $\epsilon_{i}\to 0$ because the metric is converging to smoothly to the flat metric by (29).)
 […]
@@
   3. (3)
-    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
+    $\Sigma_{g}\setminus\{z=0\}$ has genus 0.
   4. (4)
@@
 Conjectured qualitative appearance of the shrinker $\Sigma_{g}$ in case $g=1$.
-The surface shown has genus $2=2g$ and it seems to have $3$ ends.
-However, we know that $\Sigma_{g}$ has genus $2g$ and $3$ ends only when $g$ is large.
+The surface shown has genus $2=2g$ and it seems to have 33 ends.
+However, we know that $\Sigma_{g}$ has genus $2g$ and 33 ends only when $g$ is large.
 Figure courtesy of Matthias Weber.
@@
   3. (3)
-    In Theorem 55 and Proposition 58, we show, for all sufficiently large $g$, that $\Sigma_{g}$ has the remaining properties asserted in Theorem 39: it has genus $2g$, it has exactly $3$ ends, and it converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$.
+    In Theorem 59 and Proposition 62, we show, for all sufficiently large $g$, that $\Sigma_{g}$ has the remaining properties asserted in Theorem 43: it has genus $2g$, it has exactly 33 ends, and it converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$.
 To prove these theorems, we work in
@@
 Let
-$D$ =$=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq 2\},$
-$D(R)$ =$=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq R\},$
+$D$ $=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq 2\},$
+$D(R)$ $=\{(x,y,0):(x^{2}+y^{2})^{1/2}\leq R\},$
 $A^{+}$ $=\partial(B(0,R)\cap\{z>0\}),$
@@
 Let $k$ be an integer, and consider the $2(g+1)$ wedge regions
-$W_{k}:=\frac{\pi k}{g+1}<\theta<\frac{\pi(k+1)}{g+1}.$
-Let $\Omega(s)$, $10\leq s<$ 1, be a smooth, $1$-parameter family of nested, closed convex regions in $W_{0}\cap D(R)$ such that $\Omega(0)$ is a single point and such that $\cup_{s}\Omega(s)$ is all of $W_{0}\cap D(R)$.
+$W_{k}:=\frac{\pi k}{g+1}<\theta<\frac{\pi(k+1)}{g+1},$
+where $0\leq k\leq 2g+1$.
+Let $\Omega(s)$, $0\leq s<1$, be a smooth, $1$-parameter family of nested, closed convex regions in $W_{0}\cap D(R)$ such that $\Omega(0)$ is a single point, such that $\cup_{s}\Omega(s)$ is all of $W_{0}\cap D(R)$, and such that each $\Omega(s)$ is symmetric under reflection in the plane $P_{\pi/2(g+1)}$.
 For $s\in[0,1)$, define $h(s)$ by
@@
 $A^{-}_{s}$ $:=A^{-}+(0,0,-h(s)).$
-For $10<s<$ 1, we do surgery on $D\cup A^{+}_{s}\cup A^{-}_{s}$ as follows.
-For each $W_{k}$ with $k$ even, we connect $D$ to $A^{+}_{s}$ be a neck in the region $W_{k}$.
+For $0<s<1$, we do surgery on $D\cup A^{+}_{s}\cup A^{-}_{s}$ as follows.
+For each $W_{k}$ with $k$ even, we connect $D$ to $A^{+}_{s}$ by a neck in the region $W_{k}$.
 Likewise, for each $W_{k}$ with $k$ odd, we connected $D$ to $A^{-}_{s}$ by a neck in $W_{k}$.
@@
 from $A^{+}_{s}$, and then attach the annular surface
-[…]
+$\{(x,y,z):(x,y,0)\in\partial\Omega(s),\,0\leq z\leq h(s)\}.$
 The resulting surface has $(g+1)$ necks in $\{0\leq z\leq h(s)\}$ and $(g+1)$ necks in $\{-h(s)\leq z\leq 0\}$.
@@
 $\leq(g+1)^{-1}2\pi R^{3},$
-where the second-to-last inequality follows from (23).
-The same bound holds for $\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$ since $\gamma_{ij}\leq\delta_{ij}$ everywhere.
+where the second-to-last inequality follows from (30). The same bound holds for $\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$ since $\gamma_{ij}\leq\delta_{ij}$ everywhere (14).
 Consequently, since $M[s]$ has $4(g+1)$ necks,
@@
 $=4\pi+12\pi R^{2}+8\pi R^{3}.$
-We fix an $R$ for which this last expression is $<5\pi$. Thus
+We fix an $R$ for which this last expression is $5\pi$. Thus
 $\operatorname{area}_{\,\gamma}(M[s])<5\pi$
@@
 $\operatorname{area}_{\,\gamma}(M^{\prime}[s])<5\pi.$
-From the construction, it is straightforward to check that $M^{\prime}[s]$ has genus $4g$ and that $M^{\prime}[s]\setminus\{z=0\}$ has genus 00.
-Thus $M^{\prime}[s]\in\mathscr{M}_{g}^{\prime\prime}$.
-∎
+From the construction, it is straightforward to check that $M^{\prime}[s]$ has genus $4g$ and that $M^{\prime}[s]\setminus\{z=0\}$ has genus 0. Thus $M^{\prime}[s]\in\mathscr{M}_{g}^{\prime\prime}$. ∎
 Remark 45.
@@
 $\frac{\operatorname{area}(M[s]\cap B(0,r))}{\pi r^{2}}<\Lambda$
-for all $s\in(0,1)$ and all $20<r<$ 2.
+for all $s\in(0,1)$ and all $0<r<2$.
 Proof.
 Note that
-$\operatorname{area}(B(0,r)\cap(D\cup A^{+}_{s}\cup A^{-}_{s})$ $=\pi r^{2}+2\operatorname{area}(B(0,r)\cap A^{+}_{s})$
+[…]
 $\leq\pi r^{2}+2\operatorname{area}(B(0,r)\cap A^{+})$
-=$=\begin{cases}3\pi r^{2}&\text{if $r<R$}\\ \pi r^{2}+6\pi R^{2}&\text{if $r>R$}.\end{cases}$
+$=\begin{cases}3\pi r^{2}&\text{if $r<R$}\\ \pi r^{2}+6\pi R^{2}&\text{if $r>R$}.\end{cases}$
 $\leq 7\pi r^{2}.$
@@
 Definition 48.
-Let $\mathscr{F}=\mathscr{F}_{g}$ be the family of smooth surfaces $M^{\prime}[s]$ constructed in the proof of Proposition 40, where the smoothing has been done in accord with Remarks 41 and 43.
+Let $\mathscr{F}=\mathscr{F}_{g}$ be the family of smooth surfaces $M^{\prime}[s]$ constructed in the proof of Proposition 44, where the smoothing has been done in accord with Remarks 45 and 47.
 Theorem 49.
@@
 Furthermore, if we give $M$ the orientation such that $\nu_{M}(2,0,0)=\mathbf{e}_{3}$, then
-$\nu_{M(t)}(0,0,0)=-\mathbf{e}_{3}$
+$\nu_{M(t)}(0,0,0)=\mathbf{e}_{3}$
 for all $t\in[0,T_{\textnormal{pos}}(M))$, where the orientation is the standard induced orientation as defined in Section 3.
@@
 Since $(2,0,z)\in M(t)$ only for $z=0$, we see that
-$(1,0,z)\in K(t)\iff z\leq 0.$
+$(2,0,z)\in K(t)\iff z\leq 0.$
 Thus if $p$ is any point for which
@@
 then $p\notin K(t)$. Consequently, at the highest point of $M(t)\cap Z$, $\nu_{M(t)}=\mathbf{e}_{3}$, and the next highest point, $\nu_{M(t)}=-\mathbf{e}_{3}$, and so on. Thus
-$\nu_{M(t)}(0,0,0)=\begin{cases}\mathbf{e}_{3}&\text{if $n(t)$ is even,}\\ -\mathbf{e}_{3}&\text{if $n(t)$ is odd.}\end{cases}$
+(37) $\nu_{M(t)}(0,0,0)=\begin{cases}\mathbf{e}_{3}&\text{if $n(t)$ is even,}\\ -\mathbf{e}_{3}&\text{if $n(t)$ is odd.}\end{cases}$
 Thus, by (36), the parity of $n(t)$ is the same for all regular times $t\in[0,T_{\textnormal{pos}})$.
-Now at time 00,
+Now at time 0,
 $4g$ $=\operatorname{genus}(M)$
@@
 $=2n(0)$
-by Lemma 22. Thus $n(t)$ is even for all regular times $t\in[0,T_{\textnormal{pos}})$. ∎
+by Lemma 25. Thus, for all regular times $t\in[0,T_{\textnormal{pos}})$, $n(t)$ is even and so $\nu_{M(t)}(0,0,0)=\mathbf{e}_{3}$ by (37). ∎
 Definition 50.
@@
 (39) $L\cap\{0<|z|<\epsilon\}=I.$
-By the Initial Regularity Theorem (Theorem 8), there is a time $\tau>0$ such that for all sufficiently large $n$, the flow $M_{n}(\cdot)$ has no singularities on $Z$ during the time interval $[0,\tau]$.
+By the Initial Regularity Theorem (Theorem 9), there is a time $\tau>0$ such that for all sufficiently large $n$, the flow $M_{n}(\cdot)$ has no singularities on $Z$ during the time interval $0,\tau]$.
 In particular, for $t\in[0,\tau]$, $M_{n}(t)\cap Z$ consists of exactly $5$ points, and those points depend smoothly on $t$.
@@
   2. (2)
-    $\Sigma\setminus\{z=0\}$ has genus 00,
+    $\Sigma\setminus\{z=0\}$ has genus 0,
   3. (3)
@@
 Trivially, $\Sigma_{g}$ is $G_{g}$-invariant.
-Note that $M(t)\setminus\{z=0\}$ has genus 00 for all regular $t<T_{\textnormal{pos}}$ by Theorem 45.
-Hence $\Sigma\setminus\{z=0\}$ has genus 00.
+Note that $M(t)\setminus\{z=0\}$ has genus 0 for all regular $t<T_{\textnormal{pos}}$ by Theorem 49.
+Hence $\Sigma\setminus\{z=0\}$ has genus 0.
 By Theorem 39, $\Sigma\cap\{z=0\}=Q_{g}$ and $\Sigma$ has genus $g$ or $2g$.
@@
 $t\in I\mapsto M(t)$
-that admits a strong orientation $\nu$ with the following property: for each $t$, if the origin is a regular point at time $t$, then $\nu(O,t)=-\mathbf{e}_{3}$.
+that admits a strong orientation $\nu$ with the following property: for each $t$, if the origin is a regular point at time $t$, then $\nu(O,t)=\mathbf{e}_{3}$.
 Lemma 58.
@@
 $\tilde{M}(t)=M(t)\quad\text{for $0\leq t\leq\operatorname{T_{\textnormal{fat}}}(M)$}.$
-By the Local Regularity Theorem (Theorem 7), $\tilde{M}(\cdot)$ is special since each $M_{n}(\cdot)$ is.
+[…]
 Now
 (42) $t\in(-\infty,0)\mapsto|t|^{-1/2}\Sigma$
-occurs as a tangent flow to $M(\cdot)$ at the spacetime point $(0,T)$, so it also occurs as a tangent flow to $\tilde{M}(\cdot)$ at that point, since the two flows coincide up to time $T$.
-Thus there is a sequence $\lambda_{n}\to\infty$ such that the parabolically dilated flows
+[…]
 (43) $t\in[-\lambda_{n}T,\infty)\mapsto\lambda_{n}^{1/2}\tilde{M}(T+\lambda_{n}^{-1}t)$
 converge for $t<0$ to the flow (42).
-By passing to a subsequence (if necessary) we can assume that the flows converge for all $t$ a flow
+By passing to a subsequence (if necessary) we can assume that the flows (43) converge for all $t$ to a flow
 (44) $t\in\mathbf{R}\mapsto M^{\prime}(t)$
-Since the $\tilde{M}(\cdot)$ is special, so are the flows (35), and therefore (by the Local Regularity Theorem, Theorem 7) so is the flow (36). ∎
+[…]
 Theorem 59.
@@
 Note that
-$M^{\prime}(g)=|t|^{-1/2}\Sigma\quad(t<0).$
-Note that $M^{\prime}(0)$, the limit of $|t|^{-1/2}\Sigma$ as $t\to 0$ is the $P$, the plane $\{z=0\}$ with multipicity $1$.
-Hence $M^{\prime}(t)=P$ for all $t\geq 0$.
+$M^{\prime}(g)=|t|^{-1/2}\Sigma\quad(t\leq 0).$
+As a set, $M^{\prime}(0)$ is $P$ if $S$ is the sphere and $P\cup Z$ if $S$ is the cylinder.
+In either case, the corresponding Radon measure is the plane $P$ with multiplicity one.
+It follows (by unit regularity) that $M^{\prime}(t)$ is the plane $P$ with multiplicity one for all $t>0$.
 That is, $M^{\prime}(\cdot)$ is the flow
-(37) $t\in\mathbf{R}\mapsto\begin{cases}|t|^{-1/2}\Sigma&(t<0),\\ \Sigma&(t\geq 0).\end{cases}$
+(45) $t\in\mathbf{R}\mapsto\begin{cases}|t|^{-1/2}\Sigma&(t\leq 0),\\ P&(t>0).\end{cases}$
 Since $M^{\prime}(\cdot)$ is a limit of special flows, it is also special.
-Note that at each time $t=(-1)$, the singular set is the circle $J:=S\cap P$.
+Note that at time t=1t=-1, the singular set of (45) is the circle $J:=S\cap P$.
 Thus, by specialness, there is a closed set $K$ such that $K\setminus J$ is a smooth manifold-with-boundary.
-At the origin, $-\mathbf{e}_{3}$ points out of $K$.
-It follows that $K$ consists of the closure of the region above $P$ and inside $S$, together with the closure of region the region below $P$ and outside $S$.
+At the origin, $\mathbf{e}_{3}$ points out of $K$.
+It follows that $K$ consists of the closure of the region below $P$ and inside $S$, together with the closure of region the region above $P$ and outside $S$.
 See Fig.
 9.
-Figure 9. The region $K$ (in the case that $S$ is a sphere)
-By specialness, at time $t=1$,
-Thus $\nu(p,-1)=\mathbf{e}_{3}$ at all points $p$ in $\{z=0\}$ that lie outside $S$.
+Figure 9. The region $K$ (in the case that $S$ is a sphere).
+Thus $\nu(p,-1)=-\mathbf{e}_{3}$ at all points $p$ in $\{z=0\}$ that lie outside $S$.
 By the same argument (or simply by continuity)
-$\nu(p,t)=\mathbf{e}_{3}$
+$\nu(p,t)=-\mathbf{e}_{3}$
 for all $t<0$ and $p$ in $\{z=0\}$ outside $|t|^{-1/2}S$.
@@
 for all $p\neq 0$ in $\{z=0\}$. Therefore, also by continuity,
-$\nu(p,t)=\mathbf{e}_{3}$
-for all $p$ in $\{z=0\}$ and $t>0$. In particular, $\nu(0,t)=\mathbf{e}_{3}$ for all $t>0$, contrary to specialnesss of $M^{\prime}(\cdot)$. ∎
+$\nu(p,t)=-\mathbf{e}_{3}$
+for all $p$ in $\{z=0\}$ and $t>0$. In particular, $\nu(0,t)=-\mathbf{e}_{3}$ for all $t>0$, contrary to specialnesss of $M^{\prime}(\cdot)$. ∎
 9. Limits of Shrinkers
@@
 The proof of [brendle]*Theorem 2 shows that $\Sigma_{g}^{+}$ is connected.
-Since $\Sigma_{g}^{+}$ has genus 00, $\Gamma_{g}$ divides $\Sigma_{g}^{+}$ into two connected components $U_{g}$ and $U_{g}^{\prime}$.
-Note that $U_{g}$ and $U_{g}^{\prime}$ are each $G_{g}^{+}$-invariant.
-Since $\Sigma_{g}$ is connected, the closure of at least one component, which we may assume is $U_{g}^{\prime}$, contains points in $Q_{g}$.
-It follows that $\overline{U_{g}^{\prime}}$ contains an entire ray of $Q$.
-By $G_{g}^{+}$ invariance, $\overline{U_{g}^{\prime}}$ contains all of $Q$.
+Since $\Sigma_{g}^{+}$ has genus 0, $\Gamma_{g}$ divides $\Sigma_{g}^{+}$ into two connected components $U_{g}$ and $U_{g}^{\prime}$.
+Now $Q=Q_{g}$ is in the closure of $\Sigma_{g}\setminus Q$, and, therefore, by symmetry, in the closure of $\Sigma_{g}^{+}$.
+Thus the closure of one of the components, say $\overline{U_{g}^{\prime}}$, of $\Sigma_{g}\setminus\Gamma_{g}$, contains a point of $Q\setminus\{O\}$.
+Hence $\overline{U_{g}^{\prime}}$ contains the corresponding connected component of $Q_{g}\setminus\{O\}$.
+By $G_{g}^{+}$ symmetry, $\overline{U_{g}^{\prime}}$ contains all of $Q_{g}\setminus\{O\}$ and therefore all of $Q_{g}$.
 It follows that $\Gamma_{g}$ divides $\Sigma_{g}$ into two components, namely $U_{g}$ and
@@
 Proof.
-By passing to a subsequence, we can assume that the $\Sigma_{g}$ converge as sets to a limit set $\Sigma$.
+By passing to a subsequence, we can assume that the $\Sigma_{g}$ converge as sets to a limit set $\Sigma$ that is rotationally invariant about $Z$.
 We can also assume that there is a closed set $K$ (the curvature blowup set) such the convergence $\Sigma_{g}$ to $\Sigma$ is smooth (perhaps with multiplicity) in compact subsets of $\mathbf{R}^{3}\setminus K$ and such that
@@
 Since $Q_{g}\subset\Sigma_{g}$, it follows that $\{z=0\}$ is contained in $\Sigma$, and thus that
-(38) $\Sigma$ =$=\{z=0\}\cup\Sigma^{+}\cup\Sigma^{-}$
-=$=\{z=0\}\cup\Sigma^{+}\cup\rho_{X}\Sigma^{+},$
+(46) $\Sigma$ $=\{z=0\}\cup\Sigma^{+}\cup\Sigma^{-}$
+$=\{z=0\}\cup\Sigma^{+}\cup\rho_{X}\Sigma^{+},$
 where $\Sigma^{+}=\Sigma\cap\{z>0\}$, $\Sigma^{-}=\Sigma\cap\{z<0\}$, and $\rho_{X}$ denotes rotation by $\pi$ about the $x$-axis, $X$.
-By assumption, $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$ has genus 00.
+By assumption, $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$ has genus 0.
 By the compactness theorem [white-compact]*Theorem 1.1 for minimal surfaces of locally bounded area and genus, $\Sigma^{+}$ is smooth and $K\cap\{z>0\}$ is a discrete subset of $\{z>0\}$.
@@
 Proof of Claim 1.
-Suppose $\Sigma^{+}$ is nonempty.
+Suppose $\Sigma^{+}$ is nonempty. Note that
+(48) $\overline{\Sigma^{+}}\cap\{z=0\}\neq\emptyset,$
+as otherwise $\Sigma^{+}$ would be a smooth shrinker (without boundary) in $\{z>0\}$.
+Note also that $\overline{\Sigma^{+}}$ is a surface of revolution.
 Let $C^{\prime}$ be a horizontal circle in $\Sigma^{+}$.
-Let $C_{q}^{\prime}$ be smooth, simple closed curves in $\Sigma_{q}$ that converge smoothly to $C^{\prime}$.
-By Lemma 56, $C_{q}^{\prime}$ bounds a region $E_{n}$ in $\Sigma_{n}$ that is either a disk or an annulus, that lies in $\{z>0\}$, and that is properly embedded in $\mathbf{R}^{3}$.
+Let $C_{g}^{\prime}$ be smooth, simple closed curves in $\Sigma_{g}$ that converge smoothly to $C^{\prime}$.
+By Lemma 60, $C_{g}^{\prime}$ bounds a region $E_{g}$ in $\Sigma_{g}$ that is either a disk or an annulus, that lies in $\{z>0\}$, and that is properly embedded in $\mathbf{R}^{3}$.
 Passing to the limit, we see that $C^{\prime}$ bounds a region $E$ in $\Sigma^{+}$ that is either a disk or an annulus, and that is properly embedded in $\mathbf{R}^{3}$.
-The other end $E^{\prime}$ of $\Sigma^{+}$ must diverge to infinity, converge to the origin, or converge to a circle $C(r)$ in the plane $\{z=0\}$.
-The end $E^{\prime}$ cannot diverge to infinity, because then $\Sigma^{+}$ would be shrinker (without boundary) in $\{z>0\}$.
-The end $E^{\prime}$ cannot converge the origin, for if it did, then $\Sigma^{+}\cup\{0\}$ and $\{z=0\}$ would violate the strong maximum principle at the origin.
+By the strong maximum principle, $E$ lies in $\{z>0\}$. Since $E$ is properly embedded in $\mathbf{R}^{3}$,
+(49) $\overline{E}\cap\{z=0\}=\emptyset.$
+Now consider the other portion $E^{\prime}$ of $\Sigma^{+}$ bounded by $C^{\prime}$. By (48) and (49),
+(50) $\overline{E^{\prime}}\cap\{z=0\}\neq\emptyset.$
+Note that $E^{\prime}$ is a surface of revolution generated by a curve $e^{\prime}$ in the plane $y=0$, and that $e^{\prime}$ has one endpoint $p$ in $C^{\prime}\cap\{y=0\}$.
+By (50), the closure of $e^{\prime}$ contains a point $q$ in $\{z=0\}$.
+Since $\Sigma$ has finite entropy, the curve $e^{\prime}$ has locally finite length.
+Thus $e^{\prime}$ is a curve joining $p$ to $q$.
+Now $q$ cannot be the origin, since otherwise $\Sigma^{+}\cup\{0\}$ and $\{z=0\}$ would violate the strong maximum principle at the origin.
 (One can invoke various versions of the maximum principle here.
 For example, the classical strong maximum principle suffices, because $\Sigma^{+}\cup\{0\}$ is a smooth minimal surface with respect to the shrinker metric by [structure]*Corollary 2.7.)
-Thus the end $E^{\prime}$ converges to a circle $C(r)$ in the plane $\{z=0\}$. This completes the proof of Claim 1. ∎
+Thus $q=(x,0,0)$ for some $x\neq 0$.
+Therefore, $\overline{\Sigma^{+}}$ is a properly embedded disk or annulus whose boundary is the circle $C(r)$ (where $r=|x|$) in the plane $\{z=0\}$.
+This completes the proof of Claim 1.
+∎
 Recall (Corollary 28) that there exist $r_{g}\in(0,\infty)$ and $R_{g}\in[r_{g},\infty]$ such that
@@
 Proof.
+Since $X\subset Q_{g}\subset\Sigma_{g}$, it follows that $p\in\Sigma_{g}$.
 By rotational symmetry by $\pi$ about $X$, if $L$ is a closed vertical line segment centered at $p$, then $L\cap\Sigma_{g}$ has an odd number of points (namely, the point $p$ and equal numbers of points above and below $p$.) ∎
@@
 is the corresponding vector that points into $\Sigma^{-}$. By stationarity,
-$-m\mathbf{e}_{1}+m^{\prime}\mathbf{e}_{1}+\nu^{+}+\nu^{-}=0.$
+$m\mathbf{e}_{1}+m^{\prime}\mathbf{e}_{1}+\nu^{+}+\nu^{-}=0.$
 Taking the dot product with $\mathbf{e}_{1}$ gives
@@
 Now $\Sigma_{g}\cap\overline{B(0,r)}$ converges smoothly and with multiplicity $m$ away from the origin to $\overline{D(r)}$.
-If $m=1$, the convergence is smooth by Allard’s Regularity Theorem. Thus we assume that $m>1$.
+If $m=1$, the convergence is smooth by Allard’s Regularity Theorem. Thus we may assume that $m>1$.
 For all sufficiently large $g$, $\Sigma_{g}\cap\partial B(0,r)$ consists of $m$ simple closed curves.
 Let $\Gamma_{g}^{0}$ be the curve that passes through the points of $Q_{g}\cap\partial B(0,r)$.
-Let $\Gamma_{g}$ be one of the other components.
+Let $\Gamma_{g}$ be one of the other curves.
 By symmetry, we may assume that $\Gamma_{g}$ lies in $\{z>0\}$.
-Let $\Sigma_{g}$ be the component of $\Sigma_{g}\setminus\Gamma_{g}$ that lies in $\{z>0\}$.
-(The component exists by Lemma 56).
-Since $\Sigma^{g}\setminus B(0,r)$ is connected and contains points in $\{z\leq 0\}$, we see that $\Sigma_{g}$ is contained in $B(0,r)$.
+Let $\Sigma_{g}^{\prime}$ be the component of $\Sigma_{g}\setminus\Gamma_{g}$ that lies in $\{z>0\}$.
+(The component exists by Lemma 60.) Since $\Sigma_{g}\setminus B(0,r)$ is connected and contains points in $\{z\leq 0\}$, we see that $\Sigma_{g}^{\prime}$ is contained in $B(0,r)$.
 We have shown: each component of $\Sigma_{g}\cap\partial B(0,r)$ other than $\Gamma_{g}^{0}$ bounds a connected component of $\Sigma_{g}\cap B(0,r)$.
@@
 Now it has at least $2$ components and at most $3$ (since $\Sigma_{g}\cap\partial B(0,R)$ has $3$ components.) The number of components is odd (by symmetry) and thus it is $3$.
-But now (exactly as in the proof of Claim 4) each of those components converges smoothly to the disk $\overline{D(0,R)}$, so $\Sigma_{g}\cap B(0,R)$ has genus 00, contradicting $r_{g}<R$.
+But now (exactly as in the proof of Claim 4) each of those components converges smoothly to the disk $\overline{D(0,R)}$, so $\Sigma_{g}\cap B(0,R)$ has genus 0, contradicting $r_{g}<R$.
 This completes the proof in Case 1.
-Case 2: $\Sigma^{+}$ is nonempty. By Claim 1, its boundary is a circle $C(\rho)$. By Claim 3, $\Sigma^{+}$ meets $\{z=0\}$ orthogonally, so
-$\Sigma:=\Sigma^{+}\cup\Sigma^{-}\cup C(\rho)$
-is a smoothly embedded, rotationally invariant shrinker $\Sigma$. Thus $\Sigma$ is a sphere or a cylinder.
+Case 2: $\Sigma^{+}$ is nonempty. By Claim 1, its boundary is a circle $C(\rho)$. By Claim 3, $\Sigma^{+}$ meets $\{z=0\}$ orthogonally, so the surface
+$\widetilde{\Sigma}:=\Sigma^{+}\cup\Sigma^{-}\cup C(\rho)$
+is a smoothly embedded, rotationally invariant shrinker. Thus $\widetilde{\Sigma}$ is a sphere or a cylinder.
 Now choose $R>0$ such that $R$ is not equal to $r_{\infty}$ or $R_{\infty}$, and large enough that $\Sigma_{g}\cap B(0,R)$ is connected for all $g\in\mathscr{G}$. (Such an $R$ exists by Lemma 65.)
-If $\Sigma$ is a sphere, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m=2c+1$ simple closed curves, $c$ of which lie in $\{z>0\}$.
+If $\widetilde{\Sigma}$ is a sphere, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m=2c+1$ simple closed curves, $c$ of which lie in $\{z>0\}$.
 Also, $\Sigma_{g}\cap Z^{+}$ consists $c+1$ points, $c$ of which are close to the origin and one of which is close to $\Sigma^{+}\cap Z$.
@@
 $\operatorname{genus}(\Sigma_{g}\cap B(0,R))=(2c+1)g$
-If, on the other, $\Sigma$ is a cylinder, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m+2$ simple closed curves, $m$ of which lie near the circle $C(R)$ and two of which are near the circles of intersection of the cylinder $\Sigma$ and the sphere $\partial B(0,R)$.
+If, on the other hand, $\widetilde{\Sigma}$ is a cylinder, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m+2$ simple closed curves, $m$ of which lie near the circle $C(R)$ and two of which are near the circles of intersection of the cylinder $\widetilde{\Sigma}$ and the sphere $\partial B(0,R)$.
 Note that $c+1$ of those curves lie in $\{z>0\}$.
@@
 (53) $\operatorname{genus}(\Sigma_{g}\cap B(0,R))=(2c+1)g.$
-We have shown that (42) holds whether $\Sigma$ is a sphere or a cylinder. Since $\Sigma_{g}$ has genus $\leq 2g$, we see that $c=0$. This completes the proof of Theorem 57 in Case 2. ∎
+We have shown that (53) holds whether $\widetilde{\Sigma}$ is a sphere or a cylinder. Since $\Sigma_{g}$ has genus $\leq 2g$, we see that $c=0$. This completes the proof of Theorem 61 in Case 2. ∎
 Proposition 62.
@@
 Suppose that $M_{i}$ converges smoothly (and with multiplicity one) to $M$ on compact subsets of $\mathbf{R}^{3}\setminus\overline{B(0,R)}$.
-Then the convergence $M_{i}\to M$ is uniform at infinity in the following sense: if $\lambda_{i}\to 0$, then $\lambda_{i}\Sigma_{i}$ converges smoothly (with multiplicity $1$) on compact subsets of $\mathbf{R}^{3}\setminus\{0\}$ to $C\setminus\{0\}$.
+Then the convergence $M_{i}\to M$ is uniform at infinity in the following sense: if $\lambda_{i}\to 0$, then $\lambda_{i}M_{i}$ converges smoothly (with multiplicity $1$) on compact subsets of $\mathbf{R}^{3}\setminus\{0\}$ to $C\setminus\{0\}$.
 Proof.
@@
 (It may have more components.) Let $\Sigma_{n}^{\prime}$ be a surface in $\mathbf{R}^{3}\setminus B(0,r_{n})$ of least shrinker area that separates $C_{n}$ and $C_{n}^{\prime}$.
-Then $\Sigma_{n}^{\prime}$ is a surface whose boundary is in $\partial B(0,r_{n})$. It is smoothly embedded away from the boundary, and its shrinker area is at most
+Thus $\Sigma_{n}^{\prime}$ is a stable minimal surface with respect to the shrinker metric.
+Now $\Sigma_{n}^{\prime}$ is a surface whose boundary is in $\partial B(0,r_{n})$. It is smoothly embedded away from the boundary, and its shrinker area is at most
 $\operatorname{area}_{s}(\Sigma_{n})=\operatorname{entropy}(\Sigma_{n})\leq\Lambda.$
-Now the $\Sigma_{n}^{\prime}$ are stable, so (after passing to a subsequence, they converge smoothly in $\mathbf{R}^{3}\setminus\{0\}$ to a stable shrinker $\Sigma^{\prime}$ (possibly with multiplicity).
-By [gulliver-lawson], $\Sigma^{\prime}$ is smoothly embedded everywhere. But then $\Sigma^{\prime}$ is unstable, a contradiction.
+Since the $\Sigma_{n}^{\prime}$ are stable, they converge smoothly (after passing to a subsequence) in $\mathbf{R}^{3}\setminus\{0\}$ to a stable shrinker $\Sigma^{\prime}$ (possibly with multiplicity).
+By [gulliver-lawson], $\Sigma^{\prime}$ is smoothly embedded everywhere.
+But then $\Sigma^{\prime}$ is unstable, a contradiction.
+(Every smooth shrinker without boundary is an unstable minimal surface with respect to the shrinker metric.)
 The proof for $R$ is essentially the same. ∎
-11. Genus of Shrinkers
+Appendix A Genus of Shrinkers
 Chodosh, Choi, and Schulze proved that that all shrinkers at a spacetime point have the same genus [cms]*Proposition H.8. Here, we give a different proof that is perhaps somewhat simpler.
 Lemma 66.
-Suppose that $T\leq\operatorname{T_{\textnormal{fat}}}$ and that $p\in M(T)$. Let $\mathscr{S}$ be the collection of all shrinkers at $(p,T)$, i.e., the the set of all subsequential limits of
+Suppose that $T\leq\operatorname{T_{\textnormal{fat}}}$ and that $p\in M(T)$. Let $\mathscr{S}$ be the collection of all shrinkers at $(p,T)$, i.e., the set of all subsequential limits of
 $(T-t)^{1/2}(M(t)-p)$
@@
 is a Morse function, and the only critical points are local maxima.
-(In other words, at each critical point $q$, both principal curvatures with respect to the inward pointing normal are $>1/|q|$.)
+(In other words, at each critical point $q$, both principal curvatures with respect to the inward pointing normal are >1/|q|>1/|q|.)
 Proof.
 Suppose not.
-Then there exist $\Sigma_{i}\in\mathscr{S}$, and critical points $q_{i}\in\Sigma_{i}$ such that $R_{i}:=|q_{i}|\to\infty$ and such that one of the principal curvatures at $q_{i}$ (with respect to the inward pointing normal) is $\leq 1/|q_{i}|$.
+Then there exist $\Sigma_{i}\in\mathscr{S}$, and critical points $q_{i}\in\Sigma_{i}$ such that $R_{i}:=|q_{i}|\to\infty$ and such that one of the principal curvatures $\kappa_{i}$ at $q_{i}$ (with respect to the inward pointing normal) satisfies:
+(54) $\kappa_{i}\leq 1/|q_{i}|.$
 Consider the tangent flow:
 (*) $t\in(-\infty,0)\mapsto|t|^{1/2}\Sigma_{i}.$
-The mean curvature of $\Sigma_{i}$ at $q_{i}$ is $-q_{i}/2$. Thus the mean curvature of $R_{i}\Sigma_{i}$ at $R_{i}q_{i}$ is $-q_{i}/(2|q_{i}|)$.
-Now translate the flow (*11) in spacetime by $(-R_{i}q_{i},R_{i}^{2})$ to get
+The mean curvature of $\Sigma_{i}$ at $q_{i}$ is $q_{i}/2$. Thus the mean curvature of $R_{i}\Sigma_{i}$ at $R_{i}q_{i}$ is $q_{i}/(2|q_{i}|)$.
+Now translate the flow (*A) in spacetime by $(-R_{i}q_{i},R_{i}^{2})$ to get
 $t\in(-\infty,R_{i}^{2})\mapsto$ $|R_{i}^{2}-t|^{1/2}\Sigma_{i}-R_{i}q_{i}$
@@
 $=|(1-(t/R_{i}^{2})|^{1/2}R_{i}(\Sigma_{i}-q_{i})+|(1-(t/R_{i}^{2})|^{1/2}R_{i}q_{i}-R_{i}q_{i}$
-$=|(1-(t/R_{i}^{2})|^{1/2}R_{i}(\Sigma_{i}-q_{i})-(t+o(t))\frac{q_{i}}{|q_{i}|}$
+$=|(1-(t/R_{i}^{2})|^{1/2}R_{i}(\Sigma_{i}-q_{i})-\frac{1}{2}(t+o(t))\frac{q_{i}}{|q_{i}|}$
 By [bamler-kleiner], this converges (perhaps after passing to a subsequence) to a flow that is regular at almost all times, and the convergence is smooth for almost all $t$.
@@
 (55) $t\in\mathbf{R}\mapsto\Sigma^{\prime}-tq$
-is mcf, i.e., such that $\Sigma^{\prime}$ is a translator with velocity $-q$. Thus the mean curvature at each point of $\Sigma^{\prime}$ is $-q^{\perp}$.
+is a mean curvature flow, i.e., such that $\Sigma^{\prime}$ is a translator with velocity q-q.
+Here q-q is the limit of $q_{i}/(2|q_{i}|)$.
+Thus the mean curvature at each point of $\Sigma^{\prime}$ is $q^{\perp}$.
 Let
 $t\in(-\infty,0)\mapsto|t|^{-1/2}\Sigma^{\prime\prime}$
-be tangent flow at $\infty$ to the flow (43). Thus, there are $t_{n}\to-\infty$ such that
+be a tangent flow at $\infty$ to the flow (55). Thus, there are $t_{n}\to$-$\infty$ such that
 $\Sigma^{\prime}_{n}:=|t_{n}|^{-1/2}\Sigma^{\prime}(t_{n})=|t_{n}|^{-1/2}(\Sigma^{\prime}-t_{n}q)$
@@
 By [bamler-kleiner], the convergence is smooth with multiplicity one.
-Note that the mean curvature vector $H_{n}$ at each point of $\Sigma^{\prime}_{n}$ is $(-t_{n}q)^{\perp}$.
+Note that the mean curvature vector $H_{n}$ at each point of $\Sigma^{\prime}_{n}$ is $(-|t_{n}|^{1/2}q)^{\perp}$.
 Thus
-$|H_{n}|=|\nu\cdot H_{n}|=|t_{n}|\,|q^{\perp}|,$
+$|H_{n}|=|\nu\cdot H_{n}|=|t_{n}|^{1/2}\,|q^{\perp}|,$
 or
-$|q^{\perp}|=|t_{n}|^{-1}|H_{n}|.$
-Letting $n\to\infty$, we see that all points $q$ of $\Sigma^{\prime\prime}$, $|q^{\perp}|=0$.
-That is, $\Sigma^{\prime}$ is translation-invariant in the $q$ direction.
-Thus $\Sigma^{\prime}$ is the shrinking cylinder, and therefore $\Sigma^{\prime}$ is the bowl soliton [chh].
-But that is a contradiction. ∎
+$|q^{\perp}|=|t_{n}|^{-1/2}|H_{n}|.$
+Letting $n\to\infty$, we see that at all points of $\Sigma^{\prime\prime}$, $|q^{\perp}|=0$.
+That is, $\Sigma^{\prime\prime}$ is translation-invariant in the $q$ direction.
+Thus $\Sigma^{\prime\prime}$ is the shrinking cylinder, and therefore $\Sigma^{\prime}$ is the bowl soliton [chh].
+But that is a contradiction, since one of the principal curvatures of $\Sigma^{\prime}$ at 0 with respect to the direction q-q is $\leq 0$, by (54). ∎
 Corollary 67.
@@
 Proof.
-By the lemma and by standard Morse Theory, each component of $M\setminus B(0,r)$ is diffeomorphic to a closed disk or to $S^{1}\times[0,1)$. (Here $B(0,r)$ is the open ball.) ∎
+By Lemma 66 and standard Morse Theory, each component of $M\setminus B(0,r)$ is diffeomorphic to a closed disk or to $S^{1}\times[0,1)$. (Here $B(0,r)$ is the open ball.) ∎
 Theorem 68.
@@
 $\operatorname{genus}(\Sigma_{i})=\operatorname{genus}(\Sigma)$
-for all sufficiently large $i$. Let $R$ be as in the Lemma, and let $r>R$ be such that $\partial B(0,r)$ intersects $\Sigma$ transversely.
+for all sufficiently large $i$. Let $R$ be as in Lemma 66, and let $r>R$ be such that $\partial B(0,r)$ intersects $\Sigma$ transversely.
 Thus $\partial B(0,r)$ intersects $\Sigma_{i}$ transversely for all sufficiently large $i$, say $i\geq I$.